A Complete Bibliography of Publications in American Documentation

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

28 March 2025 Version 1.05

Title word cross-reference

#9 [Web65].

$ [SM67]. $.75 [Ano64b]. $1.50 [Art65, Har63]. $10.00 [Wad68, Cha63, Tau61a]. $10.60 [New63a]. $12.00 [Har68c, Boh64]. $12.50 [Ano64c, Gar65a]. $12.95. [Ano68e]. $15.00 [Tru67, Ano68g, Try69]. $18.75 [Ano69b]. $2 [Boh62]. $2.50 [Ano68f, Lit64]. $20.00 [Tru67]. $25.00 [Bal69]. $3.50 [Har68a]. $30.00 [Eva66]. $390 [Ric68]. $4.00 [New69a, Pel69]. $4.50 [Tru67, Ano69c, Boh65]. $4.95 [Hin69]. $460 [Ric68]. $5.00 [New63b, Vas69]. $5.50 [Ano69a]. $6 [Ric68]. $6.00 [Dai64]. $7.50 [Pra63]. $7.50. [FC68]. $8.50 [Ano69f]. $8.50. [Ano68h]. $8.75 [Ano65a]. $8.95 [Tau65a]. $9.00 [Ano69e]. ^{1} [Sha65]. ^{2} [Sha65]. ^{3} [Sha65].

00 [Ano64b]. 000 [Ano64c, Ano64d, Dai64, Lit64].

1 [Ano66a, Ano68d, Ano68c, Ano68b, Ano68a, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ash66, Boh64, Emb65, Har63, Har67c, Hay66a, Lan65b, Lan67, O’C63, Pra63, Tau65a, Tre66, Vai67a, Vai67b]. 1/65 [Emb65, Lan65b, Tau65a]. 1/65-1R [Tau65a]. 1/65-2R [Lan65b]. 1/65-3R [Emb65]. 1/66 [Ano66a, Ash66, Hay66a, Tre66]. 1/66-1R [Ash66]. 1/66-2R [Ano66a]. 1/66-3R [Hay66a]. 1/66-4R [Tre66]. 1/67 [Ano68d, Ano68c, Ano68b, Ano68a, Har67c, Lan67, Vai67a, Vai67b]. 1/67-1R [Ano68a, Lan67]. 1/67-2R [Ano68b, Vai67a]. 1/67-3R [Ano68c, Vai67b]. 1/67-4R [Ano68d, Har67c]. 1/69 [Ano69a, Ano69b, Ano69c, Ano69d]. 1/69-1R [Ano69a]. 1/69-2R [Ano69b]. 1/69-3R [Ano69c]. 1/69-4R [Ano69d]. 10 [Ano65a, Art65, Boh65, Dar66, Eva66, Har63, Hay66b, Hil66, New63b, Per66a, Sey66]. 10/63 [Har63, New63b]. 10/63-1 [Har63]. 10/63-2 [New63b]. 10/65 [Ano65a, Art65, Boh65]. 10/65-1R [Art65]. 10/65-2R [Boh65]. 10/65-3R [Ano65a]. 10/66 [Dar66, Eva66, Hay66b, Hil66, Per66a, Sey66]. 10/66-1R [Per66a]. 10/66-2R [Sey66]. 10/66-4R [Hil66]. 10/66-6R [Hay66b]. 11th [Lit66]. 177 [PV63]. 1950 [Cle63]. 1955 [SD55]. 1959 [JT61a]. 1960 [Gul61]. 1961 [Ano61a, Ano62h, Boh62]. 1962 [Ano62a]. 1964 [Ano65b, Mou65]. 1965 [Eva66, Sny66]. 1966 [Gre67, Har67b, Tru67, Wad68]. 1R [Ano68e, Ano68f, Har68a, Ano68a, Ano69a, Art65, Ash66, Bat67a, Gar65a, Har67a, Hay67, Lan67, Met65, Per66a, Rub66, Tau65a, Try69, Vas69].

2 [Ano69e, Ano69f, Bat67a, Cla67, Gre67, Har69, Har67b, Mou67, New63b, Pel69, She67]. 2/67 [Bat67a, Cla67, Gre67, Har67b, Mou67, She67]. 2/67-1R [Bat67a]. 2/67-2R [She67]. 2/67-3R [Mou67]. 2/67-4R [Har67b]. 2/67-5R [Cla67]. 2/67-6R [Gre67]. 2/69 [Ano69e, Ano69f, Har69, Pel69]. 2/69-1 [Ano69e]. 2/69-2 [Har69]. 2/69-3 [Pel69]. 2/69-4 [Ano69f]. 2d [Lan67, Ran67]. 2nd [Bal69]. 2R [Ano68h, Ano68g, Lew66, Ric68, Ano65b, Ano66a, Ano68b, Ano69b, Ath65b, Bal69, Boh65, Lan65b, Lin66, Rad69, Rog67, Sey66, She67, Tru67, Vai67a].

3 [Bal69, Har68a, Har68b, Hay67, Log67, New69a, Pel69, Pre69, Ric68, Tru67, Try69]. 3/67 [Hay67, Log67, Tru67]. 3/67-1R [Hay67]. 3/67-2R [Tru67]. 3/67-3R [Log67]. 3/68 [Har68a, Har68b, Ric68]. 3/69 [Bal69, New69a, Pre69, Try69]. 3/69-1R [Try69]. 3/69-2R [Bal69]. 3/69-3R. [Pre69]. 3/69-4R [New69a]. 38th [She67]. 3R [Dar66, FC68, Har68b, Har68c, Whi66, Ano65a, Ano68c, Ano69c, Emb65, Hay66a, Lan65a, Log67, Mou67, Ran67, Ste69, Vai67b, War66, Web65]. 3R. [Pre69].

4 [Ano64b, Ano64c, Ano64d, Ano68e, Ano68h, Ano68f, Ano68g, Ano69f, Anz67, Ath65b, Bra67, Dai64, FC68, Gar65a, Har68c, Har67a, Hin69, How65, Lan65a, Lin66, Lit64, Nit67, O’C63, Rad69, Ran67, Rog67, Rub66, Ste69, Vas69, Wad68, War66]. 4/63 [O’C63]. 4/63-1 [O’C63]. 4/64 [Ano64b, Ano64c, Ano64d, Dai64, Lit64]. 4/64-00 [Ano64b]. 4/64-000 [Ano64c, Ano64d, Dai64, Lit64]. 4/65 [Ath65b, Gar65a, How65, Lan65a]. 4/65-1R [Gar65a]. 4/65-2R [Ath65b]. 4/65-3R [Lan65a]. 4/65-4R [How65]. 4/66 [Lin66, Rub66, War66]. 4/66-1R [Rub66]. 4/66-2R [Lin66]. 4/66-3R [War66]. 4/67 [Anz67, Bra67, Har67a, Nit67, Ran67, Rog67]. 4/67-1R [Har67a]. 4/67-2R [Rog67]. 4/67-3R [Ran67]. 4/67-4R [Nit67]. 4/67-5R [Anz67]. 4/67-6R [Bra67]. 4/68 [Ano68e, Ano68h, Ano68f, Ano68g, FC68, Har68c, Wad68]. 4/69 [Hin69, Rad69, Ste69, Vas69]. 4/69-1R [Vas69]. 4/69-2R [Rad69]. 4/69-3R [Ste69]. 4/69-4R [Hin69]. 432198 [Lan65a]. 45 [Har69]. 48 [O’C63]. 4R [Lit66, Wad68, Ano68d, Ano69d, Har67c, Har67b, Hil66, Hin69, How65, New69a, Nit67, Tre66]. 4th [FC68].

500 [Ric68]. 5R [Eva66, Anz67, Cla67].

63-1 [Har63, O’C63, Pra63]. 63-2 [New63b]. 64-00 [Ano64b]. 64-000 [Ano64c, Ano64d, Dai64, Lit64]. 65-1R [Art65, Gar65a, Met65, Tau65a]. 65-2R [Ano65b, Ath65b, Boh65, Lan65b]. 65-3R [Ano65a, Emb65, Lan65a, Web65]. 65-4R [How65]. 6570th [Ano64b]. 66 [Dar66, Eva66, Lew66, Lit66, Whi66]. 66-1R [Ash66, Per66a, Rub66]. 66-2R [Ano66a, Lin66, Sey66]. 66-3R [Hay66a, War66]. 66-4R [Hil66, Tre66]. 66-6R [Hay66b]. 66-IR [Sny66]. 67-1R [Ano68a, Bat67a, Har67a, Hay67, Lan67]. 67-2R [Ano68b, Rog67, She67, Tru67, Vai67a]. 67-3R [Ano68c, Log67, Mou67, Ran67, Vai67b]. 67-4R [Ano68d, Har67c, Har67b, Nit67]. 67-5R [Anz67, Cla67]. 67-6R [Bra67, Gre67]. 68 [Ano68e, Ano68h, Ano68f, Ano68g, FC68, Har68a, Har68b, Har68c, Ric68, Wad68]. 69-1 [Ano69e]. 69-1R [Ano69a, Try69, Vas69]. 69-2 [Har69]. 69-2R [Ano69b, Bal69, Rad69]. 69-3 [Pel69]. 69-3R [Ano69c, Ste69]. 69-3R. [Pre69]. 69-4 [Ano69f]. 69-4R [Ano69d, Hin69, New69a]. 6R [Bra67, Gre67, Hay66b].

7 [Ano65b, Lew66, Lit66, Met65, Pra63, Sny66, Web65, Whi66]. 7/63 [Pra63]. 7/63-1 [Pra63]. 7/65 [Ano65b, Met65, Web65]. 7/65-1R [Met65]. 7/65-2R [Ano65b]. 7/65-3R [Web65]. 7/66 [Lew66, Lit66, Sny66, Whi66]. 7/66-IR [Sny66].

în [Sta61].

A.I.Ch.E. [Mor62]. a/M [Ste69]. Abbott [Ano64b]. abbreviating [Sav63]. Abbreviations [Fis59]. ABC [Alt67, AR69]. Aberdeen [KP57]. abilities [Man66]. abridged [Kut64]. abroad [Kot62]. Abstract [CM62, Del64, MK62]. abstracter [BC63]. abstracters [Pla60b]. Abstracting [Ath65b, Mil57, Bou62, CHJ61, Doy62, KP57]. Abstracts [Bro55, Pet59, BC63, Kna65, LD60, RRS61b, Pet59, Pla60b, Spe67, Van65, WA62]. absurdum [McG58b]. AC [Mit53]. academe [DK66]. Academic [Gar65a, Lun69]. Academy [JT61a, SB59, Sta61]. acceptable [BC63]. Access [Dra54, Ben69, JS69, Man69, Nol59]. accounting [Hel69, Ohl57]. achievements [Bey60b, Tat50b]. achieving [Per65]. acquisition [Mil53, Orn63]. Acquisitions [Jac52]. Actifilm [Jon61]. action [Coo68, Jon61]. activities [Ano62y, Cla56, Mat62, Mor62, Par62, Pol50, Tar62]. AD [Lan65a, McG58b, Eli69c]. Adaptation [WD59]. Adaptive [SAF65]. Adatok [Emb65]. Addison [Gar65a]. Addison-Wesley [Gar65a]. address [Sch62a]. addressed [Flo67]. ADI [Ano62x, Ano52e, Ano52g, Ano62b, Ano62h, Ano62t, Ano62u, Ano62-28, Ano62-29, Bin63, Cua65, Eva50]. Administration [Ano68h, Boh62, Lit64, Pra63, Sco51]. Advanced [Cha63]. Advisory [Ano55a]. AEC [Gra51]. Aeronautics [Ano68h]. Aerospace [Ano64b]. AFB [Lan65a]. Affairs [Har67c, Lit66]. affected [Sop68]. age [Daw68, DK66, Jon54]. agency [Bar56]. agreements [O’C69]. ahead [Hei61b]. aid [Ben69, Cro56, Luh61]. aids [BL55, Sla63]. AIP [Ano65b, Mou65]. AIP/DRP [Ano65b, Mou65]. AIP/DRP-CI [Mou65]. Air [Ano64b, Lan65a, TN68]. Alberto [BF57]. alerting [BBJB69]. Alexander [FC68]. algorithm [Dav69]. Allen [Ano68c, Rub66]. allied [Ken60]. alpha [Fei63]. alphabetic [JS69, Sha62]. alphabetical [Sla63, Sla64, Ber56a, Lew66]. also [Mos56, Tru67]. Alto [Gar65a]. amenable [KBP54b]. America [Bis59]. American [Ano65b, Ano68g, Ano69f, Ath65b, Bat67a, Boh62, Cla67, Gul61, La 63, Lit64, Mou65, Nit67, Try69, Web65, Ano53a, Ano58i, Ano61a, Ano62a, Ano62n, Ano63a, Ano64a, Ano65j, Ano67g, Ano68p, Ano68q, Bal50c, Bor55, Bur59a, Bur59b, Bur59c, Den50, Dub62, Eva51, Kin59, McM61, Nit67, SG69, Urq51, WBO^{+}61, Tat50b]. Americas [Ath65b]. Ami [Ano65b, Mou65]. among [BH57b, Kot62]. AMRL [Ano64b]. AMRL-TDR-63-8 [Ano64b]. analogue [Sch64b]. Analogy [HG65]. Analysis [BH67, Flo67, Wåh66, Alt67, Ano68g, BF64, BE65, CR69, CHJ61, Gar54, Hoy69, Kle64, Lan69b, Lei69, Mac53, MS66, MS67, Mor54, Ove63, Per50, PM59, SS61, Sti60, WP62]. analytic [Kes65]. analytical [Car69]. analyze [Yag65]. Andre [Ano65a]. Anglo [Nit67]. Anglo-American [Nit67]. animadversions [Bor54]. Anleitung [Ste69]. anniversary [Ano63k]. Annotated [Gre67, LK56]. announced [Tru67]. Announcement [Ano54o, Ano54p, Ano62b, Bra51, Car69]. Announcements [Ano52a, Ano53a, Ano53b, Ano53c, Ano54a]. Annual [Ano53a, Gul61, She67, Tat54, Ano52e, Eva51, ML67, Tru67, Ano62h, Ano68a, Bat67a, Try69]. answer [O’C68b, O’C69, Tag65]. answer-providing [O’C68b, O’C69]. Answerer [Lan65b]. answering [SKM64]. answers [Sch69a]. Anthony [Tau61a]. appearing [Ath64]. Appendix [Ath65b]. Application [Ano55a, HH60, MH53, Fie56, Lip60, Lit64, Sti58, TMK55, Uhl64, Yag65, Gar65a]. Applications [Ano69c, McC61]. applied [Moo51]. appraisal [Tat50a]. appreciation [Sew62]. approach [Bis53, Hei61a, JN58, Lib69, Orr59, PKB54, PKM59, Slo68, TN68]. approaches [Man61]. approaching [BH63]. April [Ano66a, Lan65b]. arbitrary [Uhl69]. Archival [Rid50]. archive [How65]. Archon [Har68a, Har68c, Vas69]. Area [Wad68, LBL67, MW61, Wad68]. areas [MK62]. Armed [Bar56]. army [Ano52l]. array [Per65]. Art [Art65, Ano62-30, Bou61, Mar62, Mar66]. Artandi [Bra67, Met65, Pel69]. Arthur [Har67a, Ano60m]. Articles [Ano52b, Bis56, Gro68, Lei69, Pet59]. articles-cited [Gro68]. articulation [Wal68]. Arvid [FC68]. Asa [Hay67]. Asia [Eva66, Mat62]. asking [Tay62]. ASLIB [Hay66b, LM64, She67, Ree65]. ASM [Fos58]. aspect [Vic60]. aspects [Ber54, Man64, OE58]. assembly [McC68]. assessment [Bey60b]. assistance [Ano69a, Cob56]. associated [Pot68]. Associates [Ano69d, Web65]. Association [Ano62d, Gre67, Har63, La 63, Nit67, Ran67, Web65, Ass55, Doy62, JC67, Ano62o, Hys53, Tat54]. associations [Les69, TPBK55]. associative [Che63, DD65, Sla63]. ASTIA [Gra51]. Atherton [Gro67, Sew62]. attacking [Wal67]. Attendance [BGM56]. Attitudes [FCHB65]. attributes [Man69]. Aurel [Sta61]. Australian [Dra54]. author [Bra64, Lec68, Ano66a, Ano62c]. authority [Ath65b, Way68]. authors [Ano62p, SSO65]. Automata [Har67a]. Automated [BBJB69, KB68, Whi66, BH67, Rau67, Ano68e]. Automatic [Art64, CHJ61, Lan65b, PKM59, Bor67, Dam65, Gar55, GG57, Les60, Opl59, PKB55a, Sal63, Sal65b, Sal69, Sha56, Art65, Tau61a]. automatically [AW69]. Automation [LK56, Pla60a, Har69]. auxiliary [Ano62t, Dav51b, Dav51a, Dav52, Ano62v]. Availability [Fag52, Can52]. Avenue [Ano69d, Ath65b]. Avramescu [Sta61].

B [Lan65b, Lan67]. Babbage [Gar69]. back [Tob59a]. Bar [Gar65a, Man64, Moo57]. Bar-Hillel [Gar65a, Man64, Moo57]. Barbara [Ano62d]. Barber [Har67c]. Barden [Ano62d]. Barhydt [Ano69a]. Barnard [Ano64b]. Barnett [Sey66]. Barrer [Gre67]. bars [Cum54]. Base [Ano64b, New61, TBW68, Ike67]. Baseball [Lan65b]. based [Coo68, FL69, Rop68, Vas69]. basic [Uhl68, VZ68, New63a]. Batty [Har68a]. be [Har63, Tru67]. Beatriz [Tre66]. Becker [Boh64]. beginning [Lec68]. beginnings [Emb65]. behavior [Kap65]. behind [SS59]. Ben [Ano65b, Mou65]. Ben-Ami [Ano65b, Mou65]. benchmark [Man69]. benefits [OE58]. Bensing [O’C63]. Bereich [Ste69]. Bernard [Vas69]. Bertha [Ano62d]. Bethlehem [Mou67]. between [HG65, JT61b, Kes63, Oat68, Sal69]. Bibliographic [Cla67, Ega51, Kes63, Kin59, LH69, She50, She51, SB59, War52, Ano51a, Bro52, Cla56, Fel57, Har68c, Jor59, Kes65, Min57, Min60, Pin57, SE50b, Sti58]. bibliographical [Bey60b, Cam56, Car54, Ham68, Que63]. Bibliographies [New69b]. Bibliography [Dar66, Gre67, Ste69, Ano62c, Boh62, Bro53a, Bro53b, Bui62, Emb65, Lit64, LK56, MM59, Spe67, Ano62w]. big [Sch62b]. binary [Dav69, Man69]. binding [Pow51a]. Bingley [Har68a, Vas69]. Biographical [Sch69b]. Biological [Ash66, Boh62, Har67b, Wei66, Heu56b, Woo57]. biologist [Lei69]. biology [Woo57]. Biomedical [Ano64b, Anz67, Lei69, Rop68, SSO65, SK63, FL69]. Biweekly [Tru67]. blank [Ber58, Har68b]. blocks [KT67]. Bobbs [Ano69c]. Bobbs-Merrill [Ano69c]. Boeing [WS64]. boo [Ano62q, Uhl64]. Book [Ano62d, Ano62c, Ano64b, Ano64c, Ano64d, Ano65b, Ano65a, Ano66a, Ano68d, Ano68c, Ano68b, Ano68a, Ano68e, Ano68h, Ano68f, Ano68g, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Anz67, Art65, Ash66, Ath65b, Bal69, Bat67a, Boh62, Boh64, Boh65, Bra67, Cha63, Cla67, Cle63, Dai64, Dar66, Emb65, Eva66, FC68, Gar65a, Gre67, Har66, Har68a, Har68b, Har68c, Har69, Har67a, Har67c, Har67b, Hay66a, Hay66b, Hay67, Hil66, Hin69, How65, JT61a, KT67, La 63, Lan65b, Lan65a, Lan67, Lew66, Lin66, Lit64, Lit66, Log67, Mac62b, Met65, Mou65, Mou67, New63a, New69a, Nit67, Pel69, Per66a, Pre69, Rad69, Ran67, Ric68, Rog67, Rub66, Sey66, She67, Sny66, Sta61, Ste69, Tau61a, Tau65a, Tre66, Tru67, Try69, Vai67a, Vai67b]. Book [Vas69, Wad68, War66, Web65, Wel61, Whi66, Art64, Lec68, Rid51a, Sha53, WS64]. Book-indexes [KT67]. Books [Har66, Har68a, Har68c, New63a, Vas69, Kil61, Pou56, Pow58, Har66]. boolean [Bra66, MB68, Ike67]. border [Coh59]. border-punched [Coh59]. Borko [Ano68e]. Boston [Ano62h]. bound [La 63, Tru67]. Bourne [Ano64c]. Bowles [Ano68d]. Box [Har63]. Boyle [Sav65]. Bracken [Har67b]. Bradford [Gro67]. breakthrough [Pow58]. Brian [Per66a]. Brief [And64, Ano65c, Ano67a, Ath64, BE65, Bra66, Bro68, Dav68, Dav69, FL69, Fla69, Fra66, Gro67, Gro68, Hol66, Hur69, Ike67, JH66, Jen65, Ken64, Kik67, KC69, Kna65, Kor64, Kut64, Lip65a, Man69, MB68, O’C64a, Per65, Pie64, PV67, Rau67, Ric67, Ros65a, Ros65b, Sav65, Sav66, Sha66, SM67, SR68, Swa65, Swa68, Swa69, Tau68, TJ69, TEC64, Van65, WW67, Woo64, Wor69, Mac60a, Que63]. Britannica [Try69]. British [Den50]. broad [Ano62q, Per53]. broad-subject [Ano62q]. Brown [Ano68g]. browser [Tre68]. Brunswick [Bra67, Cle63, Lew66, Met65, Per66a, Rog67]. Bryan [Ano69f]. Bucharest [Sta61]. Budapest [Emb65]. building [KT67, Lar52, Sch54, Ste69, Wal69]. buildup [LBL67]. Bull [Lit64]. burden [She59]. Bureau [Ano64d, Ano68f, Art65, McM62]. buried [Smi55]. Burke [FC68]. Burks [Har67a]. Business [Ano62h, Whi66, Cue68, Ega53, MH53]. Butterworth [Lan67]. Butterworths [Lin66]. by-laws [Ano52b]. bylaws [Ano62a].

C [Ano69a, Boh62, Bra67, Har68a, Har67b, Lan65b, Lan67, Met65, O’C63, Per66a, Vai67a, Vai67b, Sha65]. Cabot [Dav58]. calculation [Man69]. Call [Ano67b, Ano68i]. Calvin [Wis52]. Cambridge [Ano68c, Sey66, Sny66, Tau61a]. camera [Blu50, FS57, Sha50]. Canadian [Nit67]. cancer [Hof52, SPS68]. cancer-virology [SPS68]. Candea [Sta61]. capacity [Bra62]. Carbon [Ano57n]. Card [Web65, Dew59, EFB57, Gar54, Gar55, Gil58, HH60, Jon60, Kuv67, Lew69, MH53, Mal58, McM61, MM59, PW58, Pou56, Seb53, Wac52, Gra51]. cardiovascular [OC62b, PW58]. cards [BHR67, Coh59, Heu56b, Hof52, Lip60, Mor60, Ohl57, Per51, SK63, Uhl64]. Carl [Ano64b, Sny66]. Carlos [Ano68a, Bat67a, Try69]. Carter [FC68]. Case [Ano53b, Ano69a, HH53, Lar52, Lec68, TEC64, Tru67]. case-bound [Tru67]. case-problem [Lar52]. Case-reserve [Ano53b]. cash [Nee62]. CATALOG [Gra51, Ano68f, Har66, Tru67, Cla56, Gul51, KL67, Kuv67, Lew69, Lil54, Ric56, Sch57, Wad68, Web65, Har66]. cataloger [Ben69]. Cataloging [Aro52, Car52, Fra52, Nit67, Ano54f, Bra64, Mas54]. catalogs [Dew59, Eva52, WS64]. Catalogue [Ano66a, Kil66, SK63, Ano66a]. cataloguing [All59, Har68b]. Categories [Mos64]. cell [Lei69]. cell-biologist [Lei69]. census [GD63]. Center [Ano68f, Ash66, Boh62, Emb65, Fuk67, Lit64, Pol50, Wad68, BH67, Cot68, FCHB65, Hel69, Shi63, Wil52, Ano55b, Ano57a]. centers [Lun69, New61, Sim62]. Central [Ath65b, Bin63]. Centre [Par62]. century [Kno60]. certain [Hin61]. chain [Uhl64]. Chang [Ano69a]. change [Daw68]. Changeable [Cum54]. chapter [Bin63]. character [JT61b]. Characteristics [Ano69e, JHG66, CBF59, Mil52]. Characterized [Wåh66]. characters [Pri69, SK63]. Charlene [Anz67]. Charles [Ano64c, Ano69a, Har67b, New63a, Gar69]. Checklist [Ran67]. Chemical [Aue55, Mac58, Ruh64, Ber57b, CBF59, Cos61a, DG57, Fel64, Opl59, OB59, Rot56, WW67, Woo57, Bur59a, Pet59, Pla60b, Spe67, Van65]. chemical-biological [Woo57]. chemical-group [Ber57b]. chemist [Orn54]. chemists [Mai60, Str57]. Chicago [La 63, Web65, Ano51a, Nit67]. choices [GB69]. Chomsky [Lan65b]. Chronological [Rei54]. CI [Ano65b, Mou65]. CI-3 [Ano65b]. ciphers [DG57]. circulated [Pin67]. circulation [Tru65]. Citation [Ada55, Ano65b, Gar63, Spe67, GS63, Kap65, Lip62, Lip65b]. Citations [Bur59a, Bur59b, Bur59c, Yag65, Mou65]. cited [Ada67, Gro68]. City [Har63, Hil66]. Claire [Ano69b, New69a]. Clapp [Boh65]. Class [PBK54]. Classifica [Ste69]. Classification [Gre67, Per66a, Sch54, Ste69, Wåh66, Ano62d, AC61, BN69, Bor64, Dai57, Dav68, Fai58, Fos58, Har54, Hof52, Hys53, Kra64, Kyl58, Mos64, Ric54, Ric61, Str56, Tya61, Har68a, Mac62b]. classifications [Hay57]. Classificatory [Sla63]. Classified [New61, Bow54]. clearinghouse [Bau66]. Cleveland [Ano69a, Hay66a, Swa65]. Cliffs [Ano68d]. clinical [HH60, SL68]. Clive [Har68a, Vas69]. clumping [DD65]. Co [How65]. Coblans [Hay66b]. code [Gil58, KPB54, McC65, Mel62, PW58, PV63, PBK54, Tau62, WAF^{+}58, Woo57]. CODEN [Bis54, Bis57]. codes [Fel64, Per51]. Coding [RSF^{+}67, Buc61, GG57, Lam61, Sti60, Tau62, Uhl64, Vic59b, WP50, Wis52]. Cohan [Har63]. collected [Lei69]. collection [Ano62d, Ano62q, Den50, Eva52, Fla69, Har51, Man61, Orn52, Seb53, BPK54]. collections [Har68b, Met51]. College [FC68]. Collins [New63b]. Colo [New63b]. colon [Har68a, Jon54, Tya61]. Colonel [Lan65a]. Colorado [New63b]. Columbia [Dai64, FC68, CG62]. Columbus [New63a]. column [O’C62a]. Combined [Per68]. combining [Har63]. comes [Jon54]. Coming [DK66]. Command [Ano64b]. Comments [Ano57a, Cua65, Moo57, Vic51, Wol59]. commercial [CBF59, Mar66]. Committee [Ano50a, Ano52l, Ano55a, Ano59m, Ano62u, Ano62-27, Ano63a, Bal50a, Bal50b, Dar66, HMC63, Hys53, Rey61, Ano58i, Ano62v, Ano62w, Ano62x, Ano62y, Ano62z, Ano65j, Tho62]. committees [Ano62-29]. common [CG62]. Communication [Ano51j, Ano55b, Ano57a, Boh62, Har67b, Hei61a, Kot62, New69a, Cla53, HH53]. Communications [And64, Ano65c, Ano67a, Ath64, BE65, Bra66, Bro68, Dav68, Dav69, FL69, Fla69, Fra66, Gro67, Gro68, Hol66, Hur69, Ike67, JH66, Jen65, Ken64, Kik67, KC69, Kna65, Kor64, Kut64, Lip65a, Man69, MB68, New63b, O’C64a, Per65, Pie64, PV67, Rau67, Ric67, Ros65a, Ros65b, Sav65, Sav66, Sha66, SM67, SR68, Swa65, Swa68, Swa69, Tau68, TJ69, TEC64, Van65, WW67, Woo64, Wor69, Ano50f, Ano50g, Ano50h, Ano51k, Ano51j, CS62, Cla50, Lar52, New63b]. Company [Ano62c, Ano69c, Ano69f, Gar65a, Har68b, Hin69, Ran67, Ano57n]. Comparative [SSO65, Ano69f, Cle63, Kle64, Ste61]. comparing [JC67, Jor68, Kna65]. Comparison [Kes65, JS69, Kra64, Lyn69, Mon65, Pol68, RB68, Sal69, SSS61, Flo67]. Comparisons [RRS61a]. compatibility [Hys63, Lip63]. Compilation [Lip62, Sla64]. Compiled [Ano62d, Dar66, Gre67]. Complete [Har66, Fel64]. composition [Cum54, Sil50a]. compounds [CBF59, WAF^{+}58]. Comprehensive [Sag66]. compression [Cot68, Lyn69]. Computational [Ano68g]. Computer [Ano62c, Ano69f, BKB62, BK63, BML64, Bes61, Car68, CR69, Cos62, FL69, Hay67, HH67, KL67, McC68, PKM59, WS64, Ano68c, Art64, BF58, BKRL66, Fel64, Kuv67, Man69, MS66, MS67, OB59, Rop68, SS61, SS62, Sch63, SM63, WGP58, Way68, Wor69, Sey66, Tau65a, Vas69, Har68c]. computer-based [Rop68, Vas69]. Computer-produced [KL67]. computer-program [BKRL66]. Computerized [TJ69]. Computers [CG62, Kil66, Mar66, McC62, Pri69, Ano68d, Pel69]. concept [Cos61c, DO69, Fis66, Mon62, Mor54, Ruh64, Wad58]. concerning [O’C68a]. concordance [Bes61]. conditions [Sti58]. Conference [Ano55c, Mou67, She50, Sny66, Ano51a, Ano53b, Bel69a, Bel69b, JT61a, SD55, Sla69, Swa65, She67]. conferences [Bro52]. confluence [Man69]. Congress [Ash66, Nit67, Ano62p, Hol50, Hol51, Cla55, Den50, Gul51]. Congresses [Ano62o]. Connecticut [Har68a, Har68c, Mac62b, Vas69]. conservation [Dav56]. considerations [BN69, CB55, Eng61]. consistency [Coo69, Hur69, RS64, Sav65, ZD69]. constitution [Ano52b, Ano62a]. constraints [Bra66]. construction [PBK54]. consultants [Ano52l]. consumer [Bau52, Bra52, Wil52]. containing [Bow54]. Contemporary [Ull53, Jor59]. content [RRS61a]. contents [Bra51]. context [Kra64, Luh60, RB68]. continuing [Hut55, Kin59]. contol [LH69]. contractor [KB68]. Contributing [Dar66]. Contribution [Emb65, Hof67]. Contributions [Bro53a, Bro53b, Cua64]. Control [Cla67, BKB62, BML64, Bis59, FL69, KMW66, KB68, Lip63, Rop68, SB59, Sti58, Wal69, War56, Wei66]. controls [Cla52, Tau52a]. Convention [Oat68, Ano62b, ML67]. conventional [Lyn69]. conversion [Ham68, Lip63]. Cooperation [Mat62, Boh62, Hys53]. cooperative [Tau52a, Met51]. Coordinate [Bro59, Blo57, Boo60, Cos62, Gul56a, Her57, Hol66, Mon65, Per68, Pla60a, Sla63, TGW52, TW53, Tau61b, Wad58, WD59, Bra67]. coordinated [Bey60a, HH60]. Coordination [WBO^{+}61, Cos61c, Jen65, Mon62]. copy [Tru65]. copying [Sei56]. copyright [Hei68, Laz68, Sop68]. corner [Ano50b, Ano50c, Ano51b, Ano51c, Boo50, Moo50]. Corporate [Bra64]. Corporation [New69a, Mor60]. Correction [Ano64e]. correlates [GB69]. correlating [Woo57]. Correlation [O’C64b, CBF59, Rid51b, SD55]. Correlative [Ber56a, Ber57a, BH57b, Ber57b, Ber58, Ber60, Ber57a, Ber57b]. Cost [BF64, MS66, MS67, Cha50, Hel69, Hir62, Lan69b, Mon65, Ohl57, Rid51a]. Costello [Bra67]. costs [Ano54f, Lan62, Lip63, Ove63]. Council [Ano59m, Dar66, Den50, JT61a]. counterparts [Kot62]. counts [Gro68]. coupling [Kes63, Kes65]. Cox [Har68c]. Cranfield [LM64, Ree65, Ric63b]. Craven [Har63]. creativity [Mai60]. Crimes [Kes59]. crises [TEC64]. crisis [BH63]. Criteria [BC63, KBLP55, Slo68]. Critical [Hay67, Kle64]. Criticisms [Lil54]. Cros [Vai67a]. Cross [Ano64d]. Cross-References [Ano64d]. Cuadra [Ano68a, Bat67a, Try69]. Culbertson [Cha63]. cumulated [Tru67]. cumulation [Tru67]. cumulative [Gul51, KT67]. Current [Fel57, Tru67, Ano58i, Har68b, MK62, Sag66]. curricula [Art69]. curricular [Sla69]. curriculum [Sto69]. curtain [SS59]. cutter [AH63]. Czechoslovak [Nem57].

D [Cle63, Har68a, Lan65a, Ste69]. D.C [Ano68h, Ano68f, Art65, Ash66, Boh62, Har67c, Har67b, JT61a, Lan67, Lit64, New69a, Tru67]. D.C. [Cla55, Boh62, Wel61]. Dagobert [Ste69]. Dagron [Lut50]. Data [Emb65, Bec68, BS69, Cha63, Coh59, Cor54, Die62, Ega53, Fla69, Gro67, Ham68, Heu56a, KP56, Ken58, KC69, Mal58, San68, Sar68b, Sav67, Sha56, SM63, TBW68, Ull53, Wac52, Woo57, FC68, Hay67, Hin69]. Data-Retrieval [Hay67]. David [Ang58]. Davison [DG57]. day [Ano68g]. dead [Cam56]. December [Ano68h, Ano68f, Wad68]. decentralized [EP68, WW67]. decimal [Har54, Str56]. defense [Lan62]. deferred [Ano52e]. defining [Ros65a]. definition [PBK54, Ric60, KPB54]. Definitions [Ano55d]. Delegation [Fai58]. demonstration [Ano55r]. Denison [Ano62d]. dental [Lyn69]. Denver [Ano64d]. Department [Ano62c, Ano64d, Wel61, Lan62]. dependency [SKM64]. depth [Man61, SM63]. descriptor [RB68]. Descriptors [Ano64d, Har68b, Jas62, Lib67, Ano68f]. design [Con52, Del64, Hil64b, PKB55b, PM59, Wil69]. designing [Cor54, KBLP55]. Desmond [Gre67]. Detailed [Man64, Woo57]. determination [Sop68, Tru65]. Determining [Tob59a, Jah64]. detour [Cam56]. Deutsch [Ano65a]. Deutschen [Ste69]. Development [HR61, Ano58i, BF58, Bou61, Ega51, EP68, HH67, Jon60, KPB54, KP57, Met51, PKB55b, Sto69, Yov69, Ano58i, Ano64d, Ken60]. Developments [Dav51b, Lee52, Meh57, Tya61]. device [Per65]. devices [LM64, War66]. Devil [Van63]. diagrs [Ste69]. diary [RH63]. diazotype [Ben57]. Dictating [BL55]. Dictionary [LBL67, Tre66, TH69, SSS61, Van63, Ven68, Wag60, Wei66]. Different [Wåh66]. digital [OB59, Sha56]. dimensional [Hof52]. dimensions [LH69]. Direct [Sch64b, Fie56]. directions [McG58b]. Director [Ano62h]. Directory [Ash66]. disagreements [O’C67, O’C69]. disbursements [Nee62]. discarding [Pou56]. discipline [Yov69]. disciplines [Eva56]. Discovering [Gro51]. Discussion [She67, Man64]. Display [OB59, Ber60, CM62, Wor69]. displayed [TBW68]. disseminating [RH63, Tar62]. dissemination [Lee52, Luh61, SAF65, Sag66, She53, Sop68, TM62, Tum53]. dissertations [Har68d]. distinction [JT61b]. Distribution [Ano62c, ZS67, HW64, Lee52, MS66, MS67, Mil53]. distribution-acquisition [Mil53]. divided [Ric56]. Division [Ano64b, Ano68h, Gre67, Ano68a, Bat67a, Eas51, Hil66, Web58]. Document [PV63, Uhl68, Bar53, Bra62, DD65, Fei63, JS69, Jam59, KB68, Lan69b, Lec68, Les69, Lib67, Lip63, Man61, Sal63, SSO65, SWCT68]. Documentaires [Vai67a]. documentalists [Ano60a]. documentarea [Sta61]. Documentary [Ano64b, Per66b, Tat50c, KTM57, Pow51b]. Documentation [Ano53a, Ano55b, Ano57a, Ano59m, Ano61a, Ano62a, Ano63a, Ano64a, Ano65j, Ano67g, Bal50c, Bat67a, Bec68, Bey55, Dub62, Emb65, Eva51, Gre67, Gul61, Hay66b, McM61, Nem57, Par62, Pol50, PV67, SG69, Sei56, SE50a, She67, Vai67b, Wep61, WSS54, Ano55d, Ano58i, Ano62n, AC61, Bat67b, Bey57, Bey60a, Bey62, Bis53, Bro53a, Bro53b, Bro55, Bro56, Cla56, Cob56, Eng62, Eva56, GG57, Gra68, Hir62, KI59, Kui51, Kyl63, Lie54, Mac58, Mas56, McG58b, Meh57, Orr59, Pla60a, Roy57, Sch54, Sch64b, Sco51, Sha54, Ste69, Ste55, Tat50b, Tau52b, TMK55, Urq56, Van63, Wag60, WBO^{+}61, Ano62u, FC68, Rey61, Sta61, Tat50b]. Documents [Ano68h, Ano68f, Art65, Tru67, Ano51h, BKRL66, Eas51, Har51, Mas54, McC65, O’C68b, O’C69, Ruh64, Sav66]. Documentum [McG58b]. Dohme [KMW66, SS61]. dokumentació [Emb65]. Dokumentációs [Emb65]. Dokumentation [Ste69, Ste69]. domestically [Kot62]. Don [Cha63]. Dopkowski [Lit64]. Dougherty [Log67]. Dr [Bra67, Emb65]. DRP [Ano65b]. DRP-CI [Mou65]. drug [Spe67]. dual [Ric63a]. duality [Mor54]. Duplication [Eva56, BHR67]. Dutch [How65]. dynamics [Mas56].

e.V [Ste69]. earliest [Lut51]. Early [Emb65]. easing [She59]. East [Lit66]. ecology [Don57]. economic [Sop68]. Economics [Kun63, HS68]. Ed [Bal69, Met65, New69a, Vas69, How65, Lit64, New63b, Ran67, Sny66, Tau65a]. edge [Hof62]. edge-notched [Hof62]. Edited [Ano69b, Bra67, Eva66, Gre67]. Editing [Coh55, BK63, Sla64, Ven68].

Edition [Ano64d, Ano68h, FC68, Kut64, Lan67, Pow51b, Mac62b]. editions [Ano51h, Car50]. Editor [Ano68d, Ano68b, Ano68a, Bat67a, Boh62, Har67a, Mou67, Try69, Ano68e, Blu59, Ano50b, Ano50c, Ano51b, Ano51c, Ano61b, Ano62g, Ano62i, Ano69g, Ath65a, Bea69, Boo50, Bre64, CLA^{+}63a, Cla55, Eva61a, Eva61d, Eva61b, Eva61c, Ken62b, LL63, Mac62a, Mai61, MOP^{+}63, MBG^{+}66, McG60, Moo50, OHBS66, Pet59, Sch64a, Sha61, Stu69, WFY^{+}63]. Editorial [Ano52c, Ano52d, Ano53d, Ano53e, Ano53f, Ano53g, Ano54b, Ano54c, Ano54d, Ano54e, Ano55e, Ano55f, Ano55g, Ano55h, Ano60a, Ano60c, Ano60b, Ano60d, Ano62e, Ano62f, Ano63b, Ano65d, Cob65, Dar66, EK64, Eli64a, Eli64b, Eli65b, Eli65a, Eli66a, Eli66b, Eli67b, Eli67a, Eli68a, Eli68b, Eli69b, Eli69c, Eli69a, Ell67, Eva61a, Kyl63, Mac62a, Mac63a, Mac63b, Sch69b, Tre66, Wel64, Wel65, Wel66]. Editors [Anz67, Har68c, Art61, Ber69, Blu50, CDM65, Eva61d, FP66, FFBB67, Gul67, HHK68, HSM^{+}68, KTW64, LL68, New69b, ODVB68, RVS64, SVS^{+}67, Sal65a, SBF65, SB66, SFMO67, VMB69, Wai69]. Edmund [Ano68d]. Eds [Hay66a]. Education [Ano62u, Ano68f, Hay66a, Hay69, Mou67, Wel61, Ano69a, Bon62, Che65, DO69, HG65, Swa65, Tag65, Har67b, Pre69, FC68]. Educational [FC68, Sto69, Ric68, Sop68]. Edward [Cha63]. Effect [She52, Tru65]. Effective [Tri64]. Effectiveness [Swe69, AW69, Bor61, Coo68]. efficiency [Lan69a, ZS67]. Electric [Ano62c, Lev64]. electrical [Mil57]. Electronic [KP56, Kin60, PKM59, Bec68, Ken58, Luh61]. electronically [O’C63]. Electronics [Eng62, Don57, Mil57]. Elements [Boh64]. elimination [TEC64]. Elizabeth [Art65]. Elmhurst [Whi66]. Elsevier [Ano69f, How65]. Emergence [She53]. emerging [Hod68]. Encyclopaedia [Try69]. end [Cam56, Lec68]. ended [Fla69]. endocrine [OC62b]. Engineering [Lan65a, Mou67, BF58, Bur59a, Bur59b, Bur59c, Hei61a, Phe58, Pin67, Ash66]. engineers [And64, JHG66, Lev64]. Englewood [Ano68d]. English [How65, Tre66, Ano68g, Bes61, Sav63, SM63, SKM64]. enlargement [Hol51]. enneametry [San68]. enriching [Bor55]. entry [Fie56]. Equations [Ike67]. equipment [Hol62, Luh61, MH53, PKB55a]. ERIC [Ano68f]. Errata [Ano55i, Ano56a, Ano65e]. Erratum [Ano66b]. ESI [BK63]. esophagus [Ano60d]. essay [McG58b, Que63]. Essays [Gar65a]. estimate [Bou62]. Europe [Bet59, Bis59]. European [Bor54]. Eva [Ran67]. evades [Car63]. evaluating [JH66, Mac53, Ove62, RH63]. Evaluation [Blo57, Cot68, Fel63, Lil54, Ran67, War56, Ano69d, Bar53, Ber65, BF64, Del64, GS63, Lan65a, Lan69a, Mon65, Sal65b, Sav63, Tau62, Ano65b, Ano69e]. Even [Har68b]. Even-numbered [Har68b]. events [Har68b]. example [Tau62]. exchange [GG64, Mac55]. Executive [Ano62h]. Exeter [She67]. existence [Wad68]. Existing [Cam56]. Expected [Coo68]. Experience [Kor64, Dav51a, Gra52b, Hys61]. Experiences [Sha56, Hof62]. experiment [Bor61, Dam65, Ham68, Hys61, Kna65]. Experimental [Mou65, Bor67, Eva56, Lip62, Ric67]. Experiments [Sey66, Sny66, DD65, Lut51, RK58]. Exploitation [Ken60, Ove62].

Exploiting [Bau52]. extensive [Bor50]. Extracting [VZ68]. extreme [Bal69]. eye [Mil68b].

F [Ano69e, BF57, Gre67, Lan65b, Sha65, Sny66, Tre66, Vai67a, Vai67b, Whi66]. FAC [Mit53]. FAC-tronic [Mit53]. Faceted [Per66a]. facilitate [BKRL66]. Facsimile [Cro56, Ano69d, TH69]. fact [Poh65]. Factors [Can52, Ros65a, GS63, PKB55b]. facts [Lip65a, Sop68]. Faculty [Lan65a]. Fairthorne [Lin66, Ano60m]. Farewell [EM64]. fast [KB68]. fast-turnaround [KB68]. fax [Con51a]. February [Ano69d, Ano62a]. Federal [Ros65a]. felismerések [Emb65]. female [HH60]. Festschrift [Eva66]. fiction [Poh65]. Fid [Bry55, HMC63, Sha54]. field [Bou62, Mit53, Ste69, Wil52]. fields [HS68, Mil57]. figure [Dyk59]. figure-of-merit [Dyk59]. figuring [Car56]. file [Lei68, Tag65, WD59]. files [Dai56, Gil58, New61, Sho62]. filing [BF58]. Film [Sil50a, BHR67, Fis50]. films [KC69, Ano53c]. filmsort [McM61]. final [Ree65]. finance [Ega53]. findings [Emb65, GG64]. first [PKM59, Tru67]. Fisher [Ran67]. Flexowriter [BKB62]. flow [Ste69]. footage [Car56]. footnote [Kap65]. Force [Ano64b, Lan65a]. Foreign [Mar60, Har63, Orn63, Tru67]. Foresight [MM59]. Foreword [Tat51]. form [Che63]. formal [KP57]. formation [RRS61b]. Former [Ano62g, Tat50c]. forms [O’C67, Rid51b]. formulas [OB59]. Fort [New63b]. Forte [Ano68c]. Forward [She67, AH63]. found [Kna65]. Foundation [Bra67, Cla67, Ste55]. Foundations [DO69, Min60]. four [RRS61a]. frames [Kue66]. framework [JC67, Jor68]. Francis [Ano68g]. Frankfurt [Ste69]. Fred [Log67]. Frederick [Dai64, War66]. freeway [Cam56]. Fremont [Mac62b]. French [How65]. frequency [CR69, Gro68, LBL67]. fringe [OE58]. Fry [Web65]. FS [Ano68f, FS57]. function [Fuk67]. functions [LKPB55, PKB55a, Ros65b]. fundamental [Min57]. Fundamentals [Ano65a, Cor54]. further [Ric61]. Future [Boh65, Ano62x, Dic66, Hei68, Per53, Vic60].

G [Bal69, Tau61a]. GAF [SWCT68]. game [Ken64, Tre68]. Gardin [Met65, Vai67a, Vai67b]. Garvin [Tau65a]. gathering [Mai60]. Gauthier [Vai67a, Vai67b]. Gauthier-Villars [Vai67a, Vai67b]. General [Ano62c, Wåh66, Vai67a, Gil58, ML67, PKB54, SB59, Uhl69, WD59, Ano62c]. generalized [Lib69, SS62, Sch63]. generated [AW69, Sla69]. generation [BKB62, BML64, SR68]. generator [Eng61]. generic [Tau62]. genesis [Mil52]. genitalia [HH60]. George [Ano64b, Web65]. Geosciences [Hil66]. German [How65]. Germany [Str56]. Gesellschaft [Ste69]. Gil [Tre66]. given [Ano69d]. glossary [Ste69]. goals [Moo60]. gödel [Buc61]. Goldhor [Pra63]. Goldwyn [Hay66a]. good [PV63]. Gordon [Ano69a, DG57]. Government [Ano68f, Art65, Ash66, Boh62, Lit64, Lit66, Rey61, Can52, Ros65a]. Goy [Tre66]. Graduate [Bra67, Lew66, Met65, Per66a]. Grafton [Ano65a]. grail [Car63]. graphic [Kuv67, Lew69]. graphics [Kuv67, McC68]. graphs [Coh55]. Green [Lan65b]. Grolier [Vai67b]. Grose [Har68c]. Ground [KP57]. group [Ber57b]. grouping [Per68]. groups [DG57, HH53]. growing [Sch62b]. Growth [HS68, Koc69]. Guest [Cob65, Eli69a, Kyl63, Sch69b]. guide [Har68b, SPS68, Ste69, Ano62d]. guideposts [Sop68]. Guides [Hil66].

H [Pre69]. H. [Ano69b]. Habitability [Wat68]. half [BK60]. half-life [BK60]. Hall [Ano68d]. Halo [Bec63]. Hamden [Har68a, Har68c, Vas69]. Hampshire [Bro59]. hand [Per51]. hand-sorted [Per51]. Handier [Ano67a]. handling [Boh62, BH67, Har68c, Per53, Que63, Sha62, Ano64c]. Hans [Rog67]. Harman [Sny66]. Harold [Ano68e]. Harper [Pra63]. Harvard [Tau61a]. Hawkin [La 63]. Hayes [Boh64]. Haykin [Ang58]. heading [Ano62d, Kra64, Way68]. headings [Dai57, O’C64b]. Health [Dar66, Tru67, Wel61]. Hebrew [Hay67]. Heiliger [Cha63]. Heinritz [Log67]. held [Mou67, Swa65]. Henry [Ano68g]. Her [Pre69]. Herbert [Hay66b, How65, Pra63]. Herman [New63a, New63b, BS69]. Herstellung [Ste69]. heuristic [Ken64]. Hierarchical [Ric60, Sal63]. hierarchy [Per66b, Ath65b]. High [Sha50, Ano60b, MK62, Pow50]. Higman [Ano69f]. Hill [Ano68b, Hin69, Tau65a]. Hillel [Gar65a, Man64, Moo57]. hindsight [MM59]. historian [Jor59]. historical [Bou61, Dav56, Gar63, PM59, She59]. History [Emb65, SG69, Jon60, Kin59, TEC64]. hobgoblin [Coo69]. holder [Sha53]. Holdings [Wad68]. Hollerith [BS69]. holy [Car63]. horizons [Ell51]. hospitality [Per65]. Hotel [Ano62h]. Houghton [Vas69]. House [Eva66]. housing [Her52]. Hrsg [Ste69]. human [RS64, Ruh64]. humanistic [Ano68d]. humans [Kna65]. Hungarian [Bey60a, Bey62]. Hungary [Bey55].

I.R [Jon63]. I.R. [Jon63]. Ibusmae [Bea69]. ideas [Ass55, Tat50b]. Identification [Ath64, Bow54, Sti58, Bor67, McC65, War52]. Identifying [Cua64]. II [Eva66, BML64, Ber57a, Bur59b, KBP54a, MS67, Met65, Res61, Sla69, Tin68]. III [BH57b, Bur59c, Her57, KBP54b]. Ill [Whi66]. Illinois [Boh65, Har67a, SB59]. illiteracy [Woo64]. illus [Ste69]. Illus. [Sta61]. illustrated [SB59]. im [Ste69]. images [TM62]. Impact [Ano65b]. implication [Eva56]. Implications [Tau53, Bak69, HS68, Rau67, Tum53]. importance [Gro68]. Imprecision [Tau68, Tin66, Tin68]. impressions [Urq51]. Improvement [Lip65b, She50, Fel64]. Improvements [Vor65]. inadequacy [Man61]. Inaugural [Sch62a]. including [Fla69]. inclusion [Lip65b]. incoming [Sch62a]. inconsistency [Tin66]. incorporation [Ano52b]. increase [Tru67]. independent [O’C69]. Index [Ano65b, Del64, AL68, Ano62q, Ano65a, ABC64, Bro68, Fis66, Hin69, Hut55, Jah64, JS69, KT67, Kuv67, LM64, Lew69, Lip62, Luh60, O’C62a, Ort61, Pla60a, Pot68, Sav65, SPS68, SS62, Sha66, Van65, Vas69, Vor65, Wac52, Wel61, Dar66, Spe67]. Index-abstract [Del64]. indexed [Har68b]. indexer [Ben69, Hur69]. indexer/cataloger [Ben69]. indexes [Ada55, Ano52l, Ber56a, Ber56b, Ber57a, BH57b, Ber57b, Ber58, Ber60, Cos62, Gar55, Gar63, HW64, JHG66, KBP54b, Kna65, KT67, LM64, Lip65b, Mar62, Per68, Sla64]. Indexing [Ano64d, Art65, Ber65, Cla67, Doy62, Lou66, Mor60, SKM64, War66, ZD69, Ano68h, Art64, AW69, BL55, Blo57, Boo60, Bou62, Bro59, CHJ61, Cos61b, Dam65, EFB57, Fie56, GS63, Gul56a, Har68b, Her57, Hol66, KBP54a, KP57, Kes65, Kra64, LD60, Lip60, Lit64, MW61, Man61, Mil57, Mon62, Mon65, MS62, O’C61, O’C64b, PM59, Ree62, Ree65, RB68, Ruh64, Sal69, SS61, SSO65, SM63, Sla63, TGW52, TW53, Tau61b, Tin66, Tin68, Tri64, Wad58, WD59, Wha61, ZS67, Bra67, Gre67, War66]. indexing-abstracting [KP57]. India [Meh57]. Indian [Par62]. Indianapolis [Ano69c]. Indication [Lew66]. indicators [Lan68, Lip65b, RRS61a]. indices [Buc61]. individual [Bis57, Daw68, Eng61, Ano61a]. Industrial [HH53, New63b, Fag52, Fra52, Mai60]. industry [Ano53b, Ega53, Gra52a, Hay60, Lar52, McM61, Str57, Swa68]. inflation [Lon68]. informal [Gar65b]. Information [Ada56, Ano64c, Ano65a, Ano68e, Ano68h, Ano68f, Ano68p, Ano68q, Ano69b, Ano69e, Ash66, Ath65b, Bat67a, Boh64, Bor68, Bra67, Che65, Cla61, FC68, Fuk67, Har63, Har67b, Hay66a, Hay60, Hil66, Hof62, Kle69, Lew66, Lin66, Mai60, Met65, Mor62, Mou67, Nol59, Per66a, Per50, Poh65, Pol50, Rog67, Rub66, Sch69a, Sny66, Swa68, Vic59a, Whi66, Yov69, And64, Ano50a, Ano50d, Ano51d, Ano51e, Ano51f, Ano51g, Ano52h, Ano52i, Ano52j, Ano52k, Ano53b, Ano59m, Ano62c, Ano68a, Art69, Ass55, BN68, Bak69, Bal50a, Bal50b, BH57a, BH63, Bar56, BF58, Bec63, Ben69, Boh55, Boh62, Bon62, Bou61, BF64, BS68, Bro52, BH67, Buc61, Bui62, Car68, Cla63b, CM62, Cle63, Cos61a, Cot68, Cua64, Cue68, Dav58, DO69, Don66, Ebe66]. information [EP68, Eng61, Eva68, FCHB65, Fel63, Fla69, Gar54, GG64, Gar65a, Giu69, GVB64, Har68b, Hay69, Hei61b, HG65, Hel69, Her62, Hil69, Hod68, Hof67, HA65, Hoy69, Hys63, Jam59, JH66, JT61a, Joh62, JN58, Ken62a, KBLP55, Ken60, Ken64, Kin62, Kle64, KMW66, KTM57, Lan65a, Les60, Lev64, Lib69, Low68, LH69, Luh53, Luh61, Lun69, Lyn69, Mac53, Mac55, Mac63b, Mal62b, Man61, McC68, McC62, Mel68, Mil68a, Moo60, ML67, New61, New69a, O’C68a, Oat68, Oet68, Orr61, Par66, Pel69, Pol68, Pre69, Que63, RH63, Ric67, Ros65a, SAF65, SSS61, Sch62b, Sch63, Sch69b, Sch64b, Seb53, Sha62, She53, SD55, Shi63, Sim62, SL68, Sop68, Ste61, Swa65, Swa69, Swe69, SS59, TN68, Tar62, Tau68, Tho68, TBW68, TEC64, Tri64, Try69]. information [Tum53, Ven68, Vic60, Wal67, Wal68, WBO^{+}61, Wil52, WW67, ZS67, Ano62-27, Ano63a, Ano64b, Ano65j, Ano69a, Ano69c, Ano69e, Anz67, DK66, Kik67, Lew66, Mou67, Ano68h, Boh62, Boh64]. Information-retrieval [Mor62, Fel63, Sch64b]. information-searching [Tri64]. information-transfer [Lib69]. Informatsii [Rad69]. Informatsiya [Rad69]. infract [Lip65a]. inquiry [Hod68]. INSDOC [Par62]. Installing [Ken58]. Institut [Rad69]. Institute [Ano63a, Ano65b, Ath65b, Bat67a, Boh62, Gul61, Lan65b, Lan65a, Mou65, New63b, Ano53a, Ano53c, Ano58i, Ano61a, Ano62a, Ano64a, Ano65j, Ano67g, Bal50c, Dub62, Eva51, McM61, SG69, WBO^{+}61]. institutional [Cha50, She53]. instructions [BC63]. instrumentation [WSS54]. insurance [Cue68]. integrated [Bis53, Orr59]. Integrating [Dav68]. integration [Ben52]. Intellectual [Bra67, Lew66, Met65, Per66a, Rog67]. interaction [Oat68]. interdisciplinary [Blu59]. interest [MK62, Ove62]. Interim [Ano62h, Hys53]. interindexer [Coo69]. Interior [Ano64d]. interlibrary [Kin60]. intermittent [Sha50]. International [Ber54, Bro56, Hys53, Mac55, Bro52, Dub62, JT61a, Mac62b, Mos56, SE50b, Ano62y, Lit66]. interpretation [Sav67]. Interscience [Ano68a, Bat67a, Hil66, Rub66]. intersections [Uhl68]. interview [RH63]. INTREX [Sny66]. Introducere [Sta61]. Introducing [Tat50b]. introduction [Ano68c, Har68a, Sta61]. Introductory [Jah69]. inventories [Tob59a]. inverted [Cos62, Jon60]. investigation [SK63]. investigations [O’C61]. involving [HH53]. IR [Sny66]. IRIA [Wol59]. iron [SS59]. Israel [Gar65a, Hay67]. issue [Mou65]. Issued [Ash66]. Italian [How65]. Iterative [Wil69]. IV [Ber57b, BPK54]. Ivan [Emb65]. IX [PKB55a].

J [Cle63, FC68, Har67c, Hay66a, Hin69, Log67, Met65, Per66a, Sny66, Vai67a, Vai67b]. Jan [Tru67]. Janaske [Boh62]. Janda [Ano69c]. January [Ath65b, Nee62]. Japan [Ano59m, Fuk67, Kik67]. Japanese [Kot62]. Jean [How65]. Jefferson [Lan65a]. Jersey [Ano68d, Bra67, Lew66, Met65]. Jerusalem [Gar65a, Hay67]. Jessica [O’C63]. JICST [Fuk67]. John [Ano64c, Ano65a, Ano68a, Ano68e, Ano69e, Bat67a, Boh64, Bra67, Har67a, Hil66, Lew66]. joint [Gre67]. Jolley [Hin69]. Jonker [War66]. Joseph [Boh64, Pra63]. journal [Bis56, Bou62, Kun63, Lei69, Tat50c]. journals [Ben52, Bra51, Bur59a, Bur59b, Bur59c, Car69, NT58, Ric63a, Tob59b]. Joyce [Sny66]. Jr [Bra67, Lan65b]. judgments [GB69, RS64]. Judson [Ang58]. July [Ano61a, Swa65].

Kaplan [Hil66]. Kasher [Hay67]. Kaula [Eva66]. Kee [Ano69a]. Keenan [Gro67]. Kendall [DG57]. Kenneth [Ano69c, Har63]. Kent [Rub66]. Key [Luh60, Cua64, Kna65, RB68, TM62]. key-word-in-context [RB68]. keypunching [Ham68]. keysort [Mor60]. keyword [JS69, Kra64]. keyword-in-context [Kra64]. Keywords [Ano64d, Ada67, CR69]. kezdeteihez [Emb65]. Kimber [Har69]. Kinds [Wåh66]. Kingdom [Urq56]. Klassifikationssysteme [Ste69]. Klassifikationssystemen [Ste69]. knowledge [All59, Ano55c, Dye56, Hut55, Koc69, Moo51, San68, She52]. Kodak [TMK55]. Könyvtar [Emb65]. Központ [Emb65]. Kucera [Ano68g]. KWIC [Kra64, AL68, Fis66, Luh60, Ruh64, Van65]. KWIC-machine [Ruh64].

L [Hin69, Lit64, Pra63, Pre69, Tau65a]. Labor [Lip63]. Laboratories [Ano64b, Sha59, Dav58, KMW66, SS61]. Laboratory [Ano64b, Gre67, Lan65b, Ano57n, Mai60]. Lancaster [Ano69e]. Landrum [Dai64]. Language [Ber56b, Har63, Oet68, Tau65a, Alt67, Ano68e, Bou62, Cho57, GVB64, LM64, PKB55b, SM63, Tau61a, Gar65a, Ano68c, Har63]. languages [LB56, Ano68b, Ano69f]. large [Aro52, BF58, BF64, Bra62, Nol59, Rop68, Sha59, Sho62]. large-scale [BF58]. laundry [Stu69]. L’Automatisation [Vai67a]. Law [O’C63, Mos56, Nem57, Ano60c, Gro67, Sav65]. laws [Ano52b, Man66]. Learned [Den50, Tob59b]. legal [Dye56, O’C63, Sop68]. Lehigh [Mou67, Ano60b]. length [Coo68]. Leonard [Har63]. Lester [Gre67]. Letter [Art61, Ath65a, Ber69, Bre64, Cla55, McG60, Sch64a, WFY^{+}63]. Letters [Ano61b, Ano62i, Ano69g, Bea69, CLA^{+}63a, CDM65, FP66, FFBB67, Gul67, HHK68, HSM^{+}68, Ken62b, KTW64, LL68, LL63, Mai61, MOP^{+}63, MBG^{+}66, New69b, OHBS66, ODVB68, Pet59, RVS64, SVS^{+}67, Sal65a, SBF65, SB66, Sha61, SFMO67, Stu69, VMB69, Wai69, Blu50]. levels [Cla52, Oat68]. Levéry [Vai67b]. Lévy [Vai67a]. lexical [RRS61a]. Lexicon [How65]. Lexington [Lan65b]. liberal [Che65]. librarian [Hea52, Hoo52, Laz68, Wil51a]. Librarians [Dar66]. Librarianship [Don66, Tre66, Ano58i, Ber54, Giu69, Har63, Tau52b]. Libraries [AS67, Ano62d, Cla56, Gre67, Hys53, Ran67, SS59, Wad68, Bet59, Cla61, Dav67, Mor55, Pow50, Slo68, Har69, Pra63]. Library [Ano66a, Ano69d, Ash66, Bey60a, Bra67, Don57, Emb65, Eva66, FL69, Gre67, Har66, Hin69, Kil66, La 63, Lew66, Log67, Meh57, Met65, Nit67, Per66a, Pie64, Sha59, Tru67, Wad68, Web65, Wel61, Ano52l, Ano57n, Art69, Bor55, Bro59, BGM56, Car52, Dai64, DSTW66, Fra52, Gul56b, Hol50, Hol51, Kno60, Kor64, KL67, Lan62, LK56, McC61, Mit53, NT58, Phe58, Pin67, Pow58, Pre69, Rop68, She56, War56, Wil51b, Boh65, Cla55, Den50, Eva66, Gul51, Kil61, McG58a, Ran67, Cha63]. library-school [Art69]. Lierature [Ano64d]. Lieutenant [Lan65a]. life [BK60, Cue68]. Limited [Ano62c, MW61]. Lincoln [Lan65b]. line [Ben69, DSTW66, Jon63, McC68, Wor69]. linguistics [Gar65b]. linkage [Vic60]. Linking [Sar68a]. links [AH63, AW69, Cos61a, Lan65a, Mon62, Mon65, Tau61b]. Linofilm [TJ69]. Lipetz [Ano65b, Mou65]. List [Ano64d, Ano54n, Ano61a, Hir50, Pou56, Way68, Ath65b]. Listing [Wad68]. lists [Ano62d]. Literature [Ano54g, Ano54h, Ano54i, Ano55j, Ano55k, Ano55l, Ano55m, Ano56b, Ano56c, Ano56d, Ano56e, Ano57c, Ano57d, Ano57b, Ano57e, Ano58b, Ano58a, Ano58c, Ano58d, Ano59a, Ano59b, Ano59c, Ano59d, Ano60e, Ano60f, Ano60g, Ano60h, Ano61c, Ano61d, Ano61e, Ano61f, Ano62j, Ano62k, Ano62l, Ano62m, Ano63e, Ano63c, Ano63d, Ano63f, Ano64f, Bou64a, Bou64b, Bou64c, Bou65, BL65, Bro53c, Bro53d, Bro54, Cla67, Lam65a, Lam65b, Ada55, Ano52l, AC61, Aue55, Bau52, Ben52, BPK54, Bin63, Bor56, Bou62, Bra64, BBJB69, Bro55, Car69, Fos58, GS63, Hof52, HS68, Hys53, Hys61, KBP54a, KBP54b, KPB54, KBLP55, Ken60, Kyl58, Lan69b, Lei68, Lei69, Lip62, Lip65b, LKPB55, Luh60, Lyn69, MH53, Man64, Man66, Mar60, Mil52, Mil68b, O’C63, OE58, Orn63]. literature [Ort61, Ove62, PKB54, PBK54, PKB55a, PKB55b, PM59, Ree65, Rot56, Sag66, SPS68, Str57, VZ68, WD59, WP62, Smi55, Mou65]. literatures [BK60]. loan [Ano62d]. loans [Kin60]. Locate [FC68]. logic [Bra66, SKM64]. Logical [Cho57, TW53, TPBK55, Uhl69]. logically [All59]. logician [Ano57a, BH57a]. London [Ano65a, Gar65a, Hay66b, Lin66, Pre69, She67, Ano54q]. looking [Eli69c, She67]. Looseleaf [Ath65b]. L’Organization [Vai67b]. Lou [Ran67]. Louis [Cha63, Har66]. Low [Cha50, Ohl57]. Low-cost [Cha50, Ohl57]. Ltd [Gar65a]. Luhn [Ano69b]. Lydia [Ano60i].

M [Boh64, Cha63, Har68c, Hay66a, Log67, New63a, Ste69, Sha65]. M.I [NT58]. M.I.T. [BGM56]. MacDonald [Ano69f]. Machine [BPK54, Die62, Hys61, KBP54a, KBP54b, KPB54, KBLP55, LB56, LKPB55, Mac60a, PKB54, PBK54, PKB55a, PKB55b, Sla64, Ano65c, Bal55, CHJ61, EFB57, Gar54, Ken60, Kor64, MH53, Per50, PK56, PM59, RSF^{+}67, Sha62, She52, LKPB55, PKB55b, WD59, Ruh64]. Machinelike [MS62]. Machines [Ano55a, GG57, AS67, BL55, Bor64, Bra68, CG62, JT61b, McC61, O’C61, Ohl57, Per53, RRS61b, San68, Sti58]. magnetic [Jam59, Ric68]. maintained [JHG66, WS64]. Maintenance [Orn52, Wei66]. Majesty [Pre69]. Making [KBP54b]. malignant [HH60]. man [Bra68, Kor64, Oat68]. Management [Log67, Mou67, BF58, KB68, She51]. Managerial [Hel69]. mandate [Ano60a]. manipulation [Opl59, Wha61]. manipulative [HW64]. manual [Sal69, Sha54]. Manuscript [McG58a]. manuscripts [Den50]. March [Pin67, SD55]. marginal [Ohl57, Seb53]. Marilyn [Har67b]. Marlis [Vic60]. Marshall [Ano62c]. Mary [Art65, FC68]. Mason [Anz67]. Mass [Lan65b, Sey66, Sny66]. Massa [Tre66]. Massachusetts [Gar65a, Lan65b, Tau61a]. Masthead [Ano54j, Ano54k, Ano54l, Ano54m, Ano55n, Ano55o, Ano55p, Ano55q, Ano56f, Ano56g, Ano56h, Ano56i, Ano57f, Ano57g, Ano57h, Ano57i, Ano59e, Ano59f, Ano59g, Ano59h, Ano63g, Ano63h, Ano63i, Ano63j, Ano64g, Ano64h, Ano64i, Ano64j, Ano65f, Ano65g, Ano65h, Ano65i, Ano66c, Ano66d, Ano66e, Ano66f, Ano67c, Ano67d, Ano67e, Ano67f, Ano68j, Ano68k, Ano68l, Ano68m, Ano69h, Ano69i, Ano69j, Ano69k]. Matching [Tag65, Uhl69]. material [BKB62, Bow54, Dav56, Gro51, Ric67]. materials [Gul56b, Pin67, Rid50, Sil50b, Sil51, War56, Ano69d]. Mathematical [Sti60, Min60]. mathematics [Tau65b]. matrix [JH66, Yag65]. matter [Per53]. Maximum [SM63, ZS67]. Maximum-depth [SM63]. Maxterm [Fei63]. Maxterm-miniterm [Fei63]. May [Gre67, Har63, Har67b, Mou67, Bin63, Pin67]. McGraw [Ano68b, Hin69, Tau65a]. McGraw-Hill [Ano68b, Hin69, Tau65a]. meaning [McG60, Tin66]. means [Ass55, Bau52, Bra52]. measure [Coo68, Tru65, Uhl68]. measured [Tin66]. Measures [Pol68, JC67]. Measuring [Bor64]. Mechanical [Hol51, Hol66, AW69, BH51, Gar54, Moo51, Vic51, Bur59b]. mechanised [Pla60a]. mechanism [CB55]. mechanization [AR69, Kor64, Man64, Wel61]. Mechanized [Hay66b, RK58, Rub66, Bou61, O’C61, Uhl69]. Media [Mou67]. Medical [Ano64b, Kil61, Ano52l, Flo67, Kin62, LBL67, MW61, O’C64b, Wad68, Wor69]. Medicine [Har66, Tru67, Wel61]. Medicus [Spe67]. MEDLARS [Lan69a]. meet [AL68]. Meeting [Ano62h, Ano52e, Ano52f, Ano53a, Ano54q, Cua65, Eva51, McM61, Tat54]. meetings [Ano52g, Bau66, Bro56]. meets [Bry55]. Melton [O’C63]. members [Wad68]. Membership [Ano54n, Ano61a, Ano62z]. memory [Che63, Nol59]. men [Bor64, RRS61b]. Mental [Dar66]. Merck [KMW66, SS61]. merit [Dyk59]. Merrill [Ano69c, New63a]. message [Hei65, Sop62]. Messages [Hei65]. meta [GVB64]. meta-language [GVB64]. metabolism [Orr61]. Metal [Mor60]. metallurgical [Fos58, Hys53, MH53, Ree65, WP62, Bur59c]. metallurgists [Ove62]. metallurgy [Ken60]. Metals [Hys53]. Metcalfe [Lew66]. method [Alt67, Car56, CM62, Luh53, SS62, Sch63, SL68, Uhl69]. Methodology [Kle64, Sla69]. Methods [Hay66b, Her62, Tar62, War66, Boh55, Con52, Eva68, Hol51, Mac53, Sal69, Sav63, She52, Sil51, Sti60, Swe69, Ano64c]. Metron [Lit66]. Metropolitan [Wad68]. Metuchen [Pel69]. Mexico [Tre66]. Michael [Har68c, Sey66]. micro [Irv51, TM62]. micro-images [TM62]. micro-reproduction [Irv51]. microcard [Bis56]. Microcards [Kui51, Rid51a, Bow54, Rid50]. microdocument [TM62]. microfacsimile [Con51b, Gra52b]. Microfiche [Ric68, Bis61, De 58, Gre67, Kue66]. Microfilm [Pow51b, Pow51a, Tat54, Bor50, Bor51, Car50, Car56, Dic66, Ell51, FS57, KL67, Pow50, Sei56, Tat50a]. microfilming [Har68d, McG58a]. microfilms [Hir50]. microforms [Hei61a]. microphotographed [Ano51h]. microphotographic [Ano50e]. microphotography [Lut51, Mas54, Bal69]. Microprint [Bon51]. microreproduction [Bor56]. microtext [Ano65c, Rid51b]. microthesaurus [TN68]. mid [Kno60]. mid-twentieth [Kno60]. Middletown [Mac62b]. military [Car52]. Minerals [Hil66]. miniature [McM61]. Minicard [Ano55r, KTM57, TMK55]. Minimatrex [Jon63]. minimize [TBW68]. Mining [Hil66]. miniterm [Fei63]. Minneapolis [Anz67]. Minnesota [Anz67, O’C63]. minor [Van63]. Minutes [Ano52e, Ano52f, Ano52g]. mire [Lip65a]. Miscellaneous [Ano54o, Ano54p, Ano62n, Ano67g, Ano68n, Ano68o, Ano68p, Ano68q]. mission [Shi63]. mission-oriented [Shi63]. Mo [Har66]. model [Bat67b, CS62, Ebe66, Lei68, Lib69, ML67]. Modèle [Vai67a]. models [Hil64b, Sal63]. Modern [Har63, And64, Cla53, Dai64, Eva56, Gar65b]. modified [Fel63]. Monograph [Art65]. Monographs [Ano69f]. mooers [Fel63, Ano60c, Wis52]. Mortimer [CG62]. Moskva [Rad69]. Motion [KC69]. movement [Dai64]. Multi [Hof52, Ric54, Vic60]. multi-aspect [Vic60]. Multi-dimensional [Hof52]. multi-plane [Ric54]. Multiple [WP50, Wis52, Alt67, Tru65]. Müszaki [Emb65]. myth [CG62].

N [Cla67, Cle63, Har63, Per66a, Sha65, Wis52]. N.J [Pel69]. N.W [Boh62].

N.Y [Ano69b, Bal69]. NACA [Gra51]. names [Bor67]. NASA [Ano68h, Wai69, Ano68h, BE65]. Nath [Eva66]. National [Ano68h, Art65, Ash66, Bra67, Gre67, HMC63, JT61a, Par62, Tru67, Wel61, Ada56, Bor50, Ebe66, Hea52, Hod68, HA65, Mac63b, Orn63, Sch57, SE50b, WW67, Ano57n, Ano59m, Tat54, Tru67]. Nations [Cob56, Har51]. natural [Alt67, Hof67, Tau65a]. nature [Cla52]. Nauchnaya [Rad69]. Nauchnoi [Rad69]. naval [Blo57]. NBS [Art65]. Need [Cla67, Lan68, O’C68a]. needed [Hay57]. needing [Low68]. needs [Con51b, Vic60]. negative [Heu56a, McG60]. Nelson [Ano68g, Ano69d]. Netherlands [Pol50]. network [EP68, Wal68]. networks [Hod68]. Neumann [Har67a]. News [Ano55s, Ano55t, Ano55u, Ano56j, Ano56k, Ano56l, Ano56m, Ano57k, Ano57j, Ano57l, Ano57m, Ano58h, Ano58e, Ano58f, Ano58g, Ano59i, Ano59j, Ano59k, Ano59l, Ano60j, Ano60k, Ano60l, Ano61g, Ano61h, Cas61, Ano51k, Ano51j, Ano50f, Ano50g, Ano50h, Cla50]. newsletter [Ano51i]. newspaper [Bor67, Slo68]. next [Moo60]. NFSAIS [Jen65]. Nigel [Har68c]. No [Ano64b, Ano68f, Dai64, Dar66, La 63, Lec68, Lit64, Vas69, O’C61]. nonconventional [Wha61]. Norman [Cle63]. norms [Kap65]. North [Nit67]. notation [Fei63, Dai56]. notched [Hof62]. Note [Ano55s, Ano55t, Ano55u, Ano57b, Ano57j, Ano58e, Ano58f, Ano58g, Ano59i, Ano59j, Ano59k, Ano59l, Ano60j, Ano60k, Ano60l, Ano63c, Ano63d, Ano58a, Jas62, Ros65b, Sav63, Tau62, Tau65b]. Notes [Ano54g, Ano54h, Ano54i, Ano55j, Ano55k, Ano55l, Ano55m, Ano56b, Ano56c, Ano56d, Ano56e, Ano56j, Ano56k, Ano56l, Ano56m, Ano57c, Ano57d, Ano57e, Ano57l, Ano57m, Ano58c, Ano58d, Ano59a, Ano59b, Ano59c, Ano59d, Ano60e, Ano60f, Ano60g, Ano60h, Ano61d, Ano61e, Ano61f, Ano61g, Ano61h, Ano62j, Ano62k, Ano62l, Ano62m, Ano63e, Ano63f, Ano64f, Bou64a, Bou64b, Bou64c, Bou65, BL65, Bro54, Fis50, Heu56a, JT61b, Lam65a, Lam65b, Tau61b, Ano50f, Ano50g, Ano50h, Ano51k, Ano51j, Ano57k, Ano58b, Ano58h, Ano61c, Bro53c, Bro53d, Cas61, Cla50, TPBK55, Kin59]. notice [Tat50c]. Notices [Ano52a, Ano52l]. notion [Hil64a]. novel [ABC64]. November [Ano62h]. NS [PV63]. NSF [Ben69]. Nuclear [Gre67, BH67, Cot68, LD60, Wac52]. number [Kue66, Lam61, Tob59a]. numbered [Har68b]. numbers [McC65]. numeric [Fei63, TJ69]. numerical [Per51]. Nurses [Cla67]. Nursing [Cla67, Cla67].

o [Sha65, Pow58]. Oak [Gre67]. objectives [Ano62y, GG64]. Observations [Pin57]. obtained [Har63]. occurrence [LB63]. October [Gul61, Nee62]. Ode [Woo64]. Oettinger [Tau61a]. Office [Ano55a, Ano68h, Ano68f, Art65, Ash66, Mat62, Pre69, Wea55, Web58]. officers [Ano62-28]. Offset [Har68d]. Ohio [Ano64b, Ano69a, Lan65a, New63a, Bin63, Swa65]. old [Bes61]. OMK [Bey60a]. On-line [Ben69, DSTW66, Wor69]. one [JHG66]. Ontario [Ano66a]. open [Fla69]. open-ended [Fla69]. operating [Ebe66, Lan69a]. Operation [Ran67, Bak69, Dav52, JH66, Jon60]. operational [Bau66, Cos61c, Ken60, Ove63, KBLP55, PKB55a]. Operations [Rey61, Bor50, Bor51, Lan62, Ano55a, Log67]. Opinion [Coo69, Lib69, Low68, O’C68a, Sav67]. opportunities [Hoo52]. Optimal [Bak69, Gil58, Tob59a]. optimum [Kue66]. order [Lib67]. ordering [Coo68, Dyk59, Tob59b]. ordnance [Blo57]. organic [WAF^{+}58]. Organization [Ano50d, Ano51d, Ano51e, Ano51f, Ano51g, Ano52h, Ano52i, Ano52j, Ano52k, Ano62-27, Ano63a, Ano65j, Bra67, Lew66, Met65, Per66a, Ran67, Ano50a, Ano51a, Aro52, Bal50a, Bal50b, Fuk67, Gul56b, Kin59, Lei68, LKPB55, Min57, Min60, Moo51, Pin57, She52, Sho62, Tau53, Tho68, War52, Har68c]. Organizational [Bau66]. Organizing [Gra53, Har51, Boh55, Eva68, Har68b, Her62]. Orgnization [Rog67]. oriented [Shi63, SL68, Swa69]. origin [Bou62]. Original [Ano51h, Sil50b]. Origins [Mos64]. Országos [Emb65]. Otdyelnyi [Rad69]. other [HS68, Man61]. our [Hut55]. output [Dyk59, MB68, OB59, Ove63]. Overhage [Sny66].

P [Ano64c, Ano69b, Boh62, Ano62c, Ano68g, CG62, Cha63, La 63, Pra63, Ste69, Pow58]. P.O [Har63]. Pa [Mou67]. Pacific [Ano62o, Ano62p]. page [Eva61a, Eva61d, Mac62a]. Pages [Eva61b, Eva61c, Har68b, Har68c, New63a]. Palo [Gar65a]. Pandex [Ric68, Ric68]. Paper [Coo69, Lib69, Sav67, Ano69c, Ath64, Kle69, Low68, Moo57, ML67]. Papers [Ano67b, Ano68i, O’C68a, She67, Ano50e, Ath64, Kes63, Sla69, Yag65]. paradigm [Bor61]. paramedical [Wad68]. Parameters [Sha62]. Paris [Vai67a, Vai67b, Lut50]. part [BML64, BKB62, BK63, RRS61b, Res61, Sla69, Tin68]. Patent [Ano55a, Mon62, Cos61a, New61, PKM59, Pla60b, Wea55, WD59]. patents [Bal55]. pathology [PW58]. pattern [LD60]. patterns [BMR65, Fai56, Mai60, Sco53]. Patterson [Ano64b, Lan65a]. Paul [Boh62, Tau65a]. PC6 [BML64]. Peek [Ano62q, Uhl64]. Peek-a-boo [Ano62q, Uhl64]. pentathlon [Hil69]. People [Sta61, MS62, Res61]. perceiving [JT61b]. perfect [Tri64]. performance [Fel63]. Pergamon [Har69]. Periodical [Cla67]. periodicals [Wad68, McC61]. permuted [Vor65]. personal [Bor67, JHG66]. Personnel [Har63, Hir62, Mac53, Rau67, Swa65, Hay66a]. pertinency [Ree62]. Peter [Tre66]. Petroleum [Ath65b, WBO^{+}61, Orn54]. pharmaceutical [KMW66, OE58]. pharmacodynamic [Die62]. Philip [Lit64]. photo [Jam59]. photo-magnetic [Jam59]. Photochromic [TM62]. photocomposing [BKB62, BML64]. Photocopying [Laz68, Cha50, La 63]. photoduplication [Hol50]. photofabrication [Bal69]. photography [Bal69]. Photoreproduction [Ost61]. phototypesetter [TJ69]. Photronic [Hue51]. Physical [Ash66, Ath64, CBF59, RSF^{+}67, Pet59]. physician [SL68]. Physics [Ano65b, Mou65, Ano65b]. PICS [KMW66]. picture [KC69]. pilot [Ano69d]. Pings [Cla67]. Pioneer [Ano69b, BS69]. place [Par66]. Plan [Cla67, Bor50, Fel63, Lan62]. plane [Ric54]. Planning [Bor51, Sny66, Sch57, SD55]. plans [Jen65]. plastics [ABC64]. Plus [Har66, Ath65b]. point [Eas51, JS69, Wil51a]. Poland [Wep61]. Policy [Lit66, Mac63b]. Political [Ano69c, Lit66]. Polities [Har67c]. pollution [TN68]. polydimensional [Joh62]. polygraph [Wil51b]. Polzovics [Emb65]. Pont [Cos61a]. Poole [Dai64]. Pop [Coo66]. position [Kle69, Vic60]. Possibilities [Tau52a, Wal68]. possible [Rid51b, Tru65]. postcoordinate [Lan68]. Posting [Gul56a, WD59]. Pp [Tau61a, Ano64b, Ano64c, Ano65a, Ano68e, Ano68f, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Art65, Ath65b, Bal69, Boh65, Dai64, Eva66, FC68, Gre67, Har68a, Har68b, Har69, Hin69, Lit64, New63b, New69a, Pel69, Sta61, Tau65a, Try69, Vas69, Wad68, Ash66]. pp. [Ano62d, Ano64d, Ano68d, Ano68c, Ano68b, Ano68a, Anz67, Bat67a, Bra67, Cla67, Dar66, Emb65, Har67c, Har67b, Hay66a, Hay66b, Hay67, Hil66, How65, Lan65b, Lan67, Lew66, Lin66, Lit66, Log67, Mou67, Per66a, Pre69, Ran67, Rog67, Rub66, Sey66, She67, Sny66, Tre66, Vai67a, Vai67b, War66, Web65, Whi66, Har66, JT61a]. Practical [Gra52b, Ano55c, BS68, Dye56, TM62, Pra63]. practice [PV63]. practices [Bor54, Lev64]. predictions [Moo60]. Preface [Ano53h, Bel69c]. Preliminary [Ano68h, Gar54, Mac62b, SD55, MK62, Orr61, She50, SK63, Alt67]. Prentice [Ano68d]. Prentice-Hall [Ano68d]. Preparation [Spe67, Gar55, Wei66, BKB62]. Prepared [Nit67, Kna65, Way68]. present [Ano68g, BH51, Bou61, Vic60]. present-day [Ano68g]. presented [Lan65a]. President [Ano51i, Gul61, Sch62a]. presidential [Sop62]. Press [Ano68c, Ano68g, Ano69a, Ano69b, Boh65, Cha63, Cle63, Dai64, FC68, Gar65a, Har69, Har67c, Log67, Pel69, Sey66, Sny66, Tau61a, Whi66, CG62, Har67a]. Price [Ano69d, Ano68h, Tru67]. Prime [Lam61]. Prime-number [Lam61]. primer [Ano68c]. principal [Cla56]. principle [Fie56, Hei68]. Principles [PM59, Cos61b, Whi66]. Print [Tob59b, Hol51]. printed [Gar55, Sla64, WS64]. Printing [Ano68f, Art65, Ash66, Fel64, Har68d, Sil51]. Prithvi [Eva66]. probability [Uhl68]. problem [Bis53, Cla53, Hei61b, Jor59, Ken62a, Kin59, Lar52, Mil53, Que63, Roy57, Str57, Urq56, Wea55]. Problems [Bis61, Ham68, Bau66, Cla61, Cla63b, Cos61c, Cos61b, Giu69, Hof67, Irv51, Lou66, Low68, SE50b, Tat50b, Tau68, TMK55, Wal67, KBP54a]. procedure [All59]. procedures [BF64, MW61, Ove62, Ove63, PKM59, Sal65b]. Proceedings [Mou67, New63b, JT61a]. process [Ben57, Ber65, Pow51b, Tay62]. processes [Bar53, Hol51]. processing [Bec68, BS69, Cha63, Hod68, KP56, Ken58, Luh61, McC62, RSF^{+}67, San68, Sha56, SD55, TN68, Tau68, Ano68e]. produced [KL67, WS64]. Producing [Mel68]. production [Bis56, Bis61, Hol51, Kil66, Man66, Ohl57, SS62, Ull53]. productivity [Hin61]. products [TPBK55]. profession [Dye56, Har63]. professional [HS68, Tau53]. Program [Lan65b, Ano62t, Ano69d, Bal50c, BKRL66, Bor55, Cla56, Dav67, DO69, HMC63, Ove62, Web58, HMC63, O’C63]. programmes [Car54]. programming [Ano68c, Ano69f, Ano68b]. programs [Hoy69]. Project [Ano66a, Ath65b, Boh62, Har67b, La 63, Web65, BF58, Den50, Gre67, Pie64, LM64, Ric63b, Wel61]. projects [Bey60b]. Prolegomena [Kap65]. proof [KP57]. proofreading [TJ69]. propaganda [Mac60b]. properties [CBF59]. proposal [Lil54, Ric63a]. proposals [Que63]. proposed [Ano50e, Fel64, New61, TBW68]. prospect [Cla63b]. Prospects [Sey66]. Providence [Ano68g]. providing [O’C68b, O’C69]. Proving [KP57]. pseudo [Tau65b]. pseudo-mathematics [Tau65b]. Psychology [Mil68a, GG64]. psychopharmacologic [OC62b]. Public [Boh62, Har67c, Lit64, Lit66, Tru67, Dav67, Eas51, Pra63]. Publication [BMR65, La 63, Web65, Coh55, Con52, Dav51b, Dav51a, Dav52, Gra52b, Kun63, Lon68, OC62b, Ric63a, Sco53, Sil50b]. Publications [Ano62r, Ano62s, Ano62v, Ano67g, Ano68p, Ano68q, Ano62t, Ano68n, Bey62, Fis59, PV63, SB59, Ano68o]. Published [Try69, Ano52l, Mar62, Sch57, Wad68]. Publishers [Ano68a, Bat67a, Hil66]. Publishing [Ano69f, Eva66, Gar65a, How65, Mar66]. punch [Gil58, MM59, PW58, Per51]. punch-card [MM59]. Punched [Dew59, Heu56b, Coh59, EFB57, Gar54, Gar55, Hof52, HH60, Jon60, Lip60, MH53, Mal58, Ohl57, Pou56, Seb53, Wac52]. punched-card [Gar55, MH53]. punching [Per51]. punctuation [Fra66].

quality [ZD69]. quantitative [Lyn69, SK63, Tru65]. quarterly [Tat50b, Tru67]. Queiroz [Ano60i]. question [Flo67, Tag65, Lan65b]. Question-Answerer [Lan65b]. questions [Flo67, JC68, Kyl63, Sch69a, SKM64, Tay62, O’C68a].

R [Ano62d, Ano65a, Hil66, La 63, Sny66, Vai67a, Sha65]. R.I [Ano68g]. R10 [Boh64]. R10/1 [Boh64]. R62CD2 [Ano62c]. radiation [Lei69]. radiation- [Lei69]. Radio [Har63]. Rand [Mal58]. Random [McC65, Man69, Nol59, Uhl64, Wis52]. random-access [Man69]. Ranganathan [Eva66]. ranges [Per53]. Ranking [MB68]. ranks [Bro68]. Rapid [Bra51, McM62, WP50, Wis52]. rate [Ric68]. rationale [Sch57, Wal67]. Ray [Tre66]. Reaching [Wil52, Bau52, Bra52]. reaction [BH57a]. reactions [Ano57a]. read [Ric67]. readable [Rot56]. reader [Con51b]. Reading [Gar65a, SPS68, Ano65c]. realizable [Che63]. recall [Mon65]. receipts [Nee62]. Received [Ano62r, Ano62s]. Récherches [Vai67a]. Reclamation [Ano64d]. recognition [JT61b]. Recommendations [Bro52, Ric61, Mou67]. Record [Cla67]. Recorded [Kil61, Ano55c, Dye56, Ken60]. recording [Luh53]. records [BF58, Fis50, Har68c, Kin62, Mor60]. reduce [Ano54f, Lan62]. reduction [Orn52, Pow50]. Rees [Hay66a]. Referativnyi [Rad69]. Reference [Anz67, Eva52, Lit66, DSTW66, Flo67, JC68, Mit53, Mos56, OE58, PM59]. References [Ano64d, Ada67, MH53]. Referral [Ash66]. regarding [CB55]. regional [Par62]. related [Hol51, Mil57, Pot68]. relation [Mos56]. Relations [Ano62y, BH57b, HH53, Mos64]. Relationship [Ada67, Giu69, Hof67, Lip65b]. relationships [Kor64]. Relative [Gro68]. Relevance [O’C67, Art69, GB69, Hil64a, Mon65, Per66b, RS64, RB68, Tau65b, Vic60]. Relevancy [Ree62]. reliability [Bor64, Res61]. remarks [Jah69]. Remington [Mal58]. Remington-Rand [Mal58]. René [Lut50]. reply [Wis52]. Report [Ano54q, Ano62t, Ano62u, Ano62v, Ano62w, Ano62x, Ano62y, Ano62z, Ano62-27, Ano63a, Ano65j, Cla67, Dye56, Lan65b, Sny66, Swa65, Tho62, Web65, Ano62h, Bau52, Ben52, Bra64, Dav52, Gar54, Gul61, Hys53, KI59, Lan69a, Mal62a, Mil52, MK62, Nee62, Orr61, Ree65, She50, SE50b, She55, Sop68, WGP58, Art65]. reported [Car69]. reporting [Mil53]. Reports [Ano62-28, Ano62-29, Gra52a, Hoo52, Aro52, Can52, Car52, Con52, EFB57, Fag52, Fra52, Gra53, Hea52, Her52, KP57, Web58]. representations [Opl59]. representative [Wil52]. Representing [Pri69]. reproducer [Hue51]. Reproducing [Har67a, Sil51]. Reproduction [Web65, Ano50e, Con52, Irv51, Pow51b, Sha54, Tat50c]. Rept [Ano64b]. Republic [Sta61]. request [O’C67]. require [O’C61]. required [Hir62]. Requirements [Shi63, Cos62, Jah64, Tru65, Hoo52]. Research [Ano64b, Ano68f, Boh65, Cla63b, Dar66, Dav58, JT61a, KMW66, SS61, Ano57n, Ano58i, Ath64, BH51, Bis57, Blu59, Can52, Car52, Cos61a, Cro56, Dra54, EFB57, Fag52, Gar63, Gro51, Her52, KC69, Kyl63, Mai60, McG58b, Met51, Mil68b, Mor60, MK62, OC62b, Sar68a, Sco51, Sco53, Sha59, She59, Sil50b, Tag65, War56, WBO^{+}61, Wil52, Wor69, Ano55b, Ano57a, Har67c, Ano68d]. researchers [Man61]. Reserve [Ano55b, Ano55c, Ano69a, Har66, Hay66a, Ree65, Ano53b]. resolution [Bal69]. resolved [O’C69]. resource [Ada56, Hod68]. Resources [Ano64d, Ano68f, Ash66, Orn63, Bet59, Bor55, Cam56, HH67, Sto69]. respect [Sop68]. responsibilities [Kno60]. restoration [Bar53]. Restricted [Tum53]. restricting [Can52]. Results [O’C63, Alt67, Bro52, Fel64, Kes65, Sal65b]. Retrieval [Anz67, Ath65b, Boh64, Kin62, Lan67, O’C68b, Alt67, AR69, Ano64b, Ano69a, Ano69c, Ano69e, Ass55, BN68, BH63, BF58, Bor61, Bou61, Bra62, Buc61, Cla61, Coo68, Cos61a, Dai56, DD65, Die62, DSTW66, Fai56, Fel63, Fla69, Flo67, GVB64, HG65, Hil64b, Hof62, Jam59, Joh62, JN58, KBLP55, Ken64, Kle64, Lan68, Les69, Les60, Lib67, Mal58, Man69, Man61, MS66, MS67, Moo60, Mor62, Nol59, Pol68, Sal63, Sal65b, SSS61, Sch63, Sch64b, SM63, SL68, SWCT68, Ste61, Swe69, Tre68, Uhl69, Vas69, Ven68, Vic59a, Wol59, Ano65a, Hay67, Lin66, Rub66, Whi66, Ano62c, Cle63, Lan65a]. retriever [Eng61]. Retrieving [Ano64d, Woo57]. retrospect [Cla63b]. retrospective [Fis66]. Rev [New69a, Ran67]. Reversible [Fra66, LB63]. Review [Ano62d, Ano62c, Ano64b, Ano64c, Ano64d, Ano65b, Ano65a, Ano66a, Ano68d, Ano68c, Ano68b, Ano68a, Ano68e, Ano68h, Ano68f, Ano68g, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Anz67, Art65, Ash66, Ath65b, Bal69, Bat67a, Boh62, Boh64, Boh65, Bra67, Cha63, Cla67, Cle63, Dai64, Dar66, Emb65, Eva66, FC68, Gar65a, Gre67, Har66, Har68a, Har68b, Har68c, Har69, Har67a, Har67c, Har67b, Hay66a, Hay66b, Hay67, Hil66, Hin69, How65, JT61a, La 63, Lan65b, Lan65a, Lan67, Lew66, Lin66, Lit64, Lit66, Log67, Mac62b, Met65, Mou65, Mou67, New63a, New69a, Nit67, O’C63, Pel69, Per66a, Pre69, Rad69, Ran67, Ric63b, Ric68, Rog67, Rub66, Sey66, She67, Sny66, Sta61, Ste69, Tau61a, Tau65a, Tre66, Tru67, Try69, Vai67a]. Review [Vai67b, Vas69, Wad68, War66, Web65, Wel61, Whi66, Ano69g, Boo60, Hay60, Her57, Heu56b, KP56, Ken58, Kle64, Lan69b, PK56, Ree65, Sha54, Tat50b, Vic59a, Vic59b, AS67, Pet59]. Reviews [Har63, New63b, O’C63, Pra63]. revised [Ano62a, FC68]. Revision [FC68, Fos58]. Richard [Har69, Har67c]. Rider [Mac62b, Mac62b]. Ridge [Gre67]. Robert [Boh64, Lin66, Mou67, O’C63, Ano60m]. role [Bey60a, Blu59, Dub62, Lan68, Sco51, Ste55]. Roles [AH63, Cos61a, Lan65a, Mon62, Mon65, Tau61b]. Roosevelt [Blu50]. Rosen [Ano68b]. Roster [Ano58i]. Routine [WGP58]. rule [Car56]. Rules [Nit67]. Rumanian [Sta61]. Russian [Fis59, Fra66, Kut64]. Russians [PK56]. Rutgers [Bra67, Cle63, Lew66, Met65, Per66a, Rog67].

S [Cha63, Har68c, Lan65b, Mou67, O’C63, Tre66, Sha65]. safety [BH67, Cot68]. Sambaquy [Ano60i]. Saul [Ano68b]. Saunders [Pre69]. scale [BF58, Nol59, Orn63, Swa68]. scan [O’C62a]. scan-column [O’C62a]. Scarecrow [Cha63, Log67, Pel69, War66]. schedules [Sla63]. scheme [BN69, Dav68, Joh62]. schemes [Ano62d, Ric54, Per66a]. Scher [Pre69]. Schmidt [Ano69a]. scholar [Dic66]. Scholarly [Cla67, Sil51]. scholars [Hin61]. School [Boh62, Bra67, Har66, Lan65a, Lew66, Lit64, Met65, Per66a, Art69]. Schultheiss [Cha63]. Schultz [Ano69b, New69a]. Science [Ano59m, Ano62o, Ano69b, Ash66, Bat67a, Bis59, Bra67, Eva66, Fuk67, Gre67, Hay66a, Hil66, Mat62, Ano62p, Ano68a, Art69, Bey55, Bey57, Boh62, Bor68, BS68, BGM56, Che65, Coo66, Cua64, DO69, DK66, Don66, Giu69, Har63, Hay69, Hil69, Hoy69, JC68, Kle69, Kyl58, LD60, Lip65b, Lit64, New69a, Ort61, Poh65, Ric67, Ros65a, Sch69a, SB59, Swa65, Tar62, Try69, Yov69, Ano68p, Ano68q, Ano69c, Har63, Lit66, Low68, SB59, Spe67, Pel69]. science-fiction [Poh65]. Sciences [Ano64c, Ano68e, Ano69e, Ash66, Boh62, Boh64, JT61a, Ega51, Fel57, Heu56b, Hof67, MW61, Mac63b, Anz67]. Scientific [Ano53c, Ano59m, Kik67, Lie54, Par62, Sim62, Ada55, Ano68h, Bau66, Bey60a, Bey62, Bin63, BMR65, Bra51, Bro52, BK60, CS62, Cla63b, Cot68, Fis59, GS63, GG64, Giu69, GD63, JT61a, Kes59, Kes63, Lip62, Luh61, Mac60b, Man66, Mar60, Orn63, Orr61, PV63, Pre69, Ric63a, Roy57, Sar68b, Sco53, Shi63, Sop68, Sta61, Tar62, VZ68, Yag65, Yov69, Log67]. Scientifique [Vai67b]. scientist [Bin63, Bis57, Hof67, Sar68b]. Scientists [AL68, And64, FCHB65, JHG66, Kot62, Sch69b]. SDI [Dav67, Jor68, Sav67]. SEAC [PKM59]. Search [Stu69, War66, Bak69, BH57a, Boh55, BE65, Car63, Coo68, Dav69, Dyk59, Lei68, MB68, New61, TBW68, Uhl68]. Searching [Bal55, Hol62, Ano53b, BPK54, Har68b, Her62, Hys61, KBP54a, KBP54b, KPB54, KBLP55, Ken60, Kyl63, LKPB55, Luh53, Man64, Mil68b, OE58, Ove63, Per50, PKB54, PBK54, PKB55a, PKB55b, PKM59, RK58, Spe67, Tri64, O’C63]. second [Lib67]. Secondary [OC62b]. Secretary [Ano62h]. Secrets [SM67]. Section [Hys53]. security [Hea52]. see [Mos56]. see-also [Mos56]. Seidell [Sew62]. Selected [Ano69b, Gar65a, Gre67, Boh62, Sal65b, Lit64]. selecting [Res61]. selection [CR69, RRS61b, Ros65a, Sha54, Vic51]. Selective [Luh61, Fla69, TEC64]. selectivity [Lip65b]. selector [McM62, RK58, WP50, Wis52]. Self [Har67a]. Self-Reproducing [Har67a]. Selye [Rog67]. semantemes [BH57b]. Semantic [Mal62b, LKPB55, Mel62, Tau62, Vic59b, BH57b]. Senate [Rey61]. sense [CG62]. Sentence [RRS61b]. sentences [RRS61b, Res61]. separately [Tru67]. Sept [Ano54n]. September [Sny66, Ano54q]. sequences [Bak69]. Serendipity [Ber60]. serial [Fis59]. Serials [NT58, BMR65, FL69, GD63, Rop68]. Series [Ano62c, Ano64c, Ano68e, Ano69e, Boh64, Bra67, Eva66, Lew66, Lit66, Met65, Per66a, Rog67]. Service [Ath65b, Bra67, Dai64, Lew66, Met65, New69a, Per66a, Tru67, Cha50, Dav52, Dav58, Gra52a, Hol50, Ken60, MK62, Ove63, Sha59]. services [Ano59m, Bar56, Bro52, Cla56, Fel57, KP57, Meh57, Mel68, Mil57, Mor55, Swa68, SS59, Web58, WBO^{+}61, She50]. servicing [Gra53]. sets [Uhl68]. Seven [Gul56b]. several [HH53]. Sharp [Ano65a, KMW66, SS61]. sheet [BE65]. shelf [Pou56]. SHIEF [Che63]. Shilling [Har67b]. Ships [McM62]. Short [Ano51j]. Shorter [Ano50f, Ano50g, Ano50h, Ano51k, Ano51j, Ano52l, Cla50]. Shorthand [Rog67]. siege [Lut50]. sigils [Rei54, Bis54]. significance [Buc58]. significant [Fei63]. signs [RSF^{+}67]. Silver [Ano63k]. Simonton [Anz67]. simple [Sha53, Car56]. Simulation [Bra66, BN68, BF64, Car68]. since [Wad68]. single [Bis56, Coo68, JS69]. Sinnett [Lan65a]. sketch [SG69]. sketches [Sch69b]. SLA [Ano62d, Fos58]. SLIC [Sha66]. slide [Car56]. SLIP [WS64]. small [Car68, Kil66, Swa68]. small-scale [Swa68]. SMART [Sal65b]. SNOBOL [Ano68c]. social [Buc58, Ega51, Ega53, Fel57, Hof62, Hof67, Kyl58, Tum53, Ull53]. societies [Phe58, Den50]. Society [Ano68p, Ano68q, Try69, Cla53]. sociological [Gar63]. Soergel [Ste69]. Sold [Ano68h, Tru67]. solitary [Ano60d]. Solution [Ike67, Low68, TEC64]. solutions [Cos61c, Cos61b, Hof67, Lou66]. Some [Bor54, CB55, Cor54, Cos61c, DD65, Kyl63, Lip65a, O’C61, O’C68a, O’C69, OE58, Que63, Ric54, Sal63, TPBK55, Vic51, Wol59, BK60, HS68, Lou66, Moo60, Ano65a, BN69, Sch69a]. Somerset [Ano62h]. Sons [Ano64c, Ano68a, Ano68e, Ano69e, Bat67a, Boh64, Hil66]. sort [Ber58]. sorted [Per51]. Sorting [DG57]. source [BKB62, Dav56]. Sources [Hil66, Mou67, Gro51, Ste69, Jac52]. Southeast [Mat62]. Soviet [Bey57, Bey60b, Bui62, KI59, Lie54, Mac60b, WP62]. SP [Ano68h]. SP-7030 [Ano68h]. Space [Ano68h]. Spangler [Ano62c]. Spanish [How65, Tre66].

Spartan [Ano69b]. Special [Ano62d, Anz67, Eng61, Gre67, Hei65, Ran67, Tat50c, Tau52b, Ano52f, Gul56b, Hay57, HH60, Man66, PM59, Hys53]. Specialists [Har67b, ABC64]. Specialized [Per53, Cla53, FCHB65, HH53, LBL67, Seb53, She53]. specific [MK62]. specification [Uhl68]. Specifications [Ano65c]. Speed [Spe60, Sha50]. spelling [Uhl64]. St [Boh62, Har66]. St. [Bea69]. Stability [Koc69, Bro68, LBL67]. stack [Tru65]. standard [Ano50e, MH53, PKM59]. Standards [Art65, Bor54]. standing [Ano62-29, Ano62-27]. State [Ano62-30, Ano69d, Art65, Bey60a, Bra67, Lew66, Mar62, Mar66, Met65, New63b, SB59, BH51, Bou61]. State-of-the-Art [Art65, Bou61]. statement [SD55, Nee62]. statements [Uhl69]. States [Ash66, HMC63, Met51, SE50a, SE50b, Sim62]. Station [Har63, Blo57]. Stationery [Pre69]. Statistical [SR68, Gar54]. statistics [HH60]. status [Sop68]. Stenafax [Con51a]. Stevens [Art65, Bal69, Cle63]. stiintifica [Sta61]. stockpile [Hut55]. Storage [Ass55, Boh64, Cos61a, Irv51, Ven68, Alt67, Ano62c, Ano64b, BN68, Bra62, Buc61, Cla61, CM62, Fei63, Her52, Her62, Kle64, MS66, MS67, Nol59, Rid51a, SWCT68, TM62, Vic59a]. storing [Eva68, Woo57]. strategies [Hol62, Tri64]. strategy [Uhl68, Wil69]. Street [Lit66]. structural [CBF59, OB59, Per66b]. structure [GG64, Mos56, She53, TW53, TBW68, VZ68, Vic59b, WAF^{+}58]. structures [Cho57, Fel64, Opl59]. Stuart [Hil66]. student [Bui62]. Studies [Buc61, Dai64, Her57, Sla69]. Study [Cla67, Mos56, Ano52l, Ano69f, BKRL66, CS62, Cle63, GG64, HH53, Hur69, Hys63, MW61, Mai60, Mil68b, Pin67, Ric61, SK63, Ste61, Tya61, Web65, Hin69]. studying [BN68]. Subcontractor [KB68]. Subject [Ano66a, Ano68h, Cla52, Dai57, Lew66, Spe67, Wåh66, Way68, Ano62d, Ano62q, Bor64, Dai56, JS69, Kes65, Kra64, Lil54, Lip60, Mor54, Mos56, Ort61, Per53, Sla64, Tau52a, Ath65b]. subjects [Ken60, Wåh66]. subscribers [Tat50c]. substitute [Har68d, Pow51a]. successful [Lip60]. suggested [AC61, O’C61]. Suggestions [Ric67]. suggestiveness [Ber60]. summary [Bel69a, SD55, Ste61]. Superimposed [Per51, Sti60, Uhl64]. Superintendent [Ano68h, Art65, Tru67]. superior [Ano65c]. Supersedes [Tru67]. supplied [SSO65]. Supported [Bra67]. Supt [Ano68f]. surrogation [Lan69b]. Survey [BS68, Bey62, BC63, Car69, Gar65b, Lev64, Mac60a, Hay67, Ano64b]. Susan [Bra67, Met65, Pel69]. Switzerland [Aue55]. Symbolic [Rog67]. symposium [Ano62-30, Bin63]. syntax [CHJ61]. SYNTOL [Met65, Vai67a]. System [EFB57, Lan67, Lit66, AR69, BHR67, BF58, BK63, BML64, Bis57, BKRL66, Bor67, Bra62, BE65, BBJB69, Buc58, Car68, CBF59, DSTW66, Dyk59, Ebe66, FL69, Fel63, Fla69, Flo67, Gul56a, Hil64b, Hof62, HH60, HA65, Jam59, JH66, KP57, KMW66, KB68, Kra64, KTM57, Mal58, McM62, Mit53, Par66, PM59, Pla60a, Pot68, Ree65, RH63, Rop68, Sal65b, SS61, Sch62b, Seb53, SWCT68, TMK55, Vic60, WD59, WW67, Wol59, Wor69, Bra62, Rog67]. systematically [Sav63]. Systematics [San68]. systematization [KPB54].

Systems [Ano64b, Bra67, Jon63, Lew66, Met65, Mou67, Per66a, Rog67, Wåh66, Ano68b, Ano69e, BN68, Bak69, BH57a, Bou61, BF64, Cle63, Coo68, Cro56, GVB64, Hys63, Jon60, Jor68, KBLP55, Kle64, Lan68, Les69, Lib69, Lip63, O’C64b, Pol68, Rau67, SSS61, Ste69, Ste61, Swa69, Uhl69, Wal68, Wha61, Jon63, Wåh66, Vas69]. szakirodalmi [Emb65].

T [Ano69a, Har69, Sey66, Sha65]. T. [NT58]. Tab [Har68b]. Tab-indexed [Har68b]. tables [Cor54, Ste69]. tabulating [RSF^{+}67]. tailored [ABC64]. tape [Ric68]. tapes [BKB62, BML64]. tasks [Eva50]. Taube [CG62]. Taylor [Mou67]. TDR [Ano64b]. Teachers [FC68]. teaching [Sar68a]. Technical [And64, Ano62c, Ano64b, Bey60a, Coh59, Dav58, Emb65, Lan65b, Mor55, New63b, Ano50f, Ano50g, Ano50h, Ano51k, Ano51j, Ano57n, Ano59m, Ano68h, Bar56, Bau66, BH57b, Bet59, BKRL66, Boh55, Bou62, Bra64, BH67, BK60, Car52, Cla50, Cob56, Con52, Cot68, DSTW66, GD63, Hel69, HR61, HA65, Lan62, Les60, Luh60, Mac60b, Mas54, Mil53, New61, New63a, RH63, She55, SR68, TN68, Web58, Kik67, Pol50, Tre66]. technique [Dew59, Gul51, Jah64, Mac53, Sch57, Sho62]. techniques [Bes61, Bui62, Gar55, Jac52, Jor59, Lyn69, McC61, RH63, Tat50b, Wal69, War56, Lit64]. technological [Pre69]. Technology [Ano68h, Ash66, Boh62, Fuk67, Hei68, La 63, Lan65b, Lan65a, Web65, JC68, Lit64, Pie64, Web58, Bat67a, Hil66, Ano68a, Try69]. Tekhnicheskaya [Rad69]. Tekhnicheskoi [Rad69]. Teller [Tho62]. Tennessee [Gre67]. Tentative [Ano64d]. term [Bra66, HW64, JC67, Ike67]. Termatrex [BHR67, Jon60, Jon63]. terminologies [Flo67]. terminology [BPK54, KPB54, Tag65, TH69, How65, New69a]. terms [Ano68h, Bro68, Pot68, RB68, TGW52, Wag60, ZS67, SSO65, Ano69a]. Test [Ove62, Alt67, Blo57, O’C63, Ree65, Sal65b]. Testing [Ano69e, LM64, Mon65, Alt67, AR69, Ano58i]. tests [RSF^{+}67]. text [Nit67]. Textbook [Rub66]. texts [BK63, Bor67]. Textual [GB69]. Thailand [PV67]. Their [Gar65a, CBF59, Fel64, Hof67, Jon60, Kot62, Lou66, Ruh64, Wåh66]. theme [Bel69b, Oat68, ML67]. Theodore [Blu50]. Theoretical [Man64, BN69, Eva56]. theories [Hil69, Min57, Mos64, Pin57, Boh64]. theorizing [BH57a]. Theory [AR69, Gar65a, Lit64, War66, Dai57, Dew59, Gra68, Jon54, Har67a, Lan67]. Thermit [Mor60]. Thermo [Con51a]. Thermo-fax [Con51a]. Thesauri [Ste69]. Thesaurus [EP68, O’C64a, Ano68h, Ano69a, Ano69g, BH57b, Har68b, HH67, HR61, JN58, Wai69, Ano64d, Ano68f, New69a]. Thesis [Lan65a, Bis59]. thinking [Bra68, CG62]. thinning [Tru65]. Third [Ano69d]. Thompson [BF57]. those [Kna65, Wad68]. thought [CB55]. three [Cle63, O’C64b, Ste61]. thru [Nee62]. Time [Hir62, Lec68, TBW68]. Times [Har68b, Ano69g]. Title [Ano66a, JS69, O’C64b, Vor65]. titles [Ada67, Bor67, Gro68, Kna65, Kra64, Ruh64, SSO65, Wad68]. titles-cited [Gro68]. Titling [Car50]. Today [Eva66, Mel62]. Too [Sch62b]. tool [And64]. Tools [Boh64, She59]. Toronto [Ano66a]. tortenetenek [Emb65]. toxicity [RSF^{+}67]. Traditional [Eva68]. Training [Ano62u, Bon62, Bui62, Har67b, Pla60b, Bat67b, BS68, Giu69, Pre69, Rau67, Tau53]. transactions [SB59]. transcription [Hof52]. Transfer [Sny66, Lib69, ML67]. translation [BH51, Hol66, LB56, Mac60a, Mar60, Tau61a]. transliteration [Fra66]. transmission [Ano69d, Kin60, TH69, ZS67]. transmitted [CM62]. transportation [Hod68]. Trautman [Tre66]. Treasurer [Nee62]. treasures [Smi55]. trend [McC62]. trends [Bey62, Ull53]. Trillas [Tre66]. Trip [Mal62a, KI59]. tronic [Mit53]. trope [Ber57a]. true [Yov69]. Trustees [Ano52g]. tumors [HH60]. turnaround [KB68]. twentieth [Kno60]. twenty [Moo60]. Two [Boh55, Hil64b, Car69, Hys63, Mos64, SSS61, Ike67]. type [Cum54, Yag65]. Types [Wåh66, RRS61a]. typescript [Har68d]. Typesetting [Sey66]. typewriter [Cum54].

U.D.C. [Kut64]. U.S [Ano64d, Ano68f, Art65, Ash66, Tru67, Wel61]. U.S. [Lar52]. U.S.A [Har67b]. U.S.S.R. [Mac58, Mac60a]. UCLA [FL69]. Unanswered [JC68]. unclear [O’C67]. underlying [PKB55b]. understanding [GB69]. UNESCO [Cla56, Mat62, PV63, She50, Car54]. UNESCO/NS/177 [PV63]. unified [Ort61]. unifying [Hil69]. Union [Hir50, Sch57, Bey57, Lie54, Wad68]. Unit [TGW52, Wad58]. United [Ash66, Cob56, Har51, HMC63, Met51, SE50a, SE50b, Sim62, Urq56]. Uniterm [Cos61b, Gul56a]. units [Ath64, LKPB55]. Univ [Mit53]. Universal [Lit66, Har54, McC65]. Universities [Ano66a]. University [Ano66a, Ano68g, Ano69a, Anz67, Boh62, Boh65, Bra67, CG62, Cle63, Dai64, FC68, Har66, Har67a, Hay66a, Hay67, Hin69, Lan65a, Lew66, Lit64, Met65, Mou67, New63b, Per66a, She67, Tau61a, Bro59, Cha63, JHG66, Sha59, She56, Wil51b, Ano55b, Ano55c, Ree65]. unlimited [Kle69]. unp [Boh62]. unpublished [Gra53, Her52]. Unthinking [Bra68]. Urbana [Boh65]. USA [Bey60b]. usage [HH67, HW64]. Use [Bis57, GVB64, Wor69, BN68, Bra66, Buc61, BGM56, Ega53, Eva52, Gar54, Hol51, JHG66, Jon60, Kil61, Lib67, Mas54, McC62, McM61, Mit53, Pin67, Pow50, RH63, SSS61, Tau61b, Ull53, Hay66b, Ike67]. used [Lan65a]. User [Con51b, Swa69, Bak69, Eng61, GB69, Par66, Tru65]. User-oriented [Swa69]. Users [Sav66]. uses [CBF59]. Using [Cue68, CR69, DG57, EFB57, Mon62, Nol59, RK58, Sal65b, Spe67, Uhl68]. USNCFID [HMC63]. USSR [Bey60b, Tar62]. Utilization [Ano68h, Ano55c, Bis61, Dye56].

v [Ber58, Ste69, KPB54]. valid [Sti58]. value [Ros65b]. various [Sti60]. varying [Man61]. Vatican [McG58a]. verbal [BK63]. Vern [Cla67]. vernacular [Orn54]. Verner [Boh65]. versus [Ano51h, Bis54, Gro68, RB68, Sav66]. VI [Rog67, Ber60, PBK54]. viable [BN69]. Vickery [Lan67, Per66a]. video [Wor69]. video-display [Wor69]. view [Eas51, Fis66, Wil51a]. VII [Bra67, LKPB55]. VIII [KBLP55]. Villars [Vai67a, Vai67b]. Virgil [Sta61]. Virginia [Wil51b]. virology [Car69, SPS68]. vocabularies [Dav68]. Vocabulary [Wal69, SR68]. Vol [Bra67, Dar66, Eva66, Lan65b, Lew66, Lit66, Per66a]. vols [Har68b, JT61a, Ano66a]. Volume [Ano68a, Bat67a, Hil66, Met65, Rog67, Tru67, Try69, Bou62, Her57]. volumes [Ano68h, La 63]. vs [Rid51a, Ruh64, Wis52]. Vsesoyuznyi [Rad69]. Vypusk [Rad69].

W [Ano64b, Ano68g, Bal69, Boh65, Har68c, Har67a, Har67b, New63b, Pre69]. Walnut [Bra62]. wares [Mel68]. Washington [Ano68h, Ano68f, Art65, Ash66, Boh62, Har66, Har67c, Har67b, JT61a, Lan67, Lit64, New69a, Tru67, Wel61, Cla55]. Water [Ano64d, WA62, HH67]. way [Hay57]. Ways [Bra52, Bau52]. weak [Coo68]. weeding [Slo68]. weights [AW69, Bra66, Ike67]. Weisman [New63a, New63b]. Welfare [Wel61, Hof62]. Wesley [Anz67, Gar65a]. Western [Ano69a, Hay66a, Ano55b, Ano55c, Bet59, Bey57, Lev64, Ree65]. Wheeler [Pra63]. which [Fel64, O’C61]. Whole [Dar66]. Wiley [Ano64c, Ano68a, Ano68e, Ano69e, Bal69, Bat67a, Boh64, Hil66]. Wilfred [Ano69e]. William [Dai64, La 63, Whi66, Dai64]. Williams [Whi66]. Williamson [Dai64]. win [Hil69]. Winifred [Gre67]. within [SSS61]. without [Tri64]. Wolf [Lan65b]. Word [Les69, CR69, LBL67, Luh60, RB68, Wis52]. word-frequency [CR69]. word-in-context [Luh60]. words [Kna65, LB63, O’C64b, Pot68, Sav63, Spe60]. work [Bon62, Gul56b, Tau52a, Hay66b, Pre69]. Works [Ano69b]. World [Hin69, Bou62, Cue68]. Worldwide [GD63]. Wright [Ano64b, Lan65a]. Wright-Patterson [Ano64b, Lan65a]. writing [Ano67a, Kes59, WGP58, New63a]. WRU [RK58, Tau62].

X [PKB55b]. xerographic [CM62]. Xerox [Con51a]. xix [Tau61a].

Yale [Kil61]. year [Tru67]. years [Dav51a, Gul56b, Moo60]. Yehoshua [Gar65a]. York [Ano62d, Ano64c, Ano65b, Ano68b, Ano68a, Ano68e, Ano69b, Ano69d, Ano69e, Ano69f, Ano69g, Ath65b, Bat67a, Boh64, Cha63, Cla67, Dai64, Dar66, Eva66, Har68b, Har69, Har63, Hil66, Hin69, How65, Lit66, Log67, Mou65, Pra63, Ran67, Ric68, Rub66, Tau65a, Wad68, War66, CG62]. Yugoslavia [Bey55, Bry55].

Zatocoding [Moo51]. Zhurnal [Rad69]. zur [Ste69].

References

Anzalone:1964:NIT

[ABC64] Alfred M. Anzalone, Carol A. Brokars, and Gunther Cohn. A novel index tailored to plastics specialists. American Documentation, 15 (3):191–195, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Atherton:1961:SCL

[AC61] Pauline Atherton and Virginia Clark. A suggested classification for the literature of documentation. American Documentation, 12(1): 38–48, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Adair:1955:CIS

[Ada55] W. C. Adair. Citation indexes for scientific literature? American Documentation, 6(1):31–32, January 1955. CODEN AMDOA7. ISSN 0096-946X.

Adams:1956:INR

[Ada56] Scott Adams. Information — a national resource. American Documentation, 7(2):71–75, April 1956. CODEN AMDOA7. ISSN 0096-946X.

Adams:1967:RKT

[Ada67] Wm. Mansfield Adams. Relationship of keywords in titles to references cited. American Documentation, 18(1):26–32, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Artandi:1963:RLF

[AH63] Susan Artandi and Theodore C. Hines. Roles and links — or forward to cutter. American Documentation, 14(1):74–77, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Adams:1968:SMK

[AL68] Wm. Mansfield Adams and Laurence C. Lockley. Scientists meet the KWIC index. American Documentation, 19(1):47–59, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Allen:1959:TPL

[All59] Layman E. Allen. Toward a procedure for logically cataloguing knowledge. American Documentation, 10(4):296–315, October 1959. CODEN AMDOA7. ISSN 0096-946X.

Altmann:1967:MTN

[Alt67] Berthold Altmann. A multiple testing of the natural language storage and retrieval ABC method: Preliminary analysis of test results. American Documentation, 18(1):33–45, January 1967. CO-DEN AMDOA7. ISSN 0096-946X.

Anderson:1964:BCT

[And64] Mary Anderson. Brief communications: Technical information — A modern tool for scientists and engineers. American Documentation, 15(4):302–304, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Angell:1958:DJH

[Ang58] Richard S. Angell. David Judson Haykin, 1896–1958. American Documentation, 9(4):237, October 1958. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1950:COI

[Ano50a] Anonymous. Committee on organization of information. American Documentation, 1(3):154–163, August 1950. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1950:ECa

[Ano50b] Anonymous. Editor’s corner. American Documentation, 1(1):58, January 1950. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1950:ECb

[Ano50c] Anonymous. Editor’s corner. American Documentation, 1(3):173– 176, August 1950. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1950:OI

[Ano50d] Anonymous. Organization of information. American Documentation, 1(4):214–224, October 1950. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1950:PSM

[Ano50e] Anonymous. A proposed standard for the microphotographic reproduction of new papers. American Documentation, 1(1):46–50, January 1950. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1950:SCNa

[Ano50f] Anonymous. Shorter communications: News and technical notes. American Documentation, 1(1):51–57, January 1950. CODEN AM-DOA7. ISSN 0096-946X.

Anonymous:1950:SCNb

[Ano50g] Anonymous. Shorter communications: News and technical notes. American Documentation, 1(3):164–172, August 1950. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1950:SCNc

[Ano50h] Anonymous. Shorter communications: News and technical notes. American Documentation, 1(4):207–213, October 1950. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1951:CCB

[Ano51a] Anonymous. Chicago conference on bibliographic organization. American Documentation, 2(1):5–7, January 1951. CODEN AM-DOA7. ISSN 0096-946X.

Anonymous:1951:ECa

[Ano51b] Anonymous. Editor’s corner. American Documentation, 2(3):181– 184, August 1951. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1951:ECb

[Ano51c] Anonymous. Editor’s corner. American Documentation, 2(4):259– 268, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1951:OIa

[Ano51d] Anonymous. Organization of information. American Documentation, 2(1):59–68, January 1951. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1951:OIb

[Ano51e] Anonymous. Organization of information. American Documentation, 2(2):116–129, April 1951. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1951:OIc

[Ano51f] Anonymous. Organization of information. American Documentation, 2(3):171–180, August 1951. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1951:OId

[Ano51g] Anonymous. Organization of information. American Documentation, 2(4):252–259, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1951:OVM

[Ano51h] Anonymous. Original versus microphotographed editions of documents. American Documentation, 2(3):150, August 1951. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1951:PN

[Ano51i] Anonymous. The President’s newsletter. American Documentation, 2(4):185–201, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1951:SCS

[Ano51j] Anonymous. Short communication: Shorter communications, news and technical notes. American Documentation, 2(2):108–115, April 1951. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1951:SCN

[Ano51k] Anonymous. Shorter communications, news and technical notes. American Documentation, 2(4):241–251, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1952:ANA

[Ano52a] Anonymous. Announcements: Notices and announcements. American Documentation, 3(1):82–89, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1952:AIC

[Ano52b] Anonymous. Articles of incorporation, constitution, and by-laws. American Documentation, 3(4):193–198, ???? 1952. CODEN AM-DOA7. ISSN 0096-946X.

Anonymous:1952:Ea

[Ano52c] Anonymous. Editorial. American Documentation, 3(3):i, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1952:Eb

[Ano52d] Anonymous. Editorial. American Documentation, 3(4):i, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1952:MDA

[Ano52e] Anonymous. Minutes of the deferred annual meeting of adi. American Documentation, 3(1):1–7, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1952:MSM

[Ano52f] Anonymous. Minutes of the special meeting. American Documentation, 3(4):198–207, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1952:MTM

[Ano52g] Anonymous. Minutes of the Trustees’ meetings, ADI. American Documentation, 3(1):8–11, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1952:OIa

[Ano52h] Anonymous. Organization of information. American Documentation, 3(1):71–81, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1952:OIb

[Ano52i] Anonymous. Organization of information. American Documentation, 3(2):123–132, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1952:OIc

[Ano52j] Anonymous. Organization of information. American Documentation, 3(3):182–191, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1952:OId

[Ano52k] Anonymous. Organization of information. American Documentation, 3(4):226–235, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1952:SNC

[Ano52l] Anonymous. Shorter notices: Committee of consultants for the study of the indexes to medical literature published by the army medical library. American Documentation, 3(4):219–222, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1953:AAM

[Ano53a] Anonymous. Announcements: Annual meeting of the American Documentation Institute. American Documentation, 4(4):185–186, October 1953. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1953:ACR

[Ano53b] Anonymous. Announcements: Case-reserve conference on information searching for industry. American Documentation, 4(4):186, October 1953. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1953:AIS

[Ano53c] Anonymous. Announcements: The Institute of Scientific Films. American Documentation, 4(4):186, October 1953. CODEN AM-DOA7. ISSN 0096-946X.

Anonymous:1953:Ea

[Ano53d] Anonymous. Editorial. American Documentation, 4(1):ii, January 1953. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1953:Eb

[Ano53e] Anonymous. Editorial. American Documentation, 4(2):ii, April 1953. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1953:Ec

[Ano53f] Anonymous. Editorial. American Documentation, 4(3):136, August 1953. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1953:Ed

[Ano53g] Anonymous. Editorial. American Documentation, 4(4):ii, October 1953. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1953:P

[Ano53h] Anonymous. Preface. American Documentation, 4(3):ii, August 1953. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:A

[Ano54a] Anonymous. Announcements. American Documentation, 5(1):39, January 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:Ea

[Ano54b] Anonymous. Editorial. American Documentation, 5(1):ii, January 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:Eb

[Ano54c] Anonymous. Editorial. American Documentation, 5(2):i, April 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:Ec

[Ano54d] Anonymous. Editorial. American Documentation, 5(3):i, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:Ed

[Ano54e] Anonymous. Editorial. American Documentation, 5(4):i, October 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:HRC

[Ano54f] Anonymous. How to reduce cataloging costs. American Documentation, 5(3):146–154, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:LNa

[Ano54g] Anonymous. Literature notes. American Documentation, 5(1):30– 38, January 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:LNb

[Ano54h] Anonymous. Literature notes. American Documentation, 5(2):105– 114, April 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:LNc

[Ano54i] Anonymous. Literature notes. American Documentation, 5(3):177– 189, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:Mb

[Ano54j] Anonymous. Masthead. American Documentation, 5(1):fmi, January 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:Mc

[Ano54k] Anonymous. Masthead. American Documentation, 5(2):fmi, April 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:Md

[Ano54l] Anonymous. Masthead. American Documentation, 5(3):fmi, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:Me

[Ano54m] Anonymous. Masthead. American Documentation, 5(4):fmi, October 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:MLS

[Ano54n] Anonymous. Membership list, Sept. 1, 1954. American Documentation, 5(4):247–251, October 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:MAa

[Ano54o] Anonymous. Miscellaneous announcement. American Documentation, 5(2):115, April 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:MAb

[Ano54p] Anonymous. Miscellaneous announcement. American Documentation, 5(3):190, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1954:RML

[Ano54q] Anonymous. Report of meeting, London, September 1953. American Documentation, 5(2):101–104, April 1954. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:ACA

[Ano55a] Anonymous. Advisory Committee on Application of Machines to Patent Office Operations. American Documentation, 6(1):56, January 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:CDC

[Ano55b] Anonymous. A Center for Documentation and Communication Research at Western Reserve University. American Documentation, 6 (3):178, July 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:CPU

[Ano55c] Anonymous. Conference on the practical utilization or recorded knowledge at Western Reserve University. American Documentation, 6(4):255–258, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:DD

[Ano55d] Anonymous. Definitions of documentation. American Documentation, 6(4):254–255, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:Ea

[Ano55e] Anonymous. Editorial. American Documentation, 6(1):ii, January 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:Eb

[Ano55f] Anonymous. Editorial. American Documentation, 6(2):ii, April 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:Ec

[Ano55g] Anonymous. Editorial. American Documentation, 6(3):ii, July 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:Ed

[Ano55h] Anonymous. Editorial. American Documentation, 6(4):ii, October

1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:Ee

[Ano55i] Anonymous. Errata. American Documentation, 6(4):242, October

1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:LNa

[Ano55j] Anonymous. Literature notes. American Documentation, 6(1):40–

55, January 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:LNb

[Ano55k] Anonymous. Literature notes. American Documentation, 6(2):106–

125, April 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:LNc

[Ano55l] Anonymous. Literature notes. American Documentation, 6(3):179–

205, July 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:LNd

[Ano55m] Anonymous. Literature notes. American Documentation, 6(4):260–

268, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:Ma

[Ano55n] Anonymous. Masthead. American Documentation, 6(1):fmi, Jan-

uary 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:Mb

[Ano55o] Anonymous. Masthead. American Documentation, 6(2):fmi, April

1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:Mc

[Ano55p] Anonymous. Masthead. American Documentation, 6(3):fmi, July

1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:Me

[Ano55q] Anonymous. Masthead. American Documentation, 6(4):fmi, Octo-

ber 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:MD

[Ano55r] Anonymous. Minicard demonstration. American Documentation, 6 (4):258–259, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:NNa

[Ano55s] Anonymous. News note. American Documentation, 6(2):126–127,

April 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:NNb

[Ano55t] Anonymous. News note. American Documentation, 6(3):206–209,

July 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1955:NNc

[Ano55u] Anonymous. News note. American Documentation, 6(4):269–271,

October 1955. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:E

[Ano56a] Anonymous. Errata. American Documentation, 7(3):265, July 1956.

CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:LNa

[Ano56b] Anonymous. Literature notes. American Documentation, 7(1):53–

61, January 1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:LNb

[Ano56c] Anonymous. Literature notes. American Documentation, 7(2):138–

161, April 1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:LNc

[Ano56d] Anonymous. Literature notes. American Documentation, 7(3):235–

260, July 1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:LNd

[Ano56e] Anonymous. Literature notes. American Documentation, 7(4):333–

350, October 1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:Ma

[Ano56f] Anonymous. Masthead. American Documentation, 7(1):fmi, Jan-

uary 1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:Mb

[Ano56g] Anonymous. Masthead. American Documentation, 7(2):fmi, April

1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:Mc

[Ano56h] Anonymous. Masthead. American Documentation, 7(3):fmi, July 1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:Md

[Ano56i] Anonymous. Masthead. American Documentation, 7(4):fmi, October 1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:NNa

[Ano56j] Anonymous. News notes. American Documentation, 7(1):62–63, January 1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:NNb

[Ano56k] Anonymous. News notes. American Documentation, 7(2):162–165, April 1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:NNc

[Ano56l] Anonymous. News notes. American Documentation, 7(3):261–265, July 1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1956:NNd

[Ano56m] Anonymous. News notes. American Documentation, 7(4):351–353, October 1956. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:CLR

[Ano57a] Anonymous. Comments on “A logician’s reactions”. from the Center for Documentation and Communication Research. American Documentation, 8(2):117–122, April 1957. CODEN AMDOA7. ISSN 0096-946X. See [BH57a].

Anonymous:1957:LNc

[Ano57b] Anonymous. Literature note. American Documentation, 8(3):227– 235, July 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:LNa

[Ano57c] Anonymous. Literature notes. American Documentation, 8(1):59– 73, January 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:LNb

[Ano57d] Anonymous. Literature notes. American Documentation, 8(2):134– 143, April 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:LNd

[Ano57e] Anonymous. Literature notes. American Documentation, 8(4):314– 324, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:Ma

[Ano57f] Anonymous. Masthead. American Documentation, 8(1):fmi, January 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:Mb

[Ano57g] Anonymous. Masthead. American Documentation, 8(2):fmi, April 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:Mc

[Ano57h] Anonymous. Masthead. American Documentation, 8(3):fmi, July 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:Md

[Ano57i] Anonymous. Masthead. American Documentation, 8(4):fmi, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:NNb

[Ano57j] Anonymous. News note. American Documentation, 8(2):144–151, April 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:NNa

[Ano57k] Anonymous. News notes. American Documentation, 8(1):76–80, January 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:NNc

[Ano57l] Anonymous. News notes. American Documentation, 8(3):236–241, July 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:NNd

[Ano57m] Anonymous. News notes. American Documentation, 8(4):325–333, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1957:TLN

[Ano57n] Anonymous. The technical library in the new research laboratory of the National Carbon Company. American Documentation, 8(1): 74–75, January 1957. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1958:LNb

[Ano58a] Anonymous. Literature note. American Documentation, 9(2):143– 151, April 1958. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1958:LNa

[Ano58b] Anonymous. Literature notes. American Documentation, 9(1):59– 66, January 1958. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1958:LNc

[Ano58c] Anonymous. Literature notes. American Documentation, 9(3):215– 223, August 1958. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1958:LNd

[Ano58d] Anonymous. Literature notes. American Documentation, 9(4):312– 319, October 1958. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1958:NNb

[Ano58e] Anonymous. News note. American Documentation, 9(2):152–158, April 1958. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1958:NNc

[Ano58f] Anonymous. News note. American Documentation, 9(3):224–235, August 1958. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1958:NNd

[Ano58g] Anonymous. News note. American Documentation, 9(4):320–328, October 1958. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1958:NNa

[Ano58h] Anonymous. News notes. American Documentation, 9(1):67–72, January 1958. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1958:RCR

[Ano58i] Anonymous. Roster of current research, development and testing in documentation and librarianship, 1957–1958 Committee on Research and Development of the American Documentation Institute. American Documentation, 9(3):184–191, August 1958. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:LNa

[Ano59a] Anonymous. Literature notes. American Documentation, 10(1): 87–94, January 1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:LNb

[Ano59b] Anonymous. Literature notes. American Documentation, 10(2): 151–158, April 1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:LNc

[Ano59c] Anonymous. Literature notes. American Documentation, 10(3):

242–253, July 1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:LNd

[Ano59d] Anonymous. Literature notes. American Documentation, 10(4):

321–328, October 1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:Ma

[Ano59e] Anonymous. Masthead. American Documentation, 10(1):fmi, Jan-

uary 1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:Mb

[Ano59f] Anonymous. Masthead. American Documentation, 10(2):fmi, April

1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:Mc

[Ano59g] Anonymous. Masthead. American Documentation, 10(3):fmi, July

1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:Md

[Ano59h] Anonymous. Masthead. American Documentation, 10(4):fmi, Oc-

tober 1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:NNa

[Ano59i] Anonymous. News note. American Documentation, 10(1):95–104,

January 1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:NNb

[Ano59j] Anonymous. News note. American Documentation, 10(2):159–172,

April 1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:NNc

[Ano59k] Anonymous. News note. American Documentation, 10(3):254–265,

July 1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:NNd

[Ano59l] Anonymous. News note. American Documentation, 10(4):329–343,

October 1959. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1959:STI

[Ano59m] Anonymous. Scientific and technical information services in Japan. National Committee for Documentation Science Council of Japan.

American Documentation, 10(3):176–191, July 1959. CODEN AM-DOA7. ISSN 0096-946X.

Anonymous:1960:EMD

[Ano60a] Anonymous. Editorial: A mandate for documentalists. American Documentation, 11(4):ii, October 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:ENH

[Ano60b] Anonymous. Editorial: a new high at Lehigh. American Documentation, 11(2):ii, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:EML

[Ano60c] Anonymous. Editorial: Mooers’ Law. American Documentation, 11 (3):ii, August 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:ESE

[Ano60d] Anonymous. Editorial: “The solitary esophagus”. American Documentation, 11(1):ii, January 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:LNa

[Ano60e] Anonymous. Literature notes. American Documentation, 11(1): 56–65, January 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:LNb

[Ano60f] Anonymous. Literature notes. American Documentation, 11(2): 189–195, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:LNc

[Ano60g] Anonymous. Literature notes. American Documentation, 11(3): 247–260, August 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:LNd

[Ano60h] Anonymous. Literature notes. American Documentation, 11(4): 317–322, October 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:LQS

[Ano60i] Anonymous. Lydia de Queiroz Sambaquy. American Documentation, 11(3):205, August 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:NNa

[Ano60j] Anonymous. News note. American Documentation, 11(1):66–87, January 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:NNb

[Ano60k] Anonymous. News note. American Documentation, 11(2):196–203, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:NNc

[Ano60l] Anonymous. News note. American Documentation, 11(3):261–275, August 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1960:RAF

[Ano60m] Anonymous. Robert Arthur Fairthorne. American Documentation, 11(1):1, January 1960. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1961:ADI

[Ano61a] Anonymous. American Documentation Institute. Individual membership list — July 1961. American Documentation, 12(4):307–320, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1961:LE

[Ano61b] Anonymous. Letters to the Editor. American Documentation, 12 (2):157, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1961:LNa

[Ano61c] Anonymous. Literature notes. American Documentation, 12(1): 64–74, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1961:LNb

[Ano61d] Anonymous. Literature notes. American Documentation, 12(2): 145–155, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1961:LNc

[Ano61e] Anonymous. Literature notes. American Documentation, 12(3): 230–238, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1961:LNd

[Ano61f] Anonymous. Literature notes. American Documentation, 12(4): 296–304, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1961:NNa

[Ano61g] Anonymous. News notes. American Documentation, 12(2):156, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1961:NNb

[Ano61h] Anonymous. News notes. American Documentation, 12(3):239, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:ADI

[Ano62a] Anonymous. American Documentation Institute constitution and bylaws revised to February 1962. American Documentation, 13(2): 245–248, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:AAC

[Ano62b] Anonymous. Announcement: ADI convention announcement. American Documentation, 13(4):455, October 1962. CODEN AM-DOA7. ISSN 0096-946X.

Anonymous:1962:BRBb

[Ano62c] Anonymous. Book review: General bibliography in information storage and retrieval. By Marshall Spangler. Computer Department, General Electric Company, Technical Information Series R62CD2, 1962. 394 p. Limited Distribution from the author. American Documentation, 13(4):435, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:BRBa

[Ano62d] Anonymous. Book review: Guide to the SLA loan collection of classification schemes and subject heading lists. Compiled by Bertha R. Barden and Barbara Denison. New York, Special Libraries Association, 1961. 97 pp. American Documentation, 13(2):224–225, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:Ea

[Ano62e] Anonymous. Editorial. American Documentation, 13(2):244, April

1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:Eb

[Ano62f] Anonymous. Editorial. American Documentation, 13(4):454, Octo-

ber 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:FE

[Ano62g] Anonymous. The former Editor. American Documentation, 13(2): 243, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:IRE

[Ano62h] Anonymous. Interim report of the Executive Director and Secretary of ADI to the Annual Business Meeting, Hotel Somerset, Boston — November 8, 1961. American Documentation, 13(1):131–132, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:LE

[Ano62i] Anonymous. Letters to the Editor. American Documentation, 13 (2):238–242, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:LNa

[Ano62j] Anonymous. Literature notes. American Documentation, 13(1): 142–151, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:LNb

[Ano62k] Anonymous. Literature notes. American Documentation, 13(2): 226–237, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:LNc

[Ano62l] Anonymous. Literature notes. American Documentation, 13(3): 345–356, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:LNd

[Ano62m] Anonymous. Literature notes. American Documentation, 13(4): 438–452, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:MAD

[Ano62n] Anonymous. Miscellaneous: American documentation. American Documentation, 13(4):456, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:PSA

[Ano62o] Anonymous. The Pacific Science Association and its Pacific Science Congresses. American Documentation, 13(3):344, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:PSC

[Ano62p] Anonymous. Pacific science congress authors. American Documentation, 13(3):344, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:PBI

[Ano62q] Anonymous. Peek-a-boo index for a broad-subject collection. American Documentation, 13(2):187–196, April 1962. CODEN AM-DOA7. ISSN 0096-946X.

Anonymous:1962:PRPa

[Ano62r] Anonymous. Publications received: Publications received. American Documentation, 13(3):356, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:PRPb

[Ano62s] Anonymous. Publications received: Publications received. American Documentation, 13(4):437, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:RAA

[Ano62t] Anonymous. Report of the ADI auxiliary publications program. American Documentation, 13(1):134–135, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:RAC

[Ano62u] Anonymous. Report of the ADI Committee on Education and Training in Documentation. American Documentation, 13(1):136– 138, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:RAP

[Ano62v] Anonymous. Report of the Auxiliary Publications Committee. American Documentation, 13(1):135–136, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:RCB

[Ano62w] Anonymous. Report of the Committee on Bibliography. American Documentation, 13(1):135, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:RCF

[Ano62x] Anonymous. Report of the Committee on “the future of ADI”. American Documentation, 13(1):139, January 1962. CODEN AM-DOA7. ISSN 0096-946X.

Anonymous:1962:RIR

[Ano62y] Anonymous. Report of the International Relations Committee objectives and activities. American Documentation, 13(1):138–139, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:RMC

[Ano62z] Anonymous. Report of the Membership Committee. American Documentation, 13(1):140, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:RSCb

[Ano62-27] Anonymous. Report of the Standing Committee on Organization of Information. American Documentation, 13(1):141, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:ROA

[Ano62-28] Anonymous. Reports of officers of ADI. American Documentation, 13(1):125, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:RSCa

[Ano62-29] Anonymous. Reports of standing committees of ADI. American Documentation, 13(1):133, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1962:SAS

[Ano62-30] Anonymous. State of the art symposium. American Documentation, 13(1):1–2, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1963:ADI

[Ano63a] Anonymous. American Documentation Institute Committee on Organization of Information: Report for 1961–62. American Documentation, 14(4):299–306, October 1963. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1963:E

[Ano63b] Anonymous. Editorial. American Documentation, 14(3):187, July

1963. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1963:LNb

[Ano63c] Anonymous. Literature note. American Documentation, 14(2):169– 183, April 1963. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1963:LNc

[Ano63d] Anonymous. Literature note. American Documentation, 14(3):244– 270, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1963:LNa

[Ano63e] Anonymous. Literature notes. American Documentation, 14(1): 84–90, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1963:LNd

[Ano63f] Anonymous. Literature notes. American Documentation, 14(4): 313–327, October 1963. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1963:Ma

[Ano63g] Anonymous. Masthead. American Documentation, 14(1):fmi, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1963:Mb

[Ano63h] Anonymous. Masthead. American Documentation, 14(2):fmi, April

1963. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1963:Mc

[Ano63i] Anonymous. Masthead. American Documentation, 14(3):fmi, July

1963. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1963:Md

[Ano63j] Anonymous. Masthead. American Documentation, 14(4):fmi, Oc-

tober 1963. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1963:SA

[Ano63k] Anonymous. Silver anniversary 1937–1962. American Documentation, 14(2):94, April 1963. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1964:ADI

[Ano64a] Anonymous. American Documentation Institute. American Documentation, 15(4):274–288, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1964:BRBa

[Ano64b] Anonymous. Book review: 4/64-00 Information storage and retrieval: a survey. 1963. George W. Barnard and Carl Abbott. Biomedical Laboratory, 6570th Aerospace Medical Research Laboratories, Aerospace Medical Division, Air Force Systems Command (Technical Documentary Rept. No. AMRL-TDR-63-8), Wright-Patterson Air Force Base, Ohio. 15 pp. $.75. American Documentation, 15(2):155, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1964:BRBb

[Ano64c] Anonymous. Book review: 4/64-000 Methods of Information Handling. 1963. Charles P. Bourne. John Wiley & Sons, Inc. (Information Sciences Series), New York, 241 pp. $12.50. American Documentation, 15(2):155, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1964:BRBc

[Ano64d] Anonymous. Book review: 4/64-000 Thesaurus of Descriptors: A List of Keywords and Cross-References for Indexing and Retrieving the Lierature of Water Resources Development. 1963. (Tentative Edition) U.S. Department of the Interior, Bureau of Reclamation, Denver. 140 pp. American Documentation, 15(2):155–156, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1964:C

[Ano64e] Anonymous. Correction. American Documentation, 15(4):322, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1964:LN

[Ano64f] Anonymous. Literature notes. American Documentation, 15(1): 53–75, January 1964. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1964:Ma

[Ano64g] Anonymous. Masthead. American Documentation, 15(1):fmi, January 1964. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1964:Mb

[Ano64h] Anonymous. Masthead. American Documentation, 15(2):fmi, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1964:Mc

[Ano64i] Anonymous. Masthead. American Documentation, 15(3):fmi, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1964:Md

[Ano64j] Anonymous. Masthead. American Documentation, 15(4):fmi, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1965:BRBb

[Ano65a] Anonymous. Book review: 10/65-3R Some Fundamentals of Information Retrieval. 1965. John R. Sharp. Andre Deutsch (A Grafton Book), London. 224 pp., index. $8.75. American Documentation, 16(4):335, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1965:BRBa

[Ano65b] Anonymous. Book review: 7/65-2R Evaluation of the Impact of a Citation Index in Physics. 1964. Ben-Ami Lipetz. American Institute of Physics, New York. AIP/DRP CI-3–1964. American Documentation, 16(3):242, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1965:BCS

[Ano65c] Anonymous. Brief communications: Specifications for a superior microtext reading machine. American Documentation, 16(3):246– 247, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1965:Ea

[Ano65d] Anonymous. Editorial. American Documentation, 16(1):3, January

1965. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1965:Eb

[Ano65e] Anonymous. Errata. American Documentation, 16(4):335, October

1965. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1965:Ma

[Ano65f] Anonymous. Masthead. American Documentation, 16(1):fmi, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1965:Mb

[Ano65g] Anonymous. Masthead. American Documentation, 16(2):fmi, April

1965. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1965:Mc

[Ano65h] Anonymous. Masthead. American Documentation, 16(3):fmi, July

1965. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1965:Md

[Ano65i] Anonymous. Masthead. American Documentation, 16(4):fmi, Oc-

tober 1965. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1965:RCO

[Ano65j] Anonymous. Report for 1963–1964 of the Committee on Organization of Information to American Documentation Institute. American Documentation, 16(3):235–241, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1966:BRB

[Ano66a] Anonymous. Book review: 1/66-2R Author–Title Catalogue. Subject Catalogue. April 1965. Toronto University Library, Ontario New Universities Library Project. 2 vols. American Documentation, 17 (1):48, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1966:E

[Ano66b] Anonymous. Erratum. American Documentation, 17(1):46, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1966:Ma

[Ano66c] Anonymous. Masthead. American Documentation, 17(1):fmi, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1966:Mb

[Ano66d] Anonymous. Masthead. American Documentation, 17(2):fmi, April 1966. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1966:Mc

[Ano66e] Anonymous. Masthead. American Documentation, 17(3):fmi, July 1966. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1966:Md

[Ano66f] Anonymous. Masthead. American Documentation, 17(4):fmi, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1967:BCH

[Ano67a] Anonymous. Brief communications: Handier writing. American Documentation, 18(2):111, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1967:CP

[Ano67b] Anonymous. Call for papers. American Documentation, 18(2):ii, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1967:Ma

[Ano67c] Anonymous. Masthead. American Documentation, 18(1):fmi, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1967:Mb

[Ano67d] Anonymous. Masthead. American Documentation, 18(2):fmi, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1967:Mc

[Ano67e] Anonymous. Masthead. American Documentation, 18(3):fmi, July 1967. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1967:Md

[Ano67f] Anonymous. Masthead. American Documentation, 18(4):fmi, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1967:MPA

[Ano67g] Anonymous. Miscellaneous: Publications of the American Documentation Institute. American Documentation, 18(2):i, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:BRBd

[Ano68a] Anonymous. Book review: 1/67-1R Annual review of information science and technology. Volume 2. 1967. Carlos A. Cuadra, Editor. Interscience Publishers, a division of John Wiley & Sons. New York, 484 pp. American Documentation, 19(1):108–109, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:BRBc

[Ano68b] Anonymous. Book review: 1/67-2R Programming systems and languages. 1967. Saul Rosen, Editor. McGraw-Hill, New York, 734 pp. American Documentation, 19(1):107–108, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:BRBb

[Ano68c] Anonymous. Book review: 1/67-3R SNOBOL 3 primer — an introduction to the computer programming language. 1967. Allen Forte. The MIT Press, Cambridge, 107 pp. American Documentation, 19 (1):107, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:BRBa

[Ano68d] Anonymous. Book review: 1/67-4R Computers in humanistic research. 1967. Edmund A. Bowles, Editor. Prentice-Hall, Inc., Englewood Cliffs, New Jersey, 258 pp. American Documentation, 19 (1):107, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:BRBe

[Ano68e] Anonymous. Book review: 4/68–1R Automated language processing. 1967. Harold Borko (editor). Information Sciences Series. John Wiley and Sons, Inc., New York. 386 pp. $12.95. American Documentation, 19(2):208, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:BRBg

[Ano68f] Anonymous. Book review: 4/68–1R Thesaurus of ERIC descriptors. December 1967. U.S. Office of Education. Education Resources Information Center, Bureau of Research. Supt. of Documents Catalog No. FS 5.212.12031. Government Printing Office, Washington, D.C., 309 pp. $2.50. American Documentation, 19(4):418–419, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:BRBh

[Ano68g] Anonymous. Book review: 4/68–2R Computational analysis of present-day American English. 1967. Henry Kučera and W. Nelson Francis. Brown University Press, Providence, R.I., 424 p. $15.00. American Documentation, 19(4):419, October 1968. CODEN AM-DOA7. ISSN 0096-946X.

Anonymous:1968:BRBf

[Ano68h] Anonymous. Book review: 4/68–2R NASA thesaurus: Subject terms for indexing scientific and technical information. Preliminary edition. NASA SP-7030. December 1967. Scientific and Technical Information Division, Office of Technology Utilization. National Aeronautics and Space Administration, Washington, D.C. 3 volumes Sold by Superintendent of Documents: 1968 price, $8.50. American Documentation, 19(2):208–209, April 1968. CODEN AM-DOA7. ISSN 0096-946X.

Anonymous:1968:CP

[Ano68i] Anonymous. Call for papers. American Documentation, 19(1):ii, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:Ma

[Ano68j] Anonymous. Masthead. American Documentation, 19(1):fmi, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:Mb

[Ano68k] Anonymous. Masthead. American Documentation, 19(2):fmi, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:Mc

[Ano68l] Anonymous. Masthead. American Documentation, 19(3):fmi, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:Md

[Ano68m] Anonymous. Masthead. American Documentation, 19(4):fmi, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:MNPa

[Ano68n] Anonymous. Miscellaneous: New publications. American Documentation, 19(1):i, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:MNPb

[Ano68o] Anonymous. Miscellaneous: New Publications. American Documentation, 19(2):i, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:MPAa

[Ano68p] Anonymous. Miscellaneous: Publications of the American Society for Information Science. American Documentation, 19(1):i, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1968:MPAb

[Ano68q] Anonymous. Miscellaneous: Publications of the American Society for Information Science. American Documentation, 19(2):i, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1969:BRBa

[Ano69a] Anonymous. Book review: 1/69-1R Information retrieval thesaurus of education terms. 1968. Gordon C. Barhydt and Charles T. Schmidt, with the assistance of Kee T. Chang. Press of Case Western Reserve University, Cleveland, Ohio, 131 pp. $5.50. American Documentation, 20(1):100, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1969:BRBb

[Ano69b] Anonymous. Book review: 1/69-2R H. P. Luhn: Pioneer of Information Science; Selected Works. 1968. Edited by Claire K. Schultz. Spartan Press, New York, N.Y. 320 pp. $18.75. American Documentation, 20(1):100–101, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1969:BRBc

[Ano69c] Anonymous. Book review: 1/69-3R Information retrieval: Applications to Political Science. Kenneth Janda. Bobbs-Merrill Company, Indianapolis, Ind., 230 pp. (paper) $4.50. American Documentation, 20(1):101, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1969:BRBd

[Ano69d] Anonymous. Book review: 1/69-4R An evaluation of the New York State Library’s pilot program in the facsimile transmission of library materials. February 1968. Nelson Associates, Inc., 845 Third Avenue, New York 10022, 72 pp. Price not given. American Documentation, 20(1):101–102, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1969:BRBe

[Ano69e] Anonymous. Book review: 2/69-1 Information retrieval systems: Characteristics, Testing, and Evaluation. 1968. F. Wilfred Lancaster. John Wiley & Sons, New York (Information Sciences Series), 222 pp. $9.00. American Documentation, 20(2):173, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1969:BRBf

[Ano69f] Anonymous. Book review: 2/69-4 A comparative study of programming languages. 1967. Bryan Higman. American Elsevier Publishing Company, New York. (MacDonald Computer Monographs) 164 pp. $8.50. American Documentation, 20(2):175, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1969:LEN

[Ano69g] Anonymous. Letters to the Editor: New York Times thesaurus review. American Documentation, 20(3):281–282, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1969:Ma

[Ano69h] Anonymous. Masthead. American Documentation, 20(1):fmi, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1969:Mb

[Ano69i] Anonymous. Masthead. American Documentation, 20(2):fmi, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1969:Mc

[Ano69j] Anonymous. Masthead. American Documentation, 20(3):fmi, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Anonymous:1969:Md

[Ano69k] Anonymous. Masthead. American Documentation, 20(4):fmi, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Anziowar:1967:BRB

[Anz67] Boris R. Anziowar. Book review: 4/67-5R Information Retrieval with Special Reference to the Biomedical Sciences. 1966. Wesley Simonton and Charlene Mason, Editors. University of Minnesota, Minneapolis. 199 pp. American Documentation, 18(4):257, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Altmann:1969:TTM

[AR69] Berthold Altmann and Walter A. Riessler. Theory, testing, and mechanization of the ABC retrieval system. American Documentation, 20(1):6–15, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Aronson:1952:CLR

[Aro52] Eleanor J. Aronson. Cataloging in a large reports organization. American Documentation, 3(2):114–117, ???? 1952. CODEN AM-DOA7. ISSN 0096-946X.

Artandi:1961:LE

[Art61] Susan Artandi. Letter to the Editors. American Documentation, 12(3):239, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Artandi:1964:ABI

[Art64] Susan Artandi. Automatic book indexing by computer. American Documentation, 15(4):250–257, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Artandi:1965:BRB

[Art65] Susan Artandi. Book review: 10/65-1R Automatic Indexing: A State-of-the-Art Report. 1965. Mary Elizabeth Stevens. National Bureau of Standards, Washington, D.C. (NBS Monograph 91.) Superintendent of Documents, U.S. Government Printing Office. 223 pp. $1.50. American Documentation, 16(4):334, October 1965. CO-DEN AMDOA7. ISSN 0096-946X.

Artandi:1969:RIS

[Art69] Susan Artandi. The relevance of information science to libraryschool curricula. American Documentation, 20(4):337–338, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Adkinson:1967:LMR

[AS67] Burton W. Adkinson and Charles M. Stearns. Libraries and machines — A Review. American Documentation, 18(3):121–124, July 1967. CODEN AMDOA7. ISSN 0096-946X.

Ash:1966:BRB

[Ash66] Lee Ash. Book review: 1/66-1R A Directory of Information Resources in the United States: Physical Sciences, Biological Sciences, Engineering. 1965. (Issued by the Library of Congress, National Referral Center for Science and Technology). U.S. Government Printing Office, Washington, D.C. 352 pp. American Documentation, 17 (1):47, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Associates:1955:SRI

[Ass55] Mortimer Taube Associates. Storage and retrieval of information by means of the association of ideas. American Documentation, 6 (1):1–18, January 1955. CODEN AMDOA7. ISSN 0096-946X.

Atherton:1964:BCI

[Ath64] Pauline Atherton. Brief communications: Identification of research papers by the physical units appearing in the paper. American Documentation, 15(2):150, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Atherton:1965:LE

[Ath65a] P. Atherton. Letter to the Editor. American Documentation, 16(2): 126, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Atherton:1965:BRB

[Ath65b] Pauline Atherton. Book review: 4/65-2R Subject authority list. January 1964. American Petroleum Institute Information Retrieval Project, Central Abstracting Service, 1271 Avenue of the Americas, New York, New York 10020. Looseleaf. 245 pp. plus 74 pp. in Appendix and Hierarchy. American Documentation, 16(2):127–128, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Auer:1955:CLS

[Aue55] Erwin Auer. Chemical literature of Switzerland. American Documentation, 6(4):228–236, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Artandi:1969:EAG

[AW69] Susan Artandi and Edward H. Wolf. The effectiveness of automatically generated weights and links in mechanical indexing. American Documentation, 20(3):198–202, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Baker:1969:OUS

[Bak69] Norman R. Baker. Optimal user search sequences and implications for information systems operation. American Documentation, 20 (3):203–212, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Ball:1950:COIa

[Bal50a] Norman T. Ball. Committee on organization of information. American Documentation, 1(1):24–34, January 1950. CODEN AMDOA7. ISSN 0096-946X.

Ball:1950:COIb

[Bal50b] Norman T. Ball. Committee on organization of information. American Documentation, 1(2):100–109, April 1950. CODEN AMDOA7. ISSN 0096-946X.

Ball:1950:PAD

[Bal50c] Norman T. Ball. A program for the American Documentation Institute. American Documentation, 1(2):61–65, April 1950. CODEN AMDOA7. ISSN 0096-946X.

Ball:1955:SPM

[Bal55] Norman T. Ball. Searching patents by machine. American Documentation, 6(2):88–92, April 1955. CODEN AMDOA7. ISSN 0096-946X.

Ballou:1969:BRB

[Bal69] Hubbard W. Ballou. Book review: 3/69-2R Microphotography, photography and photofabrication at extreme resolution. 1968. G. W. W. Stevens. Wiley, Inc., N.Y. 2nd Ed. 510 pp. $25.00. American Documentation, 20(3):283–284, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Barrow:1953:EDR

[Bar53] W. J. Barrow. An evaluation of document restoration processes. American Documentation, 4(2):50–54, April 1953. CODEN AM-DOA7. ISSN 0096-946X.

Barden:1956:AST

[Bar56] William A. Barden. Armed services technical information agency. American Documentation, 7(3):188–200, July 1956. CODEN AM-DOA7. ISSN 0096-946X.

Batts:1967:BRB

[Bat67a] Natalie C. Batts. Book review: 2/67-1R Annual Review of Information Science and Technology. Volume 1. 1966. Carlos A. Cuadra, Editor. American Documentation Institute. Interscience Publishers, a division of John Wiley & Sons, New York, 389 pp. American Documentation, 18(2):115, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Batty:1967:DTM

[Bat67b] C. D. Batty. A documentation training model. American Documentation, 18(3):125–130, July 1967. CODEN AMDOA7. ISSN 0096-946X.

Bauer:1952:ERL

[Bau52] Charles K. Bauer. Exploiting report literature: Ways and means of reaching the consumer. American Documentation, 3(3):161–163, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Baum:1966:CST

[Bau66] Harry Baum. A clearinghouse for scientific and technical meetings: Organizational and operational problems. American Documentation, 17(1):28–32, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Brannon:1969:ALA

[BBJB69] Pam Barney Brannon, Donald F. Burnham, Richard M. James, and Lee Ann Bertram. Automated literature alerting system. American Documentation, 20(1):16–20, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Borko:1963:CAA

[BC63] Harold Borko and Seymour Chatman. Criteria for acceptable abstracts: a survey of abstracters’ instructions. American Documentation, 14(2):149–160, April 1963. CODEN AMDOA7. ISSN 0096-946X.

Brandhorst:1965:BCN

[BE65] W. T. Brandhorst and P. F. Eckert. Brief communications: NASA search system analysis sheet. American Documentation, 16(2):124– 126, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Beatty:1969:LES

[Bea69] William K. Beatty. Letters to the Editor: St. Ibusmae. American Documentation, 20(3):281, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Becourt:1963:HI

[Bec63] Daniel Becourt. Halo sur l’information. American Documentation, 14(4):283–288, October 1963. CODEN AMDOA7. ISSN 0096-946X.

Becker:1968:DED

[Bec68] A. M. Becker. Documentation and electronic data processing. American Documentation, 19(3):311–316, July 1968. CODEN AM-DOA7. ISSN 0096-946X.

Belzer:1969:CS

[Bel69a] Jack Belzer. The conference summary. American Documentation, 20(4):352–354, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Belzer:1969:CT

[Bel69b] Jack Belzer. The conference theme. American Documentation, 20 (4):329–330, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Belzer:1969:P

[Bel69c] Jack Belzer. Preface. American Documentation, 20(4):328, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Bennington:1952:IRL

[Ben52] John Bennington. The integration of report literature and journals. American Documentation, 3(3):149–152, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Benbrook:1957:DP

[Ben57] C. H. Benbrook. The diazotype process. American Documentation, 8(2):81–88, April 1957. CODEN AMDOA7. ISSN 0096-946X.

Bennelt:1969:LAI

[Ben69] John L. Bennelt. On-line access to information: NSF as an aid to the indexer/cataloger. American Documentation, 20(3):213–220, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Berthold:1954:IAL

[Ber54] Arthur B. Berthold. International aspects of librarianship. American Documentation, 5(4):252–253, October 1954. CODEN AM-DOA7. ISSN 0096-946X.

Bernier:1956:CIA

[Ber56a] Charles L. Bernier. Correlative indexes I. Alphabetical correlative indexes. American Documentation, 7(4):283–288, October 1956. CODEN AMDOA7. ISSN 0096-946X.

Bernier:1956:LI

[Ber56b] Charles L. Bernier. Language and indexes. American Documentation, 7(3):222–224, July 1956. CODEN AMDOA7. ISSN 0096-946X.

Bernier:1957:CIIa

[Ber57a] Charles L. Bernier. Correlative indexes II: Correlative trope indexes. American Documentation, 8(1):47–50, January 1957. CO-DEN AMDOA7. ISSN 0096-946X.

Bernier:1957:CIIc

[Ber57b] Charles L. Bernier. Correlative indexes IV: Correlative chemicalgroup indexes. American Documentation, 8(4):306–313, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Bernier:1958:CIV

[Ber58] Charles L. Bernier. Correlative indexes v. the blank sort. American Documentation, 9(1):32–41, January 1958. CODEN AMDOA7. ISSN 0096-946X.

Bernier:1960:CIV

[Ber60] Charles L. Bernier. Correlative indexes VI: Serendipity, suggestiveness, and display. American Documentation, 11(4):277–287, October 1960. CODEN AMDOA7. ISSN 0096-946X.

Bernier:1965:IPE

[Ber65] Charles L. Bernier. Indexing process evaluation. American Documentation, 16(4):323–328, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Bernier:1969:LE

[Ber69] Charles L. Bernier. Letter to the Editors. American Documentation, 20(4):387, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Bessinger:1961:CTO

[Bes61] J. B. Bessinger. Computer techniques for an old English concordance. American Documentation, 12(3):227–229, August 1961. CO-DEN AMDOA7. ISSN 0096-946X.

Betts:1959:RTL

[Bet59] Robert E. Betts. The resources of the technical libraries of Western Europe. American Documentation, 10(1):64–69, January 1959. CODEN AMDOA7. ISSN 0096-946X.

Beyerly:1955:DSY

[Bey55] Elizabeth Beyerly. Documentation of science in Yugoslavia and Hungary. American Documentation, 6(3):151–161, July 1955. CO-DEN AMDOA7. ISSN 0096-946X.

Beyerly:1957:DWS

[Bey57] Elizabeth Beyerly. The documentation of Western science in the Soviet Union. American Documentation, 8(2):98–102, April 1957. CODEN AMDOA7. ISSN 0096-946X.

Beyerly:1960:HST

[Bey60a] Elizabeth Beyerly. The Hungarian State Technical Library (OMK): Its role in coordinated Hungarian scientific documentation. American Documentation, 11(4):296–304, October 1960. CODEN AM-DOA7. ISSN 0096-946X.

Beyerly:1960:SBP

[Bey60b] Elizabeth Beyerly. Soviet bibliographical projects achievements in the USSR, assessment in the USA. American Documentation, 11 (1):44–54, January 1960. CODEN AMDOA7. ISSN 0096-946X.

Beyerly:1962:HSP

[Bey62] Elizabeth Beyerly. Hungarian scientific publications: a survey of documentation trends. American Documentation, 13(2):153–158, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Brownson:1957:AFT

[BF57] Helen L. Brownson and Bernard M. Fry. Alberto F. Thompson, 1907–1957. American Documentation, 8(4):243–245, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Barnard:1958:IFR

[BF58] G. A. Barnard and Louis Fein. An information filing and retrieval system for the engineering and management records of a large-scale computer development project. American Documentation, 9(3): 208–213, August 1958. CODEN AMDOA7. ISSN 0096-946X.

Bourne:1964:CAS

[BF64] Charles P. Bourne and Donald F. Ford. Cost analysis and simulation procedures for the evaluation of large information systems. American Documentation, 15(2):142–149, April 1964. CODEN AM-DOA7. ISSN 0096-946X.

Bush:1956:AUS

[BGM56] G. C. Bush, H. P. Galliher, and P. M. Morse. Attendance and use of the science library at M.I.T. American Documentation, 7(2): 87–109, April 1956. CODEN AMDOA7. ISSN 0096-946X.

Bar-Hillel:1951:PSR

[BH51] Yehoshua Bar-Hillel. The present state of research on mechanical translation. American Documentation, 2(4):229–237, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Bar-Hillel:1957:LRR

[BH57a] Yehoshua Bar-Hillel. A logician’s reaction to recent theorizing on information search systems. American Documentation, 8(2):103–113, April 1957. CODEN AMDOA7. ISSN 0096-946X. See comments [Ano57a].

Bernier:1957:CIIb

[BH57b] Charles L. Bernier and Karl F. Heumann. Correlative indexes III. Semantic relations among semantemes the technical thesaurus. American Documentation, 8(3):211–220, July 1957. CODEN AM-DOA7. ISSN 0096-946X.

Bar-Hillel:1963:IRA

[BH63] Yehoshua Bar-Hillel. Is information retrieval approaching a crisis? American Documentation, 14(2):95–98, April 1963. CODEN AM-DOA7. ISSN 0096-946X.

Buchanan:1967:AAH

[BH67] J. R. Buchanan and F. C. Hutton. Analysis and automated handling of technical information at the nuclear safety information center. American Documentation, 18(4):235–241, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Baker:1967:FSD

[BHR67] C. W. Baker, C. R. Haefele, and W. A. Reckhow. A film system for the duplication of Termatrex cards. American Documentation, 18 (2):55–58, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Bingaman:1963:SSL

[Bin63] Ronald Bingaman. The scientist and scientific literature — a symposium of the central Ohio chapter, ADI, May 18, 1962. American Documentation, 14(2):161–162, April 1963. CODEN AMDOA7. ISSN 0096-946X.

Bishop:1953:IAD

[Bis53] Charles Bishop. An integrated approach to the documentation problem. American Documentation, 4(2):54–65, April 1953. CO-DEN AMDOA7. ISSN 0096-946X.

Bishop:1954:CVS

[Bis54] Charles Bishop. CODEN versus SIGILS. American Documentation, 5(1):28–29, January 1954. CODEN AMDOA7. ISSN 0096-946X.

Bishop:1956:MPS

[Bis56] Charles Bishop. The microcard production of single journal articles. American Documentation, 7(1):33–35, January 1956. CODEN AMDOA7. ISSN 0096-946X.

Bishop:1957:UCS

[Bis57] Charles Bishop. Use of the CODEN system by the individual research scientist. American Documentation, 8(3):221–226, July 1957. CODEN AMDOA7. ISSN 0096-946X.

Bishop:1959:STC

[Bis59] David Bishop. Science thesis control in Europe and America. American Documentation, 10(1):51–58, January 1959. CODEN AM-DOA7. ISSN 0096-946X.

Bishop:1961:PPU

[Bis61] Charles Bishop. Problems in the production and utilization of microfiche. American Documentation, 12(1):53–55, January 1961. CO-DEN AMDOA7. ISSN 0096-946X.

Burton:1960:HLS

[BK60] R. E. Burton and R. W. Kebler. The “half-life” of some scientific and technical literatures. American Documentation, 11(1):18–22, January 1960. CODEN AMDOA7. ISSN 0096-946X.

Barnett:1963:CEV

[BK63] Michael P. Barnett and K. L. Kelley. Computer editing of verbal texts. Part I. The ESI system. American Documentation, 14(2): 99–108, April 1963. CODEN AMDOA7. ISSN 0096-946X.

Barnett:1962:CGP

[BKB62] M. P. Barnett, K. L. Kelley, and M. J. Bailey. Computer generation of photocomposing control tapes. Part I. Preparation of Flexowriter source material. American Documentation, 13(1):58–65, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Bobrow:1966:CPS

[BKRL66] Daniel G. Bobrow, R. Y. Kain, Bertram Raphael, and J. C. R. Licklider. A computer-program system to facilitate the study of technical documents. American Documentation, 17(4):186–189, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Bernier:1955:DMI

[BL55] Charles L. Bernier and Cecil C. Langham. Dictating machines as indexing aids. American Documentation, 6(4):237–238, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Bourne:1965:LNb

[BL65] Charles P. Bourne and Burton E. Lamkin. Literature notes. American Documentation, 16(2):131–160, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Bloomfield:1957:ECI

[Blo57] Masse Bloomfield. Evaluation of coordinate indexing at the naval ordnance test station. American Documentation, 8(1):22–25, January 1957. CODEN AMDOA7. ISSN 0096-946X.

Blum:1950:ECL

[Blu50] John M. Blum. Editors’ camera: “The letters of Theodore Roosevelt”. American Documentation, 1(4):181–184, October 1950. CODEN AMDOA7. ISSN 0096-946X.

Bluemle:1959:REI

[Blu59] Andrew Bluemle. The role of the editor in interdisciplinary research. American Documentation, 10(2):105–107, April 1959. CODEN AM-DOA7. ISSN 0096-946X.

Barnett:1964:CGP

[BML64] M. P. Barnett, D. J. Moss, and D. A. Luce. Computer generation of photocomposing control tapes part II the PC6 system. American Documentation, 15(2):115–120, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Bishop:1965:PPS

[BMR65] David Bishop, Arnold L. Milner, and Fred W. Roper. Publication patterns of scientific serials. American Documentation, 16(2):113– 121, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Baker:1968:USS

[BN68] Norman R. Baker and Richard E. Nance. The use of simulation in studying information storage and retrieval systems. American Documentation, 19(4):363–370, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Borden:1969:TVC

[BN69] George A. Borden and William F. Nelson. Toward a viable classification scheme: Some theoretical considerations. American Documentation, 20(4):298–301, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Bohnert:1955:TMO

[Boh55] Lea M. Bohnert. Two methods of organizing technical information for search. American Documentation, 6(3):134–151, July 1955. CODEN AMDOA7. ISSN 0096-946X.

Bohnert:1962:BRB

[Boh62] Lea M. Bohnert. Book review: Information handling and science information: a selected bibliography, 1957–1961. By Paul C. Janaske, Editor, Biological Sciences Communication Project, The American Institute of Biological Sciences, in cooperation with The American University, Center of Technology and Administration, School of Government and Public Administration, Washington, D.C. 1962. unp. $2. American Institute of Biological Sciences, 2000 P St. N.W., Washington 6, D.C. American Documentation, 13(4):435–436, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Bohnert:1964:BRR

[Boh64] Lea Bohnert. Book review: R10/1 Information Storage and Retrieval: Tools, Elements, Theories. 1963. Joseph Becker and Robert M. Hayes. John Wiley & Sons, Inc. (Information Sciences Series), New York. $12.00. American Documentation, 15(4):307–308, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Bohnert:1965:BRB

[Boh65] Lea Bohnert. Book review: 10/65-2R The Future of the Research Library. 1964. Verner W. Clapp. University of Illinois Press, Urbana. 114 pp. $4.50. American Documentation, 16(4):334–335, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Boni:1951:M

[Bon51] Albert Boni. Microprint. American Documentation, 2(3):150–152, August 1951. CODEN AMDOA7. ISSN 0096-946X.

Bonn:1962:TEI

[Bon62] George S. Bonn. Training and education for information work. American Documentation, 13(3):301–312, July 1962. CODEN AM-DOA7. ISSN 0096-946X.

Booth:1950:EC

[Boo50] Robert Edmond Booth. Editor’s corner. American Documentation, 1(2):121–125, April 1950. CODEN AMDOA7. ISSN 0096-946X.

Booth:1960:MCI

[Boo60] A. D. Booth. More on coordinate indexing, a review. American Documentation, 11(4):316, October 1960. CODEN AMDOA7. ISSN 0096-946X.

Born:1950:NPE

[Bor50] Lester K. Born. A national plan for extensive microfilm operations. American Documentation, 1(2):66–75, April 1950. CODEN AM-DOA7. ISSN 0096-946X.

Born:1951:PMO

[Bor51] Lester K. Born. Planning for microfilm operations. American Documentation, 2(1):1–5, January 1951. CODEN AMDOA7. ISSN 0096-946X.

Born:1954:SAE

[Bor54] Lester K. Born. Some animadversions on European standards and practices. American Documentation, 5(3):160–162, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Born:1955:PEA

[Bor55] Lester K. Born. A program for enriching American library resources. American Documentation, 6(4):219–225, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Born:1956:LM

[Bor56] Lester K. Born. The literature of microreproduction, 1950–1955. American Documentation, 7(3):167–187, July 1956. CODEN AM-DOA7. ISSN 0096-946X.

Bornstein:1961:PRE

[Bor61] Harry Bornstein. A paradigm for a retrieval effectiveness experiment. American Documentation, 12(4):254–259, October 1961. CO-DEN AMDOA7. ISSN 0096-946X.

Borko:1964:MRS

[Bor64] Harold Borko. Measuring the reliability of subject classification by men and machines. American Documentation, 15(4):268–273, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Borkowski:1967:ESA

[Bor67] Casimir Borkowski. An experimental system for automatic identification of personal names and personal titles in newspaper texts. American Documentation, 18(3):131–138, July 1967. CODEN AM-DOA7. ISSN 0096-946X.

Borko:1968:ISW

[Bor68] H. Borko. Information science: What is it? American Documentation, 19(1):3–5, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Bourne:1961:HDP

[Bou61] Charles P. Bourne. The historical development and present state-of-the-art of mechanized information retrieval systems. American Documentation, 12(2):108–110, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Bourne:1962:WTJ

[Bou62] Charles P. Bourne. The world’s technical journal literature: an estimate of volume, origin, language, field, indexing, and abstracting. American Documentation, 13(2):159–168, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Bourne:1964:LNa

[Bou64a] Charles P. Bourne. Literature notes. American Documentation, 15 (2):157–175, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Bourne:1964:LNb

[Bou64b] Charles P. Bourne. Literature notes. American Documentation, 15 (3):231–246, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Bourne:1964:LNc

[Bou64c] Charles P. Bourne. Literature notes. American Documentation, 15 (4):309–322, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Bourne:1965:LNa

[Bou65] Charles P. Bourne. Literature notes. American Documentation, 16 (1):41–64, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Bowers:1954:IMC

[Bow54] Ford W. Bowers. Identification of microcards containing classified material. American Documentation, 5(1):29, January 1954. CO-DEN AMDOA7. ISSN 0096-946X.

Berry:1954:MLS

[BPK54] M. M. Berry, J. W. Perry, and Allen Kent. Machine literature searching IV. Collection of terminology. American Documentation, 5(2):95–100, April 1954. CODEN AMDOA7. ISSN 0096-946X.

Bray:1951:RAC

[Bra51] Robert S. Bray. Rapid announcement of the contents of scientific journals. American Documentation, 2(4):213–216, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Bray:1952:WMR

[Bra52] Robert S. Bray. Ways and means of reaching the consumer. American Documentation, 3(3):152–154, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Bradshaw:1962:WSL

[Bra62] P. D. Bradshaw. The Walnut System: a large capacity document storage and retrieval system. American Documentation, 13(3):270– 275, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Brandhorst:1964:CAC

[Bra64] W. T. Brandhorst. Corporate author cataloging and the technical report literature. American Documentation, 15(1):35–47, January 1964. CODEN AMDOA7. ISSN 0096-946X.

Brandhorst:1966:BCS

[Bra66] W. T. Brandhorst. Brief communications: Simulation of boolean logic constraints through the use of term weights. American Documentation, 17(3):145–146, July 1966. CODEN AMDOA7. ISSN 0096-946X.

Brandhorst:1967:BRB

[Bra67] W. T. Brandhorst. Book review: 4/67-6R Coordinate Indexing. 1966. John C. Costello, Jr., Graduate School of Library Service, Rutgers, The State University (Rutgers Series on Systems for the Intellectual Organization of Information Vol. VII), Edited by Dr. Susan Artandi. Supported by the National Science Foundation. New Brunswick, New Jersey. 218 pp. American Documentation, 18(4): 257–258, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Bradbury:1968:UMH

[Bra68] Ray Bradbury. Unthinking man and his thinking machines. American Documentation, 19(4):371–374, October 1968. CODEN AM-DOA7. ISSN 0096-946X.

Brenner:1964:LE

[Bre64] Everett H. Brenner. Letter to the Editor. American Documentation, 15(3):230, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Brownson:1952:RRI

[Bro52] Helen L. Brownson. Recommendations and results of international conferences on scientific information and bibliographic services. American Documentation, 3(1):29–55, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Brownson:1953:CTBa

[Bro53a] Helen L. Brownson. Contributions toward a bibliography of documentation. American Documentation, 4(1):27–34, January 1953. CODEN AMDOA7. ISSN 0096-946X.

Brownson:1953:CTBb

[Bro53b] Helen L. Brownson. Contributions toward a bibliography of documentation. American Documentation, 4(2):69–81, April 1953. CO-DEN AMDOA7. ISSN 0096-946X.

Brownson:1953:LNa

[Bro53c] Helen L. Brownson. Literature notes. American Documentation, 4 (3):126–135, August 1953. CODEN AMDOA7. ISSN 0096-946X.

Brownson:1953:LNb

[Bro53d] Helen L. Brownson. Literature notes. American Documentation, 4 (4):174–184, October 1953. CODEN AMDOA7. ISSN 0096-946X.

Brownson:1954:LN

[Bro54] Helen L. Brownson. Literature notes. American Documentation, 5 (4):254–267, October 1954. CODEN AMDOA7. ISSN 0096-946X.

Brownson:1955:ADL

[Bro55] Helen L. Brownson. Abstracts of documentation literature. American Documentation, 6(2):63–67, April 1955. CODEN AMDOA7. ISSN 0096-946X.

Brownson:1956:IMD

[Bro56] Helen L. Brownson. International meetings on documentation. American Documentation, 7(1):49–52, January 1956. CODEN AM-DOA7. ISSN 0096-946X.

Brockway:1959:CIU

[Bro59] Duncan Brockway. Coordinate indexing at the university of new Hampshire library. American Documentation, 10(3):228–231, July 1959. CODEN AMDOA7. ISSN 0096-946X.

Brookes:1968:BCS

[Bro68] B. C. Brookes. Brief communications: The stability of ranks of index terms. American Documentation, 19(1):101–102, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Bryant:1955:FMY

[Bry55] Douglas W. Bryant. Fid meets in Yugoslavia. American Documentation, 6(2):102–103, April 1955. CODEN AMDOA7. ISSN 0096-946X.

Bracken:1968:SPT

[BS68] Marilyn C. Bracken and Charles W. Shilling. Survey of practical training in information science. American Documentation, 19(2): 113–119, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Blodgett:1969:HHD

[BS69] John H. Blodgett and Claire K. Schultz. Herman Hollerith: data processing pioneer. American Documentation, 20(3):221–226, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Buchanan:1958:SSS

[Buc58] Bruce Buchanan. A system of social significance. American Documentation, 9(2):114–142, April 1958. CODEN AMDOA7. ISSN 0096-946X.

Buck:1961:SIS

[Buc61] R. Creighton Buck. Studies in information storage and retrieval: On the use of gödel indices in coding. American Documentation, 12 (3):165–171, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Buist:1962:TBI

[Bui62] Eleanor Buist. Training in bibliography and information techniques for the Soviet student. American Documentation, 13(4):397–400, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Burton:1959:CAEa

[Bur59a] Robert E. Burton. Citations in American engineering journals I. Chemical engineering. American Documentation, 10(1):70–73, January 1959. CODEN AMDOA7. ISSN 0096-946X.

Burton:1959:CAEb

[Bur59b] Robert E. Burton. Citations in American engineering journals, II. Mechanical engineering. American Documentation, 10(2):135–137, April 1959. CODEN AMDOA7. ISSN 0096-946X.

Burton:1959:CAEc

[Bur59c] Robert E. Burton. Citations in American engineering journals III. Metallurgical engineering. American Documentation, 10(3):209– 213, July 1959. CODEN AMDOA7. ISSN 0096-946X.

Campion:1956:EBR

[Cam56] Eleanor Este Campion. Existing bibliographical resources — dead end, detour, or freeway? American Documentation, 7(2):83–86, April 1956. CODEN AMDOA7. ISSN 0096-946X.

Canter:1952:FRA

[Can52] Louis Canter. Factors restricting the availability of government research reports. American Documentation, 3(3):133–138, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Carruthers:1950:TME

[Car50] Ralph H. Carruthers. Titling of microfilm editions. American Documentation, 1(4):190–193, October 1950. CODEN AMDOA7. ISSN 0096-946X.

Carlson:1952:CTR

[Car52] Irving G. Carlson. Cataloging of technical reports in a military research library. American Documentation, 3(2):119–122, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Carter:1954:UBP

[Car54] Edward Carter. Unesco’s bibliographical programmes. American Documentation, 5(1):1–7, January 1954. CODEN AMDOA7. ISSN 0096-946X.

Carruthers:1956:SMF

[Car56] R. H. Carruthers. A Simple method for figuring microfilm footage on a slide rule. American Documentation, 7(1):44–45, January 1956. CODEN AMDOA7. ISSN 0096-946X.

Carlson:1963:HGE

[Car63] William H. Carlson. The holy grail evades the search. American Documentation, 14(3):207–212, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Caras:1968:CSS

[Car68] Gus J. Caras. Computer simulation of a small information system. American Documentation, 19(2):120–122, April 1968. CODEN AM-DOA7. ISSN 0096-946X.

Carroll:1969:ASV

[Car69] Kathleen H. Carroll. An analytical survey of virology literature reported in two announcement journals. American Documentation, 20(3):234–237, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Casey:1961:NN

[Cas61] Robert S. Casey. News notes. American Documentation, 12(1): 75–77, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Clavier:1955:SCR

[CB55] Henri Clavier and Normat T. Ball. Some considerations regarding the mechanism of thought. American Documentation, 6(4):215–219, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Carpenter:1959:SCP

[CBF59] R. A. Carpenter, C. C. Bolze, and L. D. Findley. A system for the correlation of physical properties and structural characteristics of chemical compounds with their commercial uses. American Documentation, 10(2):138–143, April 1959. CODEN AMDOA7. ISSN 0096-946X.

Curtice:1965:LE

[CDM65] Robert M. Curtice, Lauren B. Doyle, and Richard J. McClure. Letters to the Editors. American Documentation, 16(3):248, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Cass:1962:CCS

[CG62] Saul Cass and C. D. Gull. Computers and common sense: the myth of thinking machines. By Mortimer Taube. New York, Columbia University Press, 1961. 136 p. American Documentation, 13(2): 223–224, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Chapman:1950:LCI

[Cha50] Edward A. Chapman. Low-cost institutional photocopying service. American Documentation, 1(1):35–41, January 1950. CODEN AM-DOA7. ISSN 0096-946X.

Chapin:1963:BRB

[Cha63] Richard E. Chapin. Book review: Advanced data processing in the university library. By Louis A. Schultheiss, Don S. Culbertson, and Edward M. Heiliger. Scarecrow Press, New York. 1962. 388 p. $10.00. American Documentation, 14(1):83, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Cheydleur:1963:SRF

[Che63] Benjamin F. Cheydleur. SHIEF: a realizable form of associative memory. American Documentation, 14(1):56–67, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Cheydleur:1965:ISL

[Che65] Benjamin F. Cheydleur. Information science and liberal education. American Documentation, 16(3):171–177, July 1965. CODEN AM-DOA7. ISSN 0096-946X.

Climenson:1961:ASA

[CHJ61] W. D. Climenson, N. H. Hardwick, and S. N. Jacobson. Automatic syntax analysis in machine indexing and abstracting. American Documentation, 12(3):178–183, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Chomsky:1957:LSL

[Cho57] Noam Chomsky. Logical structures in language. American Documentation, 8(4):284–291, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Clapp:1950:SCN

[Cla50] Verner W. Clapp. Shorter communications: News and technical notes. American Documentation, 1(2):110–120, April 1950. CO-DEN AMDOA7. ISSN 0096-946X.

Clapp:1952:SCN

[Cla52] Verner W. Clapp. Subject controls — nature and levels of controls. American Documentation, 3(1):11–15, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Clapp:1953:PSC

[Cla53] Verner W. Clapp. The problem of specialized communication in modern society. American Documentation, 4(3):83–90, August 1953. CODEN AMDOA7. ISSN 0096-946X.

Clapp:1955:LEL

[Cla55] Verner W. Clapp. Letter to the Editor: The Library of Congress. Washington 25, D.C. American Documentation, 6(4):271, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Clapp:1956:UPL

[Cla56] Verner W. Clapp. UNESCO’s program: Libraries, bibliographic services, documentation, a catalog of principal activities. American Documentation, 7(2):127–134, April 1956. CODEN AMDOA7. ISSN 0096-946X.

Clapp:1961:ISR

[Cla61] Verner W. Clapp. “information storage and retrieval” and the problems of libraries. American Documentation, 12(3):224–226, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Carlson:1963:LE

[CLA^{+}63a] Gary Carlson, F. W. Lancaster, Hugh C. Atkinson, Ralph R. Shaw, Ronald E. Wyllys, Richard H. Orr, and T. R. Savage. Letters to the Editor. American Documentation, 14(4):328–331, October 1963. CODEN AMDOA7. ISSN 0096-946X.

Clapp:1963:RPS

[Cla63b] Verner W. Clapp. Research in problems of scientific information — retrospect and prospect. American Documentation, 14(1):1–9, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Clark:1967:BRB

[Cla67] Haydee Garcia Clark. Book review: 2/67-5R A Plan for Indexing the Periodical Literature of Nursing. Report of a Study of the Need for Bibliographic Control of the Scholarly Record of Nursing. 1966. Vern N. Pings. American Nurses Foundation, New York. 202 pp. American Documentation, 18(2):117–118, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Cleverdon:1963:BRB

[Cle63] C. W. Cleverdon. Book review: A comparative study of three systems of information retrieval. By Norman D. Stevens. Rutgers University Press, New Brunswick, N. J. 1950. American Documentation, 14(1):82, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Clark:1962:AXM

[CM62] H. E. Clark and G. R. Mott. Abstract: A xerographic method for display or storage of transmitted information. American Documentation, 13(1):118, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Coblans:1956:UNT

[Cob56] Herbert Coblans. United Nations technical assistance and documentation. American Documentation, 7(4):289–293, October 1956. CODEN AMDOA7. ISSN 0096-946X.

Coblans:1965:EGE

[Cob65] Herbert Coblans. Editorial: Guest editorial. American Documentation, 16(2):67, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Cohn:1955:EGP

[Coh55] Ernst M. Cohn. Editing graphs for publication. American Documentation, 6(2):77–86, April 1955. CODEN AMDOA7. ISSN 0096-946X.

Cohn:1959:TDB

[Coh59] Gunther Cohn. Technical data on border-punched cards. American Documentation, 10(2):116–121, April 1959. CODEN AMDOA7. ISSN 0096-946X.

Conrad:1951:TFS

[Con51a] G. Miles Conrad. Thermo-fax, Stenafax and Xerox. American Documentation, 2(2):95–97, April 1951. CODEN AMDOA7. ISSN 0096-946X.

Conrad:1951:UNM

[Con51b] G. Miles Conrad. User needs in a microfacsimile reader. American Documentation, 2(4):201–204, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Conrad:1952:NMR

[Con52] G. Miles Conrad. New methods of reproduction, publication, and the design of technical reports. American Documentation, 3(3): 168–175, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Cooper:1966:PS

[Coo66] Daniel I. Cooper. Pop science. American Documentation, 17(2): 53–56, April 1966. CODEN AMDOA7. ISSN 0096-946X.

Cooper:1968:ESL

[Coo68] William S. Cooper. Expected search length: a single measure of retrieval effectiveness based on the weak ordering action of retrieval systems. American Documentation, 19(1):30–41, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Cooper:1969:OPI

[Coo69] William S. Cooper. Opinion paper: Is interindexer consistency a hobgoblin? American Documentation, 20(3):268–278, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Cortelyou:1954:SFD

[Cor54] Ethaline Cortelyou. Some fundamentals of designing tables of data. American Documentation, 5(3):155–160, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Costello:1961:SRC

[Cos61a] J. C. Costello, Jr. Storage and retrieval of chemical research and patent information by links and roles in Du Pont. American Documentation, 12(2):111–120, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Costello:1961:UIP

[Cos61b] J. C. Costello, Jr. Uniterm indexing principles, problems and solutions. American Documentation, 12(1):20–26, January 1961. CO-DEN AMDOA7. ISSN 0096-946X.

Costello:1961:SSO

[Cos61c] John C. Costello, Jr. Some solutions to operational problems in concept coordination. American Documentation, 12(3):191–197, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Costello:1962:CRI

[Cos62] John C. Costello, Jr. Computer requirements for inverted coordinate indexes. American Documentation, 13(4):414–419, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Cottrell:1968:ECS

[Cot68] William B. Cottrell. Evaluation and compression of scientific and technical information at the nuclear safety information center.

American Documentation, 19(4):375–380, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Carroll:1969:CSK

[CR69] John M. Carroll and Robert Roeloffs. Computer selection of keywords using word-frequency analysis. American Documentation, 20 (3):227–233, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Crooks:1956:FSA

[Cro56] Elliott Crooks. Facsimile systems as an aid to research. American Documentation, 7(1):40–43, January 1956. CODEN AMDOA7. ISSN 0096-946X.

Chapin:1962:MSS

[CS62] Richard E. Chapin and Charles W. Shilling. A model for the study of scientific communications. American Documentation, 13(4):410– 413, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Cuadra:1964:IKC

[Cua64] Carlos A. Cuadra. Identifying key contributions to information science. American Documentation, 15(4):289–295, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Cuadra:1965:CAM

[Cua65] Carlos A. Cuadra. Comments on the 1964 ADI meeting. American Documentation, 16(2):105–112, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Cueto:1968:UIL

[Cue68] Manuel R. Cueto. Using information in the life insurance business world. American Documentation, 19(3):265–268, July 1968. CO-DEN AMDOA7. ISSN 0096-946X.

Cummins:1954:CTB

[Cum54] R. L. Cummins. Changeable type bars for typewriter composition. American Documentation, 5(1):13–18, January 1954. CODEN AM-DOA7. ISSN 0096-946X.

Daily:1956:NSR

[Dai56] Jay E. Daily. A Notation for subject retrieval files. American Documentation, 7(3):210–214, July 1956. CODEN AMDOA7. ISSN 0096-946X.

Daily:1957:SHT

[Dai57] Jay E. Daily. Subject headings and the theory of classification. American Documentation, 8(4):269–274, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Dain:1964:BRB

[Dai64] Phyllis Dain. Book review: 4/64-000 William Frederick Poole and the modern library movement. 1963. William Landrum Williamson. Columbia University Press (Columbia University Studies in Library Service, No. 13), New York. 203 pp. $6.00. American Documentation, 15(2):154, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Damerau:1965:EAI

[Dam65] Fred J. Damerau. An experiment in automatic indexing. American Documentation, 16(4):283–289, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Darling:1966:BRB

[Dar66] Louise Darling. Book review: 10/66–3R Mental Health Book Review Index. Vol. 10. Whole No. 15. 1965. Compiled by the Editorial Committee and Contributing Librarians. Council on Research in Bibliography, Inc., New York. 78 pp. American Documentation, 17 (4):220–221, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Davis:1951:YEA

[Dav51a] Watson Davis. 15 years experience with auxiliary publication. American Documentation, 2(2):87–89, April 1951. CODEN AM-DOA7. ISSN 0096-946X.

Davis:1951:DAP

[Dav51b] Watson Davis. Developments in auxiliary publication. American Documentation, 2(1):7–11, January 1951. CODEN AMDOA7. ISSN 0096-946X.

Davis:1952:ROA

[Dav52] Watson Davis. A report on the operation of auxiliary publication service. American Documentation, 3(4):225, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

David:1956:CHS

[Dav56] Charles W. David. The conservation of historical source material. American Documentation, 7(2):76–82, April 1956. CODEN AM-DOA7. ISSN 0096-946X.

Davis:1958:TIS

[Dav58] Barbara M. Davis. Technical information service in the Cabot Research Laboratories. American Documentation, 9(1):23–27, January 1958. CODEN AMDOA7. ISSN 0096-946X.

Davis:1967:SPP

[Dav67] Charles H. Davis. SDI: a program for public libraries. American Documentation, 18(3):139–145, July 1967. CODEN AMDOA7. ISSN 0096-946X.

Davis:1968:BCI

[Dav68] Charlks H. Davis. Brief communications: Integrating vocabularies with a classification scheme. American Documentation, 19(1):101, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Davis:1969:BCB

[Dav69] Charles H. Davis. Brief communications: The binary search algorithm. American Documentation, 20(2):167, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Dawkins:1968:IAC

[Daw68] Young P. Dawkins. The individual in an age of change. American Documentation, 19(3):269, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Dale:1965:SCE

[DD65] A. G. Dale and N. Dale. Some clumping experiments for associative document retrieval. American Documentation, 16(1):5–9, January 1965. CODEN AMDOA7. ISSN 0096-946X.

DeHaas:1958:M

[De 58] W. De Haas. The microfiche. American Documentation, 9(2):99– 106, April 1958. CODEN AMDOA7. ISSN 0096-946X.

Delucia:1964:IAE

[Del64] Al Delucia. Index-abstract evaluation and design. American Documentation, 15(2):121–125, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Dennis:1950:ACL

[Den50] Faustine Dennis. American Council of Learned Societies, British manuscripts project. The collection in the Library of Congress. American Documentation, 1(3):130–132, August 1950. CODEN AMDOA7. ISSN 0096-946X.

Dewey:1959:PCC

[Dew59] Harry Dewey. Punched card catalogs — theory and technique. American Documentation, 10(1):36–50, January 1959. CODEN AMDOA7. ISSN 0096-946X.

Davison:1957:SCG

[DG57] W. H. T. Davison and M. Gordon. Sorting for chemical groups using Gordon–Kendall–Davison ciphers. American Documentation, 8(3):202–210, July 1957. CODEN AMDOA7. ISSN 0096-946X.

Dickison:1966:SFM

[Dic66] R. R. Dickison. The scholar and the future of microfilm. American Documentation, 17(4):178–179, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Dietrich:1962:MRP

[Die62] E. V. Dietrich. Machine retrieval of pharmacodynamic data. American Documentation, 13(1):112–114, January 1962. CODEN AM-DOA7. ISSN 0096-946X.

Donohue:1966:CAA

[DK66] Joseph C. Donohue and N. E. Karioth. Coming of age in academe — Information science at 21. American Documentation, 17(3):117– 119, July 1966. CODEN AMDOA7. ISSN 0096-946X.

Debons:1969:FCE

[DO69] Anthony Debons and Klaus Otten. Foundations of a concept for an education program in information science. American Documentation, 20(4):346–351, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Donley:1957:LEE

[Don57] Albert M. Donley, Jr. Library ecology and electronics. American Documentation, 8(4):300–305, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Donohue:1966:LSI

[Don66] Joseph C. Donohue. Librarianship and the science of information. American Documentation, 17(3):120–123, July 1966. CODEN AM-DOA7. ISSN 0096-946X.

Doyle:1962:IAA

[Doy62] Lauren B. Doyle. Indexing and abstracting by association. American Documentation, 13(4):378–390, October 1962. CODEN AM-DOA7. ISSN 0096-946X.

Drake:1954:AAR

[Dra54] E. J. Drake. Access to Australian research. American Documentation, 5(4):245–246, October 1954. CODEN AMDOA7. ISSN 0096-946X.

Drew:1966:LTL

[DSTW66] D. L. Drew, R. K. Summit, R. I. Tanaka, and R. B. Whiteley. An on-line technical library reference retrieval system. American Documentation, 17(1):3–7, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Dubester:1962:RAD

[Dub62] Henry Dubester. The role of the American Documentation Institute in international documentation. American Documentation, 13(1): 115–117, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Dyer:1956:RPU

[Dye56] J. Raymond Dyer. Report on the practical utilization of its recorded knowledge by the legal profession. American Documentation, 7(3): 201–209, July 1956. CODEN AMDOA7. ISSN 0096-946X.

Dyke:1959:FMO

[Dyk59] H. Gordon Dyke. A figure-of-merit ordering system for a search output. American Documentation, 10(1):85–86, January 1959. CO-DEN AMDOA7. ISSN 0096-946X.

Eastin:1951:PVD

[Eas51] Roy Eastin. The point of view of the division of public documents. American Documentation, 2(3):160–162, August 1951. CODEN AMDOA7. ISSN 0096-946X.

Ebersole:1966:OMN

[Ebe66] J. L. Ebersole. An operating model of a national information system. American Documentation, 17(1):33–40, January 1966. CO-DEN AMDOA7. ISSN 0096-946X.

Edge:1957:SIR

[EFB57] Eleanor B. Edge, Norman G. Fisher, and Lucy A. Bannister. System for indexing research reports using a punched card machine.

American Documentation, 8(4):275–283, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Egan:1951:BDS

[Ega51] Margaret E. Egan. Bibliographic development in the social sciences. American Documentation, 2(1):11–18, January 1951. CODEN AM-DOA7. ISSN 0096-946X.

Egan:1953:USD

[Ega53] Margaret E. Egan. The use of social data by business, finance, and industry. American Documentation, 4(4):147–154, October 1953. CODEN AMDOA7. ISSN 0096-946X.

Elias:1964:Ea

[EK64] Arthur W. Elias and John F. Kennedy. Editorial. American Documentation, 15(1):3, January 1964. CODEN AMDOA7. ISSN 0096-946X.

Elias:1964:Eb

[Eli64a] Arthur W. Elias. Editorial. American Documentation, 15(2):81, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Elias:1964:Ec

[Eli64b] Arthur W. Elias. Editorial. American Documentation, 15(3):178, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Elias:1965:Eb

[Eli65a] A. W. Elias. Editorial. American Documentation, 16(3):162, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Elias:1965:Ea

[Eli65b] Arthur W. Elias. Editorial. American Documentation, 16(1):1–3, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Elias:1966:Ea

[Eli66a] A. W. Elias. Editorial. American Documentation, 17(1):1–2, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Elias:1966:Eb

[Eli66b] Arthur W. Elias. Editorial. American Documentation, 17(3):115, July 1966. CODEN AMDOA7. ISSN 0096-946X.

Elias:1967:Eb

[Eli67a] A. W. Elias. Editorial. American Documentation, 18(2):54, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Elias:1967:Ea

[Eli67b] Arthur W. Elias. Editorial. American Documentation, 18(1):1–2, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Elias:1968:Ea

[Eli68a] A. W. Elias. Editorial. American Documentation, 19(1):1–2, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Elias:1968:Eb

[Eli68b] A. W. Elias. Editorial. American Documentation, 19(3):212, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Elias:1969:GE

[Eli69a] A. Elias. Guest editorial. American Documentation, 20(3):185, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Elias:1969:E

[Eli69b] A. W. Elias. Editorial. American Documentation, 20(4):297, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Elias:1969:ELA

[Eli69c] A. W. Elias. Editorial looking at AD –1968. American Documentation, 20(1):1–5, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Ellsworth:1951:NHM

[Ell51] R. S. Ellsworth. New horizons with microfilm. American Documentation, 2(4):221–228, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Ellas:1967:E

[Ell67] A. W. Ellas. Editorial. American Documentation, 18(3):120, July 1967. CODEN AMDOA7. ISSN 0096-946X.

Elias:1964:F

[EM64] Arthur W. Elias and James D. Mack. Farewell. American Documentation, 15(1):1–2, January 1964. CODEN AMDOA7. ISSN 0096-946X.

Ember:1965:BRB

[Emb65] George Ember. Book review: 1/65-3R A Contribution to the Early History of Documentation (Adatok es felismerések a szakirodalmi dokumentació tortenetenek kezdeteihez) [Data and findings in the history of the beginnings of bibliography]. 1964. Dr. Ivan Polzovics. Országos Müszaki Könyvtar es Dokumentációs Központ [Technical Library and Documentation Center], Budapest. 80 pp. American Documentation, 16(1):39–40, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Engelbart:1961:SCI

[Eng61] Douglas C. Engelbart. Special considerations of the individual as a user, generator, and retriever of information. American Documentation, 12(2):121–125, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Engstrom:1962:ED

[Eng62] Howard T. Engstrom. Electronics and documentation. American Documentation, 13(1):119–122, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Eller:1968:TDD

[EP68] James L. Eller and Robert L. Panek. Thesaurus development for a decentralized information network. American Documentation, 19 (3):213–220, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Evans:1950:AT

[Eva50] Luther H. Evans. ADI’s tasks. American Documentation, 1(3): 127–130, August 1950. CODEN AMDOA7. ISSN 0096-946X.

Evans:1951:AMA

[Eva51] Luther H. Evans. The 1951 annual meeting. American Documentation Institute. American Documentation, 2(2):69–80, April 1951. CODEN AMDOA7. ISSN 0096-946X.

Evans:1952:RUC

[Eva52] Elma Evans. Reference use of the collection and its catalogs. American Documentation, 3(3):146–149, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Evans:1956:DET

[Eva56] Melbourne G. Evans. Duplication in experimental and theoretical disciplines and its implication for modern documentation. American Documentation, 7(1):46–48, January 1956. CODEN AMDOA7. ISSN 0096-946X.

Evans:1961:EEPa

[Eva61a] Luther H. Evans. Editorial: The Editor’s page. American Documentation, 12(1):1–2, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Evans:1961:EPa

[Eva61b] Luther H. Evans. The Editor’s pages. American Documentation, 12(3):240, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Evans:1961:EPb

[Eva61c] Luther H. Evans. The Editor’s pages. American Documentation, 12(4):306, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Evans:1961:EEPb

[Eva61d] Luther H. Evans. Editors: The Editor’s page. American Documentation, 12(2):158, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Evans:1966:BRB

[Eva66] Luther H. Evans. Book review: 10/66–5R Library Science Today: Ranganathan Festschrift Vol. II. 1965. Edited by Kaula Prithvi Nath. Ranganathan Series in Library Science, 14. Asia Publishing House, New York, 832 pp. $30.00. American Documentation, 17(4):222, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Evans:1968:TMO

[Eva68] Luther H. Evans. Traditional methods of organizing and storing information. American Documentation, 19(3):271–272, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Fagerhaugh:1952:AIR

[Fag52] Kenneth H. Fagerhaugh. Availability of industrial research reports. American Documentation, 3(3):143–145, ???? 1952. CODEN AM-DOA7. ISSN 0096-946X.

Fairthorne:1956:PR

[Fai56] R. A. Fairthorne. The patterns of retrieval. American Documentation, 7(2):65–70, April 1956. CODEN AMDOA7. ISSN 0096-946X.

Fairthorne:1958:DC

[Fai58] R. A. Fairthorne. Delegation of classification. American Documentation, 9(3):159–164, August 1958. CODEN AMDOA7. ISSN 0096-946X.

Forman:1968:BRB

[FC68] Sidney Forman and Ruby Lee Collins. Book review: 4/68–3R Documentation in education. Arvid J. and Mary A. Burke. Revision of Carter Alexander and Arvid J. Burke’s How to Locate Educational Information and Data (4th edition, revised) 1967. Teachers College Press, Teachers College, Columbia University. 413 pp. $7.50. American Documentation, 19(2):209, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Feinler:1965:AST

[FCHB65] E. J. Feinler, C. J. Cook, O. Heinz, and C. P. Bourne. Attitudes of scientists toward a specialized information center. American Documentation, 16(4):329–333, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Feichtinger:1963:MMA

[Fei63] H. George Feichtinger. Maxterm-miniterm: an alpha numeric notation for significant document storage. American Documentation, 14(1):35–38, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Fellows:1957:CBS

[Fel57] Erwin W. Fellows. Current bibliographic services in the social sciences. American Documentation, 8(3):153–167, July 1957. CODEN AMDOA7. ISSN 0096-946X.

Fels:1963:EPI

[Fel63] E. M. Fels. Evaluation of the performance of an informationretrieval system by modified mooers plan. American Documentation, 14(1):28–34, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Feldman:1964:PIP

[Fel64] Alfred Feldman. A proposed improvement in the printing of chemical structures, which results in their complete computer codes. American Documentation, 15(3):205–209, July 1964. CODEN AM-DOA7. ISSN 0096-946X.

Firschein:1967:LE

[FFBB67] O. Firschein, M. Fischler, W. T. Brandhorst, and W. T. Brandhorst. Letters to the Editors. American Documentation, 18(1):49, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Field:1956:ADE

[Fie56] Oliver T. Field. An application of the direct entry principle in indexing. American Documentation, 7(3):225–228, July 1956. CODEN AMDOA7. ISSN 0096-946X.

Fisher:1950:NFR

[Fis50] Loran L. Fisher. Notes on film records. American Documentation, 1(4):184–190, October 1950. CODEN AMDOA7. ISSN 0096-946X.

Fisher:1959:ARS

[Fis59] Eva Lou Fisher. Abbreviations of Russian scientific serial publications. American Documentation, 10(3):192–208, July 1959. CODEN AMDOA7. ISSN 0096-946X.

Fischer:1966:KIC

[Fis66] Marguerite Fischer. The KWIC index concept: a retrospective view. American Documentation, 17(2):57–70, April 1966. CODEN AM-DOA7. ISSN 0096-946X.

Fayollat:1969:BCC

[FL69] James Fayollat and Don Luck. Brief communications: Computer based serials control system, Biomedical Library, UCLA. American Documentation, 20(4):385, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Flanigan:1969:BCO

[Fla69] David E. Flanigan. Brief communications: An open-ended information retrieval system, including selective data collection. American Documentation, 20(2):169–171, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Flood:1967:AQA

[Flo67] Barbara Flood. Analysis of questions addressed to a medical reference retrieval system: Comparison of question and system terminologies. American Documentation, 18(4):216–227, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Foster:1958:RAS

[Fos58] Laurence S. Foster. Revision of the ASM–SLA metallurgical literature classification. American Documentation, 9(1):13–19, January 1958. CODEN AMDOA7. ISSN 0096-946X.

Fairthorns:1966:LE

[FP66] Robert Fairthorns and J. Poyen. Letters to the Editors. American Documentation, 17(2):107, April 1966. CODEN AMDOA7. ISSN 0096-946X.

Francis:1952:CRI

[Fra52] Edith S. Francis. Cataloging of reports in an industrial library. American Documentation, 3(2):117–119, ???? 1952. CODEN AM-DOA7. ISSN 0096-946X.

Franklin:1966:BCR

[Fra66] David Franklin. Brief communications: Reversible punctuation Russian transliteration. American Documentation, 17(3):142–145, July 1966. CODEN AMDOA7. ISSN 0096-946X.

Flemer:1957:FMC

[FS57] H. F. Flemer and Atherton Seidell. The FS microfilm camera. American Documentation, 8(2):127–128, April 1957. CODEN AM-DOA7. ISSN 0096-946X.

Fukudome:1967:JIC

[Fuk67] Takao Fukudome. The Japan Information Center of Science and Technology (JICST): Its organization and function. American Documentation, 18(3):146–152, July 1967. CODEN AMDOA7. ISSN 0096-946X.

Garfield:1954:PRM

[Gar54] Eugene Garfield. Preliminary report on the mechanical analysis of information by use of the 101 statistical punched card machine. American Documentation, 5(1):7–12, January 1954. CODEN AM-DOA7. ISSN 0096-946X.

Garfield:1955:PPI

[Gar55] Eugene Garfield. The preparation of printed indexes by automatic punched-card techniques. American Documentation, 6(2):68–76, April 1955. CODEN AMDOA7. ISSN 0096-946X.

Garfield:1963:CIS

[Gar63] Eugene Garfield. Citation indexes in sociological and historical research. American Documentation, 14(4):289–291, October 1963. CODEN AMDOA7. ISSN 0096-946X.

Garvin:1965:BRB

[Gar65a] Paul L. Garvin. Book review: 4/65-1R Language and information: Selected Essays on Their Theory and Application. 1964. Yehoshua Bar-Hillel. Addison-Wesley Publishing Company, Inc., Reading, Massachusetts, Palo Alto, London, and The Jerusalem Academic Press Ltd., Jerusalem, Israel. $12.50. American Documentation, 16 (2):127, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Garvin:1965:ISM

[Gar65b] Paul L. Garvin. An informal survey of modern linguistics. American Documentation, 16(4):291–298, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Garwig:1969:CB

[Gar69] Paul L. Garwig. Charles Babbage (1792–1871). American Documentation, 20(4):320–324, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Gifford:1969:UUC

[GB69] Carolyn Gifford and George J. Baumanis. On understanding user choices: Textual correlates of relevance judgments. American Documentation, 20(1):21–26, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Gottschalk:1963:WCS

[GD63] Charles M. Gottschalk and Winifred F. Desmond. Worldwide census of scientific and technical serials. American Documentation, 14 (3):188–194, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Gottesman:1957:MDA

[GG57] Jerome Gottesman and Edward Gottesman. Machines, documentation, and automatic coding. American Documentation, 8(2):129– 133, April 1957. CODEN AMDOA7. ISSN 0096-946X.

Garvey:1964:SOF

[GG64] William D. Garvey and Belver C. Griffith. The structure, objectives, and findings of a study of scientific information exchange in psychology. American Documentation, 15(4):258–267, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Gilbert:1958:OPC

[Gil58] Paul T. Gilbert. An optimal punch card code for general files. American Documentation, 9(2):84–98, April 1958. CODEN AMDOA7. ISSN 0096-946X.

Giuliano:1969:RIS

[Giu69] Vincent E. Giuliano. The relationship of information science to librarianship — problems and scientific training. American Documentation, 20(4):344–345, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Gray:1951:AAN

[Gra51] Dwight E. Gray. The AEC, ASTIA, NACA CATALOG CARD. American Documentation, 2(4):219–221, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Graf:1952:RSI

[Gra52a] Dorothy Graf. Reports service to industry. American Documentation, 3(3):145–146, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Gray:1952:PEM

[Gra52b] Dwight E. Gray. Practical experience in microfacsimile publication. American Documentation, 3(1):58–61, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Gray:1953:OSU

[Gra53] Dwight E. Gray. Organizing and servicing unpublished reports. American Documentation, 4(3):103–115, August 1953. CODEN AMDOA7. ISSN 0096-946X.

Graziano:1968:TD

[Gra68] Eugene E. Graziano. On a theory of documentation. American Documentation, 19(1):85–89, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Greengrass:1967:BRB

[Gre67] Alan R. Greengrass. Book review: 2/67-6R Indexing and Classification; A Selected and Annotated Bibliography. A joint project of the Nuclear Science Division and Documentation Division of the Special Libraries Association. May 14, 1966. Compiled and edited by Winifred F. Desmond and Lester A. Barrer. Oak Ridge National Laboratory Library, Oak Ridge, Tennessee. 256 pp. (on microfiche). American Documentation, 18(2):118, April 1967. CODEN AM-DOA7. ISSN 0096-946X.

Groesbeck:1951:DSR

[Gro51] Joseph A. Groesbeck. Discovering sources of research material. American Documentation, 2(2):80–83, April 1951. CODEN AM-DOA7. ISSN 0096-946X.

Groos:1967:BCB

[Gro67] Ole V. Groos. Brief communications: Bradford’s Law and the Keenan–Atherton data. American Documentation, 18(1):46, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Groos:1968:BCR

[Gro68] Ole V. Groos. Brief communications: Relative importance of articles-cited versus titles-cited in frequency counts. American Documentation, 19(1):102–103, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Garfield:1963:NFE

[GS63] E. Garfield and I. H. Sher. New factors in the evaluation of scientific literature through citation indexing. American Documentation, 14 (3):195–201, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Gull:1951:CCT

[Gul51] C. D. Gull. The cumulative catalog technique at the Library of Congress. American Documentation, 2(3):131–141, August 1951. CODEN AMDOA7. ISSN 0096-946X.

Gull:1956:PUS

[Gul56a] C. D. Gull. Posting for the uniterm system of coordinate indexing. American Documentation, 7(1):9–21, January 1956. CODEN AMDOA7. ISSN 0096-946X.

Gull:1956:SYW

[Gul56b] C. D. Gull. Seven years of work on the organization of materials in the special library. American Documentation, 7(4):320–329, October 1956. CODEN AMDOA7. ISSN 0096-946X.

Gull:1961:ARP

[Gul61] C. D. Gull. Annual report of the President of the American Documentation Institute October 1959—October 1960. American Documentation, 12(1):3–5, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Gull:1967:LE

[Gul67] C. D. Gull. Letters to the Editors. American Documentation, 18 (4):252–253, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Goffman:1964:UML

[GVB64] W. Goffman, J. Verhoeff, and Jack Belzer. Use of meta-language in information retrieval systems. American Documentation, 15(1): 14–22, January 1964. CODEN AMDOA7. ISSN 0096-946X.

Hoshovsky:1965:TNT

[HA65] A. G. Hoshovsky and H. H. Album. Toward a national technical information system. American Documentation, 16(4):313–322, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Hammer:1968:PCB

[Ham68] Donald P. Hammer. Problems in the conversion of bibliographical data — A keypunching experiment. American Documentation, 19 (1):12–17, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Harris:1951:OUN

[Har51] Alfred G. Harris. Organizing the United Nations documents collection. American Documentation, 2(3):141–149, August 1951. CO-DEN AMDOA7. ISSN 0096-946X.

Harper:1954:UDC

[Har54] Shirley F. Harper. The universal decimal classification. American Documentation, 5(4):195–213, October 1954. CODEN AMDOA7. ISSN 0096-946X.

Harvey:1963:RBI

[Har63] John F. Harvey. Reviews: 10/63-1 Science Information Personnel: the new profession of information combining science, librarianship, and foreign language. 1961. Leonard Cohan and Kenneth Craven. Modern Language Association, New York. May be obtained from Science Information, P.O. Box 624, Radio City Station, New York 19, N. Y., for $1.50. American Documentation, 14(4):312, October 1963. CODEN AMDOA7. ISSN 0096-946X.

Harris:1966:BRB

[Har66] Jessica L. Harris. Book review: Catalog of Books Plus a Complete Catalog of Reserve Books. 1965. Washington University School of Medicine Library, St. Louis, Mo. 267 pp. American Documentation, 17(1):48–49, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Harrison:1967:BRB

[Har67a] Michael A. Harrison. Book review: 4/67-1R Theory of Self-Reproducing Automata. 1966. John von Neumann. Arthur W. Burks, Editor. University of Illinois Press. American Documentation, 18(4):254, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Harvey:1967:BRBb

[Har67b] John Harvey. Book review: 2/67-4R Education and Training of Information Specialists in the U.S.A. May 1966. Marilyn C. Bracken and Charles W. Shilling. Biological Communication Project, Washington, D.C. 70 pp. American Documentation, 18(2):117, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Harvey:1967:BRBa

[Har67c] John F. Harvey. Book review: 1/67-4R The Polities of Research. 1966. Richard J. Barber. The Public Affairs Press, Washington, D.C. 167 pp. American Documentation, 18(1):51, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Harris:1968:BRBa

[Har68a] Jessica L. Harris. Book review: 3/68–1R An introduction to colon classification. 1966 C. D. Batty. Archon Books and Clive Bingley, Hamden, Connecticut. 90 pp. $3.50. American Documentation, 19 (3):357, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Harris:1968:BRBb

[Har68b] Jessica L. Harris. Book review: 3/68–3R The New York Times thesaurus of descriptors; A guide for organizing, cataloguing, indexing and searching collections of information on current events. 1968. New York Times Company, New York. 1103 pp. in 2 vols. (Evennumbered pages blank). Tab-indexed. American Documentation, 19 (3):358–359, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Harris:1968:BRBc

[Har68c] Jessica L. Harris. Book review: 4/68–3R Organization and handling of bibliographic records by computer. 1967. Nigel S. M. Cox and Michael W. Grose, Editors. Archon Books, Hamden, Connecticut, 187 pages, $12.00. American Documentation, 19(4):419–420, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Harris:1968:OPT

[Har68d] Jessica L. Harris. Offset printing from typescript as a substitute for microfilming of dissertations. American Documentation, 19(1): 60–65, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Harris:1969:BRB

[Har69] Jessica L. Harris. Book review: 2/69-2 Automation in libraries. 1968. Richard T. Kimber. Pergamon Press, New York, 140 pp. 45. American Documentation, 20(2):173–174, April 1969. CODEN AM-DOA7. ISSN 0096-946X.

Haykin:1957:WNW

[Hay57] David Judson Haykin. What is needed by way of special classifications? American Documentation, 8(1):51–52, January 1957. CODEN AMDOA7. ISSN 0096-946X.

Hayne:1960:IIR

[Hay60] Robert L. Hayne. Information in industry, a review. American Documentation, 11(1):55, January 1960. CODEN AMDOA7. ISSN 0096-946X.

Hayes:1966:BRBa

[Hay66a] Robert Hayes. Book review: 1/66-3R The Education of Science Information Personnel — 1964. 1965. A. J. Goldwyn and A. M. Rees, Eds. Western Reserve University, Cleveland. 115 pp. American Documentation, 17(1):49, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Hayes:1966:BRBb

[Hay66b] Robert M. Hayes. Book review: 10/66-6R Use of Mechanized Methods in Documentation Work. 1966. Herbert Coblans, ASLIB, London. 89 pp. American Documentation, 17(4):222–223, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Hayes:1967:BRB

[Hay67] Robert M. Hayes. Book review: 3/67-1R Data-Retrieval by Computer: A Critical Survey. 1966. Asa Kasher. The Hebrew University, Jerusalem, Israel. 76 pp. American Documentation, 18(3):187–189, July 1967. CODEN AMDOA7. ISSN 0096-946X.

Hayes:1969:EIS

[Hay69] R. M. Hayes. Education in information science. American Documentation, 20(4):362–365, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Heald:1952:RLN

[Hea52] Lcdr J. Heston Heald. The reports librarian and national security. American Documentation, 3(3):138–143, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Heilprin:1961:CEA

[Hei61a] L. B. Heilprin. Communication engineering approach to microforms. American Documentation, 12(3):213–218, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Heilprin:1961:IPA

[Hei61b] L. B. Heilprin. On the information problem ahead. American Documentation, 12(1):6–14, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Heilprin:1965:SMS

[Hei65] L. Heilprin. Special messages: Special message. American Documentation, 16(2):66, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Heilprin:1968:TFC

[Hei68] Laurence B. Heilprin. Technology and the future of the copyright principle. American Documentation, 19(1):6–11, January 1968. CO-DEN AMDOA7. ISSN 0096-946X.

Helmkamp:1969:MCA

[Hel69] John G. Helmkamp. Managerial cost accounting for a technical information center. American Documentation, 20(2):111–118, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Herner:1952:SHU

[Her52] Saul Herner. The storage and housing of unpublished research reports. American Documentation, 3(2):110–114, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Herner:1957:SCI

[Her57] Saul Herner. Studies in coordinate indexing, volume III. A review. American Documentation, 8(1):56–58, January 1957. CODEN AM-DOA7. ISSN 0096-946X.

Herner:1962:MOI

[Her62] Saul Herner. Methods of organizing information for storage and searching. American Documentation, 13(1):3–14, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Heumann:1956:NND

[Heu56a] Karl F. Heumann. Notes on negative data. American Documentation, 7(1):36–39, January 1956. CODEN AMDOA7. ISSN 0096-946X.

Heumann:1956:PCB

[Heu56b] Karl F. Heumann. Punched cards in the biological sciences, a review. American Documentation, 7(4):330, October 1956. CODEN AMDOA7. ISSN 0096-946X.

Heilprin:1965:ABI

[HG65] Laurence B. Heilprin and Frederick L. Goodman. Analogy between information retrieval and education. American Documentation, 16 (3):163–169, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Harbison:1953:IRC

[HH53] Frederick H. Harbison and Shirley F. Harper. Industrial relations — a case study of specialized communication involving several groups. American Documentation, 4(4):155–162, October 1953. CODEN AMDOA7. ISSN 0096-946X.

Hosemann:1960:ASP

[HH60] Hans Hosemann and Jürgen Haller. Application of a special punched card system to coordinated clinical statistics of malignant tumors of the female genitalia. American Documentation, 11(2): 170–172, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Hersey:1967:CUD

[HH67] David F. Hersey and William Hammond. Computer usage in the development of a water resources thesaurus. American Documentation, 18(4):209–215, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Harris:1968:LE

[HHK68] Jessica L. Harris, Karl F. Heumann, and William T. Knox. Letters to the Editors. American Documentation, 19(3):355–356, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Hillman:1964:NR

[Hil64a] Donald J. Hillman. The notion of relevance (I). American Documentation, 15(1):26–34, January 1964. CODEN AMDOA7. ISSN 0096-946X.

Hillman:1964:TMR

[Hil64b] Donald J. Hillman. Two models for retrieval system design. American Documentation, 15(3):217–225, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Hilf:1966:BRB

[Hil66] Jack W. Hilf. Book review: 10/66-4R Mining, Minerals, and Geosciences. Volume 2 of Guides to Information Sources in Science and Technology. 1965. Stuart R. Kaplan. Interscience Publishers, a division of John Wiley & Sons, New York City, 599 pp. American Documentation, 17(4):221–222, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Hillman:1969:HWP

[Hil69] Donald J. Hillman. How to win the pentathlon — unifying theories of information science. American Documentation, 20(4):335–336, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Hines:1961:PCS

[Hin61] Theodore C. Hines. The productivity of certain scholars. American Documentation, 12(4):294–295, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Hines:1969:BRB

[Hin69] Theodore C. Hines. Book review: 4/69-4R Data study. 1968. J. L. Jolley. McGraw-Hill Book Company, New York (World University Library). 225 pp. index. $4.95. American Documentation, 20(4): 389–390, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Hirsch:1950:ULM

[Hir50] Rudolf Hirsch. Union list of microfilms. American Documentation, 1(2):88–90, April 1950. CODEN AMDOA7. ISSN 0096-946X.

Hirayama:1962:TRC

[Hir62] Kenzo Hirayama. Time required, cost, and personnel for documentation. American Documentation, 13(3):313–319, July 1962. CO-DEN AMDOA7. ISSN 0096-946X.

Hubbell:1963:PUS

[HMC63] Allan F. Hubbell, Foster E. Mohrhardt, and Verner W. Clapp. The program of the United States National Committee for FID. Program committee of USNCFID. American Documentation, 14(1):78–81, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Hodgers:1968:ENR

[Hod68] Robert W. Hodgers, Jr. An emerging national resource — networks for information transportation, processing, and inquiry. American Documentation, 19(3):273, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Hoffman:1952:MDC

[Hof52] Erwin F. Hoffman. Multi-dimensional classification and the transcription of cancer literature to punched cards. American Documentation, 3(1):61–70, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Hoffer:1962:IRS

[Hof62] Joe R. Hoffer. Information retrieval in social welfare: Experiences with an edge-notched information retrieval system. American Documentation, 13(2):169–175, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Hoffer:1967:RNS

[Hof67] Joe R. Hoffer. The relationship of natural and social sciences to social problems and the contribution of the information scientist to their solutions. American Documentation, 18(4):228–234, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Holmes:1950:LCP

[Hol50] Donald C. Holmes. The library of congress, photoduplication service. American Documentation, 1(2):84–87, April 1950. CODEN AMDOA7. ISSN 0096-946X.

Holmes:1951:MME

[Hol51] Donald C. Holmes. Mechanical methods of enlargement print production and related processes in use at the library of congress. American Documentation, 2(4):216–219, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Holm:1962:SSE

[Hol62] B. E. Holm. Searching strategies and equipment. American Documentation, 13(1):31–42, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Holst:1966:BCM

[Hol66] W. Holst. Brief communications: Mechanical translation by coordinate indexing. American Documentation, 17(3):140–141, July 1966. CODEN AMDOA7. ISSN 0096-946X.

Hooker:1952:RLR

[Hoo52] Ruth H. Hooker. Reports librarian: Requirements and opportunities. American Documentation, 3(3):163–165, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Howerton:1965:BRB

[How65] Paul Howerton. Book review: 4/65-4R Lexicon of archive terminology (French, English, German, Spanish, Italian, Dutch). 1964. Jean Herbert, ed. Elsevier Publishing Co., New York. 83 pp. American Documentation, 16(2):130, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Hoyt:1969:AIS

[Hoy69] Ronald H. Hoyt. An analysis of information science programs. American Documentation, 20(4):358–361, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Holm:1961:DTT

[HR61] B. E. Holm and L. E. Rasmussen. Development of a technical thesaurus. American Documentation, 12(3):184–190, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Holt:1968:GPL

[HS68] Charles C. Holt and William E. Schrank. Growth of the professional literature in economics and other fields, and some implications. American Documentation, 19(1):18–26, January 1968. CO-DEN AMDOA7. ISSN 0096-946X.

Heilprin:1968:LE

[HSM^{+}68] Taurence M. Heilprin, R. F. Schmidt, Eugene Miller, William Hammond, Stafan Rosenborg, Johnie Sullivan, John O’Connor, Simon M. Newman, and H. B. Thompson. Letters to the Editors. American Documentation, 19(4):415–417, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Huebner:1951:PR

[Hue51] William C. Huebner. The Photronic reproducer. American Documentation, 2(4):238–240, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Hurwitz:1969:BCS

[Hur69] Frances I. Hurwitz. Brief communications: A study of indexer consistency. American Documentation, 20(1):92–94, January 1969. CO-DEN AMDOA7. ISSN 0096-946X.

Hutchisson:1955:CIO

[Hut55] Elmer Hutchisson. A continuing index to our stockpile of knowledge. American Documentation, 6(4):211–215, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Houston:1964:DTU

[HW64] Nona Houston and Eugene Wall. The distribution of term usage in manipulative indexes. American Documentation, 15(2):105–114, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Hyslop:1953:ICM

[Hys53] Marjorie R. Hyslop. International cooperation in metallurgical literature classification report of Interim Committee on Literature Classification Metals Section, Special Libraries Association. American Documentation, 4(1):1–3, January 1953. CODEN AMDOA7. ISSN 0096-946X.

Hyslop:1961:MLS

[Hys61] Marjorie R. Hyslop. Machine literature searching — from experiment to experience. American Documentation, 12(1):49–52, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Hyslop:1963:CST

[Hys63] Marjorie R. Hyslop. A compatibility study of two information systems. American Documentation, 14(4):292–298, October 1963. CO-DEN AMDOA7. ISSN 0096-946X.

Iker:1967:BCS

[Ike67] Howard P. Iker. Brief communications: Solution of Boolean Equations Through Use of Term Weights to the Base Two. American Documentation, 18(1):47, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Irvine:1951:SPM

[Irv51] Dallas Irvine. Storage problems and micro-reproduction. American Documentation, 2(2):84–86, April 1951. CODEN AMDOA7. ISSN 0096-946X.

Jackson:1952:AST

[Jac52] Eugene B. Jackson. Acquisitions: Sources and techniques. American Documentation, 3(2):94–100, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Jahoda:1964:TDI

[Jah64] G. Jahoda. A technique for determining index requirements. American Documentation, 15(2):82–85, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Jahoda:1969:IR

[Jah69] G. Jahoda. Introductory remarks. American Documentation, 20 (4):331, October 1969. CODEN AMDOA7. ISSN 0096-946X.

James:1959:PMS

[Jam59] Peter James. A photo-magnetic system for document and information retrieval. American Documentation, 10(4):286–295, October 1959. CODEN AMDOA7. ISSN 0096-946X.

Jaster:1962:ND

[Jas62] Josephine J. Jaster. A note on descriptors. American Documentation, 13(4):433–434, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Jones:1967:FCT

[JC67] Paul E. Jones and Robert M. Curtice. A framework for comparing term association measures. American Documentation, 18(3):153– 161, July 1967. CODEN AMDOA7. ISSN 0096-946X.

Jahoda:1968:UST

[JC68] G. Jahoda and Mary Culnan. Unanswered science and technology reference questions. American Documentation, 19(1):95–100, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Jensen:1965:BCN

[Jen65] Raymond A. Jensen. Brief communications: NFSAIS plans for coordination. American Documentation, 16(3):245, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Jenkins:1966:BCM

[JH66] J. Edward Jenkins and William T. Herzog. Brief communications: A matrix for evaluating information system operation. American Documentation, 17(2):103–106, April 1966. CODEN AMDOA7. ISSN 0096-946X.

Jahoda:1966:CUP

[JHG66] G. Jahoda, Ronald D. Hutchins, and Robert R. Galford. Characteristics and use of personal indexes maintained by scientists and engineers in one university. American Documentation, 17(2):71–75, April 1966. CODEN AMDOA7. ISSN 0096-946X.

Joyce:1958:TAI

[JN58] T. Joyce and R. M. Needham. The thesaurus approach to information retrieval. American Documentation, 9(3):192–197, August 1958. CODEN AMDOA7. ISSN 0096-946X.

Johnson:1962:PSI

[Joh62] H. Thayne Johnson. A polydimensional scheme for information retrieval. American Documentation, 13(1):90–92, January 1962. CO-DEN AMDOA7. ISSN 0096-946X.

Jones:1954:CTC

[Jon54] Graham Jones. The colon theory comes of age. American Documentation, 5(4):213–217, October 1954. CODEN AMDOA7. ISSN 0096-946X.

Jonker:1960:TIP

[Jon60] Frederick Jonker. The Termatrex inverted “punched card” systems the history of their development, their operation and use. American Documentation, 11(4):305–315, October 1960. CODEN AMDOA7. ISSN 0096-946X.

Jones:1961:AA

[Jon61] Richard E. Jones. Actifilm in action. American Documentation, 12 (3):222–223, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Jonker:1963:NTL

[Jon63] Frederick Jonker. The new “Termatrex” line of I.R. Systems — the “Minimatrex” line of I.R. Systems. American Documentation, 14 (4):276–282, October 1963. CODEN AMDOA7. ISSN 0096-946X.

Jordan:1959:HCP

[Jor59] Philip D. Jordan. The historian and the contemporary problem of bibliographic techniques. American Documentation, 10(4):267–269, October 1959. CODEN AMDOA7. ISSN 0096-946X.

Jordan:1968:FCS

[Jor68] John R. Jordan. A framework for comparing SDI systems. American Documentation, 19(3):221–222, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Jahoda:1969:CKT

[JS69] G. Jahoda and Mary Lou Stursa. A comparison of a keyword from title index with a single access point per document alphabetic subject index. American Documentation, 20(4):377–380, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Jenks:1961:BRB

[JT61a] I. H. Jenks and Mortimer Taube. Book review: Proceedings of the international conference on scientific information. 1959. National Academy of Sciences — National Research Council. Washington, D.C. 2 vols., 1635 pp. American Documentation, 12(1):61–62, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Jones:1961:NDB

[JT61b] Robert M. Jones and Mortimer Taube. Notes on the distinction between character recognition machines and perceiving machines. American Documentation, 12(4):292–293, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Kaplan:1965:NCB

[Kap65] Norman Kaplan. The norms of citation behavior: Prolegomena to the footnote. American Documentation, 16(3):179–184, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Komiske:1968:ACF

[KB68] Patrick T. Komiske and Albert A. Braunstein. Automated control of a fast-turnaround contractor — Subcontractor document management system. American Documentation, 19(3):331–337, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Kent:1955:MLS

[KBLP55] Allen Kent, Madeline M. Berry, Fred U. Luehrs, Jr., and J. W. Perry. Machine literature searching VIII. Operational criteria for designing information retrieval systems. American Documentation, 6(2):93–101, April 1955. CODEN AMDOA7. ISSN 0096-946X.

Kent:1954:MLSa

[KBP54a] Allen Kent, M. M. Berry, and J. W. Perry. Machine literature searching II. Problems in indexing for machine searching. American Documentation, 5(1):22–25, January 1954. CODEN AMDOA7. ISSN 0096-946X.

Kent:1954:MLSb

[KBP54b] Allen Kent, M. M. Berry, and J. W. Perry. Machine literature searching III. Making indexes amenable to machine searching. American Documentation, 5(2):92–94, April 1954. CODEN AM-DOA7. ISSN 0096-946X.

Kline:1969:BCM

[KC69] S. J. Kline and S. K. Cabeen. Brief communications: Motion picture films as research data. American Documentation, 20(4):385–386, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Kent:1958:IED

[Ken58] Allen Kent. Installing electronic data processing, a review. American Documentation, 9(3):214, August 1958. CODEN AMDOA7. ISSN 0096-946X.

Kent:1960:ERI

[Ken60] Allen Kent. Exploitation of recorded information. I. Development of an operational machine searching service for the literature of metallurgy and allied subjects. American Documentation, 11(2): 173–188, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Kenlowrycuss:1962:IP

[Ken62a] Spartacus Kenlowrycuss. The information problem. American Documentation, 13(4):377, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Kent:1962:LE

[Ken62b] Allen Kent. Letters to the Editor. American Documentation, 13 (4):453, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Kent:1964:BCH

[Ken64] Allen Kent. Brief communications: The heuristic information retrieval game. American Documentation, 15(2):150–151, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Kesling:1959:CSW

[Kes59] Robert V. Kesling. Crimes in scientific writing. American Documentation, 10(3):173–175, July 1959. CODEN AMDOA7. ISSN 0096-946X.

Kessler:1963:BCB

[Kes63] M. M. Kessler. Bibliographic coupling between scientific papers. American Documentation, 14(1):10–25, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Kessler:1965:CRB

[Kes65] M. M. Kessler. Comparison of the results of bibliographic coupling and analytic subject indexing. American Documentation, 16(3): 223–233, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Kent:1959:SDT

[KI59] Allen Kent and A. S. Iberall. Soviet documentation. A trip report. American Documentation, 10(1):1–19, January 1959. CODEN AM-DOA7. ISSN 0096-946X.

Kikuchi:1967:BCS

[Kik67] Tohru Kikuchi. Brief communications: Scientific and Technical Information in Japan. American Documentation, 18(4):250–252, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Kilgour:1961:RUB

[Kil61] Frederick G. Kilgour. Recorded use of books in the Yale Medical Library. American Documentation, 12(4):266–269, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Kilgour:1966:LCP

[Kil66] Frederick G. Kilgour. Library catalogue production on small computers. American Documentation, 17(3):124–131, July 1966. CO-DEN AMDOA7. ISSN 0096-946X.

Kingsley:1959:BOA

[Kin59] James Kingsley. Bibliographic organization of American history: Notes on a continuing problem. American Documentation, 10(4): 270–273, October 1959. CODEN AMDOA7. ISSN 0096-946X.

King:1960:ETI

[Kin60] Ellis F. King. Electronic transmission for interlibrary loans. American Documentation, 11(1):32–39, January 1960. CODEN AM-DOA7. ISSN 0096-946X.

Kincaid:1962:RIM

[Kin62] William H. Kincaid. Retrieval of information from medical records. American Documentation, 13(1):83–85, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Kozumplik:1967:CPM

[KL67] W. A. Kozumplik and R. T. Lange. Computer-produced microfilm library catalog. American Documentation, 18(2):67–80, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Klempner:1964:MCA

[Kle64] Irving M. Klempner. Methodology for the comparative analysis of information storage and retrieval systems: a critical review. American Documentation, 15(3):210–216, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Klempner:1969:ISU

[Kle69] Irving M. Klempner. Information science unlimited? … a position paper. American Documentation, 20(4):339–343, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Kolb:1966:PPI

[KMW66] Margaret C. Kolb, Jerome T. Maddock, and Barbara N. Weaver. PICS: the pharmaceutical information control system of Merck Sharp and Dohme Research Laboratories. American Documentation, 17(4):180–185, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Knable:1965:BCE

[Kna65] John P. Knable II. Brief communications: An experiment comparing key words found in indexes and abstracts prepared by humans with those in titles. American Documentation, 16(2):123–124, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Knowles:1960:RLM

[Kno60] Asa S. Knowles. The responsibilities of the library in the midtwentieth century. American Documentation, 11(3):206–209, August 1960. CODEN AMDOA7. ISSN 0096-946X.

Kochen:1969:SGK

[Koc69] Manfred Kochen. Stability in the growth of knowledge. American Documentation, 20(3):186–197, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Koriagin:1964:BCE

[Kor64] Gretchen W. Koriagin. Brief communications: Experience in man and machine relationships in library mechanization. American Documentation, 15(3):227–229, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Kotani:1962:CAJ

[Kot62] Masao Kotani. Communication among Japanese scientists domestically and with their counterparts abroad. American Documentation, 13(3):320–327, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Kent:1956:EDP

[KP56] Allen Kent and James W. Perry. Electronic data processing, a review. American Documentation, 7(4):331–332, October 1956. CO-DEN AMDOA7. ISSN 0096-946X.

Kent:1957:NIA

[KP57] Allen Kent and J. W. Perry. New indexing-abstracting system for formal reports, development and proof services, Aberdeen Proving Ground. American Documentation, 8(1):34–46, January 1957. CO-DEN AMDOA7. ISSN 0096-946X.

Kent:1954:MLSc

[KPB54] Allen Kent, J. W. Perry, and M. M. Berry. Machine literature searching V. Definition and systematization of terminology for code development. American Documentation, 5(3):166–173, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Kraft:1964:CKC

[Kra64] Donald H. Kraft. A comparison of keyword-in-context (KWIC) indexing of titles with a subject heading classification system. American Documentation, 15(1):48–52, January 1964. CODEN AM-DOA7. ISSN 0096-946X.

Kochen:1967:BIB

[KT67] Manfred Kochen and Renata Tagliacozzo. Book-indexes as building blocks for a cumulative index. American Documentation, 18(2):59– 66, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Kuipers:1957:MSD

[KTM57] J. W. Kuipers, A. W. Tyler, and W. L. Myers. A minicard system for documentary information. American Documentation, 8(4):246– 268, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Kuney:1964:LE

[KTW64] Joseph H. Kuney, George L. Trigg, and Harold Wooster. Letters to the Editors. American Documentation, 15(2):152–153, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Kuebler:1966:ONF

[Kue66] Wolf Kuebler. On the optimum number of frames per microfiche. American Documentation, 17(3):132–135, July 1966. CODEN AM-DOA7. ISSN 0096-946X.

Kuipers:1951:MD

[Kui51] J. W. Kuipers. Microcards and documentation. American Documentation, 2(2):89–95, April 1951. CODEN AMDOA7. ISSN 0096-946X.

Kuney:1963:EJP

[Kun63] Joseph H. Kuney. Economics of journal publication. American Documentation, 14(3):238–240, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Kutten:1964:BCU

[Kut64] A. Kutten. Brief communications: U.D.C. abridged Russian edition. American Documentation, 15(3):226, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Kuvshinoff:1967:GGC

[Kuv67] Boris W. Kuvshinoff. A graphic graphics card catalog and computer index. American Documentation, 18(1):3–9, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Kyle:1958:TCS

[Kyl58] Barbara Kyle. Towards a classification for social science literature. American Documentation, 9(3):168–183, August 1958. CO-DEN AMDOA7. ISSN 0096-946X.

Kyle:1963:GES

[Kyl63] Barbara R. F. Kyle. Guest editorial: Some searching questions on documentation research. American Documentation, 14(2):93, April 1963. CODEN AMDOA7. ISSN 0096-946X.

LaHood:1963:BRB

[La 63] Charles La Hood. Book review: Photocopying from bound volumes. By William R. Hawkin. American Library Association, Chicago. 1962. 208 p. ( Library Technology Project Publication, No. 4). American Documentation, 14(1):82–83, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Lamm:1961:PNC

[Lam61] Earl Lamm. Prime-number coding. American Documentation, 12 (3):172–177, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Lamkin:1965:LNa

[Lam65a] Burton E. Lamkin. Literature notes. American Documentation, 16 (3):249–279, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Lamkin:1965:LNb

[Lam65b] Burton E. Lamkin. Literature notes. American Documentation, 16 (4):343–380, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Langenbeck:1962:PRC

[Lan62] Earl H. Langenbeck. A plan to reduce costs of technical library operations in the department of defense. American Documentation, 13(3):295–300, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Lancaster:1965:BRBb

[Lan65a] F. W. Lancaster. Book review: 4/65-3R An evaluation of links and roles used in information retrieval. Thesis presented to the Faculty of the School of Engineering of the Air Force Institute of Technology, Air University. 1963. Lieutenant Colonel Jefferson D. Sinnett. Air Force Institute of Technology, School of Engineering, Wright-Patterson AFB, Ohio. AD 432198. American Documentation, 16 (2):128–129, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Lancaster:1965:BRBa

[Lan65b] F. Wilfred Lancaster. Book review: 1/65-2R The Baseball Program: An Automatic Question-Answerer. 1963. A. K. Wolf, C. S. Chomsky, and B. F. Green, Jr. Massachusetts Institute of Technology, Lincoln Laboratory, (Vol. 1, Lincoln Laboratory Technical Report 306), Lexington, Mass., April 1963. 201 pp. American Documentation, 16(1):39, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Lancaster:1967:BRB

[Lan67] F. W. Lancaster. Book review: 1/67-1R On Retrieval System Theory. 2d edition. 1965. B. C. Vickery. Butterworth, Washington, D.C. 191 pp. American Documentation, 18(1):50, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Lancaster:1968:NRI

[Lan68] F. W. Lancaster. On the need for role indicators in postcoordinate retrieval systems. American Documentation, 19(1):42–46, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Lancaster:1969:MRE

[Lan69a] F. W. Lancaster. MEDLARS: report on the evaluation of its operating efficiency. American Documentation, 20(2):119–142, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Landau:1969:CAD

[Lan69b] Herbert B. Landau. The cost analysis of document surrogation: a literature review. American Documentation, 20(4):302–310, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Larson:1952:UBI

[Lar52] C. Theodore Larson. The U.S. building industry: a case-problem in communications. American Documentation, 3(4):208–213, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Lazowska:1968:PCL

[Laz68] Edward S. Lazowska. Photocopying, copyright, and the librarian. American Documentation, 19(2):123–130, April 1968. CODEN AM-DOA7. ISSN 0096-946X.

Locke:1956:MTL

[LB56] William N. Locke and A. Donald Booth. Machine translation of languages. American Documentation, 7(2):135–136, April 1956. CO-DEN AMDOA7. ISSN 0096-946X.

Livant:1963:ORW

[LB63] William Paul Livant and John Paul Boyd. On the occurrence of reversible words. American Documentation, 14(3):234–237, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Long:1967:DBS

[LBL67] John M. Long, Howard J. Barnhard, and Gertrude C. Levy. Dictionary buildup and stability of word frequency in a specialized medical area. American Documentation, 18(1):21–25, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Lebow:1960:NIP

[LD60] Irving Lebow and Melvin S. Day. New indexing pattern for nuclear science abstracts. American Documentation, 11(2):120–127, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Lecht:1968:NTD

[Lec68] Charles Philip Lecht. No time to document: a case for the book with no beginning, no end, and no author. American Documentation, 19 (3):274–277, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Lee:1952:NDD

[Lee52] Earle V. Lee. New developments in distribution and dissemination. American Documentation, 3(3):176–181, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Leimkuhler:1968:LSF

[Lei68] Ferdinand F. Leimkuhler. A literature search and file organization model. American Documentation, 19(2):131–136, April 1968. CO-DEN AMDOA7. ISSN 0096-946X.

Leith:1969:BLA

[Lei69] John D. Leith, Jr. Biomedical literature: analysis of journal articles collected by a radiation- and cell-biologist. American Documentation, 20(2):143–148, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Leslie:1960:ART

[Les60] W. H. P. Leslie. The automatic retrieval of technical information. American Documentation, 11(3):210–221, August 1960. CODEN AMDOA7. ISSN 0096-946X.

Lesk:1969:WWA

[Les69] M. E. Lesk. Word–word associations in document retrieval systems. American Documentation, 20(1):27–38, January 1969. CO-DEN AMDOA7. ISSN 0096-946X.

Levy:1964:SIP

[Lev64] N. P. Levy. A survey on the information practices of engineers at Western Electric. American Documentation, 15(2):86–88, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Lewicky:1966:BRB

[Lew66] George I. Lewicky. Book review: 7/66–2R Alphabetical Subject Indication of Information. Vol. 3. 1965. John Metcalfe. Rutgers Series on Systems for the Intellectual Organization of Information. Graduate School of Library Service, Rutgers, the State University, New Brunswick, New Jersey. 148 pp. American Documentation, 17(3): 150, July 1966. CODEN AMDOA7. ISSN 0096-946X.

Lewis:1969:GCC

[Lew69] Elizabeth M. Lewis. A graphic catalog card index. American Documentation, 20(3):238–246, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Lubetzky:1969:BDI

[LH69] Seymour Lubetzky and R. M. Hayes. Bibliographic dimensions in information contol. American Documentation, 20(3):247–252, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Libbey:1967:USO

[Lib67] Miles A. Libbey. The use of second order descriptors for document retrieval. American Documentation, 18(1):10–20, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Libaw:1969:OPN

[Lib69] Frieda B. Libaw. Opinion paper: A new, generalized model for information-transfer: a systems approach. American Documentation, 20(4):381–384, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Liepina:1954:SDS

[Lie54] Rita G. Liepina. Scientific documentation in the Soviet Union. American Documentation, 5(2):71–79, April 1954. CODEN AM-DOA7. ISSN 0096-946X.

Lilley:1954:ESC

[Lil54] Oliver L. Lilley. Evaluation of the subject catalog. Criticisms and a proposal. American Documentation, 5(2):41–60, April 1954. CO-DEN AMDOA7. ISSN 0096-946X.

Linn:1966:BRB

[Lin66] Margaret Linn. Book review: 4/66-2R Towards Information Retrieval. 1961. Robert A. Fairthorne. Butterworths, London. 211 pp. American Documentation, 17(2):109, April 1966. CODEN AM-DOA7. ISSN 0096-946X.

Lipetz:1960:SAP

[Lip60] Ben-Ami Lipetz. A successful application of punched cards in subject indexing. American Documentation, 11(3):241–246, August 1960. CODEN AMDOA7. ISSN 0096-946X.

Lipetz:1962:CEC

[Lip62] Ben-Ami Lipetz. Compilation of an experimental citation index from scientific literature. American Documentation, 13(3):251–266, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Lipetz:1963:LCC

[Lip63] Ben-Ami Lipetz. Labor costs, conversion costs, and compatibility in document control systems. American Documentation, 14(2):117– 122, April 1963. CODEN AMDOA7. ISSN 0096-946X.

Lipetz:1965:BCS

[Lip65a] Ben-Ami Lipetz. Brief communications: Some infract, mire on facts. American Documentation, 16(4):338–339, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Lipetz:1965:ISC

[Lip65b] Ben-Ami Lipetz. Improvement of the selectivity of citation indexes to science literature through inclusion of citation relationship indicators. American Documentation, 16(2):81–90, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Little:1964:BRB

[Lit64] Thompson M. Little. Book review: 4/64-000 Selected bibliography on indexing in science and technology: Theory, application and techniques, 1963. Philip L. Dopkowski, ed. Center for Technology and Administration, School of Government and Public Administration, American University (Bull. No. 9), Washington, D.C. 101 pp. $2.50. American Documentation, 15(2):154, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Little:1966:BRB

[Lit66] Thompson M. Little. Book review: 7/66–4R International Affairs. Universal Reference System, Political Science, Government and Public Policy Series. Vol. I. 1965. Metron, Inc., 80 East 11th Street, New York. 1205 pp. American Documentation, 17(3):151, July 1966. CODEN AMDOA7. ISSN 0096-946X.

Loftus:1956:ALA

[LK56] Helen E. Loftus and Allen Kent. Automation in the library — an annotated bibliography. American Documentation, 7(2):110–126, April 1956. CODEN AMDOA7. ISSN 0096-946X.

Luehrs:1955:MLS

[LKPB55] Fred U. Luehrs, Jr., Allen Kent, J. W. Perry, and M. M. Berry. Machine literature searching VII. Machine functions and organization of semantic units. American Documentation, 6(1):33–39, January 1955. CODEN AMDOA7. ISSN 0096-946X.

Lebowitz:1963:LE

[LL63] Abraham I. Lebowitz and Philip Leslie. Letters to the Editor. American Documentation, 14(1):91, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Lancaster:1968:LE

[LL68] F. W. Lancaster and Mrs. Edward S. Lazowska. Letters to the Editors. American Documentation, 19(2):206–207, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Lancaster:1964:TII

[LM64] F. W. Lancaster and J. Mills. Testing indexes and index language devices: The ASLIB Cranfield Project. American Documentation, 15(1):4–13, January 1964. CODEN AMDOA7. ISSN 0096-946X.

Logsdon:1967:BRB

[Log67] Richard H. Logsdon. Book review: 3/67-3R Scientific Management of Library Operations. 1966. Richard M. Dougherty and Fred J. Heinritz. The Scarecrow Press, Inc., New York, 252 pp. American Documentation, 18(3):190–191, July 1967. CODEN AMDOA7. ISSN 0096-946X.

London:1968:PI

[Lon68] Gertrude London. The publication inflation. American Documentation, 19(2):137–141, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Loukopoulos:1966:IPS

[Lou66] Loukas Loukopoulos. Indexing problems and some of their solutions. American Documentation, 17(1):17–25, January 1966. CO-DEN AMDOA7. ISSN 0096-946X.

Lowry:1968:OPS

[Low68] W. Kenneth Lowry. Opinion paper. Science information problems needing solution. American Documentation, 19(3):352–354, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Luhn:1953:NMR

[Luh53] H. P. Luhn. A new method of recording and searching information. American Documentation, 4(1):14–16, January 1953. CODEN AM-DOA7. ISSN 0096-946X.

Luhn:1960:KWC

[Luh60] H. P. Luhn. Key word-in-context index for technical literature (KWIC index). American Documentation, 11(4):288–295, October 1960. CODEN AMDOA7. ISSN 0096-946X.

Luhn:1961:SDN

[Luh61] H. P. Luhn. Selective dissemination of new scientific information with the aid of electronic processing equipment. American Documentation, 12(2):131–138, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Lunin:1969:AIC

[Lun69] Lois F. Lunin. Academic information centers. American Documentation, 20(1):39–49, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Luther:1950:RDS

[Lut50] Frederic Luther. René Dagron and the siege of Paris. American Documentation, 1(4):196–206, October 1950. CODEN AMDOA7. ISSN 0096-946X.

Luther:1951:EEM

[Lut51] Frederic Luther. The earliest experiments in microphotography. American Documentation, 2(3):167–170, August 1951. CODEN AMDOA7. ISSN 0096-946X.

Lynn:1969:QCC

[Lyn69] Kenneth C. Lynn. A quantitative comparison of conventional information compression techniques in dental literature. American Documentation, 20(2):149–151, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Macha:1953:TEI

[Mac53] H. Rowland Macha. A technique for evaluating information analysis methods and personnel. American Documentation, 4(2):35–49, April 1953. CODEN AMDOA7. ISSN 0096-946X.

Mack:1955:IIE

[Mac55] James D. Mack. International information exchange. American Documentation, 6(4):225–227, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Machelson:1958:CDU

[Mac58] Sonya G. Machelson. Chemical documentation in the U.S.S.R. American Documentation, 9(1):20–22, January 1958. CODEN AM-DOA7. ISSN 0096-946X.

Machelson:1960:MTU

[Mac60a] Sonya G. Machelson. Machine translation in the U.S.S.R. — a brief survey. American Documentation, 11(2):151–154, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Machelson:1960:SST

[Mac60b] Sonya G. Machelson. Soviet scientific and technical propaganda. American Documentation, 11(3):237–240, August 1960. CODEN AMDOA7. ISSN 0096-946X.

Mack:1962:EEP

[Mac62a] James D. Mack. Editorial: The Editor’s page. American Documentation, 13(3):357, July 1962. CODEN AMDOA7. ISSN 0096-946X.

MacPherson:1962:BRB

[Mac62b] Harriet D. MacPherson. Book review: Rider’s international classification. By Fremont Rider. Middletown, Connecticut, 1961. Preliminary edition. American Documentation, 13(4):436–437, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Mack:1963:E

[Mac63a] James D. Mack. Editorial. American Documentation, 14(1):92, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Mack:1963:ENP

[Mac63b] James D. Mack. Editorial: A national policy for the information sciences. American Documentation, 14(4):275, October 1963. CO-DEN AMDOA7. ISSN 0096-946X.

Maizell:1960:IGP

[Mai60] Robert E. Maizell. Information gathering patterns and creativity a study of research chemists in an industrial research laboratory. American Documentation, 11(1):9–17, January 1960. CODEN AM-DOA7. ISSN 0096-946X.

Maichel:1961:LE

[Mai61] Karol Maichel. Letters to the Editor. American Documentation, 12 (1):78, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Maloney:1958:RRP

[Mal58] Clifford J. Maloney. A Remington-Rand punched card data retrieval system. American Documentation, 9(1):1–12, January 1958. CODEN AMDOA7. ISSN 0096-946X.

Malin:1962:TR

[Mal62a] Morton V. Malin. Trip report. American Documentation, 13(4): 401–402, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Maloney:1962:SI

[Mal62b] Clifford J. Maloney. Semantic information. American Documentation, 13(3):276–287, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Manly:1961:IVD

[Man61] Ron Manly. The inadequacy of varying the depth of indexing and other “document collection” approaches to information retrieval for researchers. American Documentation, 12(3):204–205, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Manly:1964:DDB

[Man64] Ron Manly. Detailed discussion of Bar-Hillel’s “theoretical aspects of the mechanization of literature searching”. American Documentation, 15(2):126–131, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Mantell:1966:LSA

[Man66] Leroy H. Mantell. On laws of special abilities and the production of scientific literature. American Documentation, 17(1):8–16, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Manacher:1969:BCB

[Man69] Glenn K. Manacher. Brief communications: A benchmark calculation for retrieval by confluence of binary attributes in a randomaccess computer. American Documentation, 20(2):168–169, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Martin:1960:FSL

[Mar60] Robert L. Martin. Foreign scientific literature in translation. American Documentation, 11(2):135–150, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Markus:1962:SAP

[Mar62] John Markus. State of the art of published indexes. American Documentation, 13(1):15–30, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Markus:1966:SAC

[Mar66] John Markus. State of the art of computers in commercial publishing. American Documentation, 17(2):76–88, April 1966. CODEN AMDOA7. ISSN 0096-946X.

Mason:1954:UMC

[Mas54] William T. Mason. The use of microphotography in cataloging technical documents. American Documentation, 5(3):162–165, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Mason:1956:DD

[Mas56] William T. Mason. The dynamics of documentation. American Documentation, 7(1):29–32, January 1956. CODEN AMDOA7. ISSN 0096-946X.

Mattsson:1962:USC

[Mat62] Lennart Mattsson. UNESCO Science Cooperation Office for Southeast Asia and its activities. American Documentation, 13(3):328– 333, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Mulvihill:1968:BCR

[MB68] John Mulvihill and Everett H. Brenner. Brief communications: Ranking boolean search output. American Documentation, 19(2): 204–205, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Marshall:1966:LE

[MBG^{+}66] T. A. Marshall, Jr., Robert L. Birch, Alina Gralewska, Mahlon C. Rhaney, and R. Jordan. Letters to the Editor. American Documentation, 17(3):147–148, July 1966. CODEN AMDOA7. ISSN 0096-946X.

McCann:1961:AML

[McC61] Anne McCann. Applications of machines to library techniques: Periodicals. American Documentation, 12(4):260–265, October 1961. CODEN AMDOA7. ISSN 0096-946X.

McCormick:1962:TUC

[McC62] E. M. McCormick. A trend in the use of computers for information processing. American Documentation, 13(2):182–184, April 1962. CODEN AMDOA7. ISSN 0096-946X.

McCutchen:1965:RCN

[McC65] C. W. McCutchen. Random code numbers for universal identification of documents. American Documentation, 16(2):91–96, April 1965. CODEN AMDOA7. ISSN 0096-946X.

McConnell:1968:CGA

[McC68] William A. McConnell. Computer graphics: an assembly line for information. American Documentation, 19(3):278–285, July 1968. CODEN AMDOA7. ISSN 0096-946X.

McGarry:1958:MVM

[McG58a] Daniel D. McGarry. The microfilming of the Vatican Manuscript Library. American Documentation, 9(1):50–58, January 1958. CO-DEN AMDOA7. ISSN 0096-946X.

McGeever:1958:DAA

[McG58b] Emmett R. McGeever. Documentum ad absurdum: an essay on the directions of documentation research. American Documentation, 9 (2):73–76, April 1958. CODEN AMDOA7. ISSN 0096-946X.

McGeever:1960:LEM

[McG60] Emmett B. McGeever. Letter to the Editor the meaning of “negative meaning”. American Documentation, 11(4):323, October 1960. CODEN AMDOA7. ISSN 0096-946X.

McMahon:1961:UFC

[McM61] George T. McMahon. The use of the filmsort card in industry the American Documentation Institute meeting in miniature. American Documentation, 12(3):219–221, August 1961. CODEN AMDOA7. ISSN 0096-946X.

McMurray:1962:BSR

[McM62] James P. McMurray. The Bureau of Ships rapid selector system. American Documentation, 13(1):66–68, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Mehta:1957:LDS

[Meh57] D. S. Mehta. Library and documentation services in India: Recent developments. American Documentation, 8(1):26–33, January 1957. CODEN AMDOA7. ISSN 0096-946X.

Melton:1962:SCT

[Mel62] John L. Melton. The semantic code today. American Documentation, 13(2):176–181, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Melcher:1968:PWI

[Mel68] Daniel Melcher. Producing the wares of information services. American Documentation, 19(3):270, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Metcalf:1951:CDR

[Met51] Keyes D. Metcalf. Coöperative development of research collections in the United States. American Documentation, 2(1):18–20, January 1951. CODEN AMDOA7. ISSN 0096-946X.

Metcalfe:1965:BRB

[Met65] John Metcalfe. Book review: 7/65-1R SYNTOL: Volume II. 1965. J. C. Gardin. Graduate School of Library Service, Rutgers, the State University (Rutgers Series on Systems for the Intellectual Organization of Information, Susan Artandi Ed.), New Brunswick, New Jersey. American Documentation, 16(3):242, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Maierson:1953:ASB

[MH53] Alvin T. Maierson and W. W. Howell. Application of standard business machine punched-card equipment to metallurgical literature references. American Documentation, 4(1):3–13, January 1953. CODEN AMDOA7. ISSN 0096-946X.

Miller:1952:GCR

[Mil52] E. Eugene Miller. The genesis and characteristics of report literature. American Documentation, 3(2):91–94, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Miller:1953:DAP

[Mil53] Eugene E. Miller. The distribution-acquisition problem in technical reporting. American Documentation, 4(3):95–102, August 1953. CODEN AMDOA7. ISSN 0096-946X.

Milek:1957:AIS

[Mil57] John T. Milek. Abstracting and indexing services in electronics and related electrical fields. American Documentation, 8(1):5–21, January 1957. CODEN AMDOA7. ISSN 0096-946X.

Miller:1968:PI

[Mil68a] George A. Miller. Psychology and information. American Documentation, 19(3):286–289, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Miller:1968:SSE

[Mil68b] Russell R. Miller. A study of searching the eye research literature. American Documentation, 19(3):223–239, July 1968. CODEN AM-DOA7. ISSN 0096-946X.

Minder:1957:FTB

[Min57] Thomas Minder. On the fundamental theories of bibliographic organization. American Documentation, 8(1):1–4, January 1957. CO-DEN AMDOA7. ISSN 0096-946X.

Minder:1960:MFB

[Min60] Thomas Minder. The mathematical foundations of bibliographic organization. American Documentation, 11(1):2–8, January 1960. CODEN AMDOA7. ISSN 0096-946X.

Mitchell:1953:UUA

[Mit53] Herbert F. Mitchell. The use of the Univ AC FAC-tronic system in the library reference field. American Documentation, 4(1):16–17, January 1953. CODEN AMDOA7. ISSN 0096-946X.

Most:1962:ASS

[MK62] S. Most and S. Sim Kessler. An abstract service for specific research areas of high current interest: a preliminary report. American Documentation, 13(1):86–89, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Murdock:1967:GMI

[ML67] John W. Murdock and David M. Liston, Jr. A general model of information transfer: Theme paper 1968 annual convention. American Documentation, 18(4):197–208, October 1967. CODEN AM-DOA7. ISSN 0096-946X.

Milne:1959:FHP

[MM59] Lorus J. Milne and Margery Milne. Foresight and hindsight on a punch-card bibliography. American Documentation, 10(1):78–84, January 1959. CODEN AMDOA7. ISSN 0096-946X.

Montague:1962:PIC

[Mon62] Barbara A. Montague. Patent indexing by concept coordination using links and roles. American Documentation, 13(1):104–111, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Montague:1965:TCE

[Mon65] Barbara A. Montague. Testing, comparison, and evaluation of recall, relevance, and cost of coordinate indexing with links and roles. American Documentation, 16(3):201–208, July 1965. CODEN AM-DOA7. ISSN 0096-946X.

Mooers:1950:EC

[Moo50] Calvin N. Mooers. Editor’s corner. American Documentation, 1(4): 225–229, October 1950. CODEN AMDOA7. ISSN 0096-946X.

Mooers:1951:ZAM

[Moo51] Calvin N. Mooers. Zatocoding applied to mechanical organization of knowledge. American Documentation, 2(1):20–32, January 1951. CODEN AMDOA7. ISSN 0096-946X. See reply [Wis52].

Mooers:1957:CPB

[Moo57] Calvin N. Mooers. Comments on the paper by Bar-Hillel. American Documentation, 8(2):114–116, April 1957. CODEN AMDOA7. ISSN 0096-946X.

Mooers:1960:NTY

[Moo60] Calvin N. Mooers. The next twenty years in information retrieval; some goals and predictions. American Documentation, 11(3):229– 236, August 1960. CODEN AMDOA7. ISSN 0096-946X.

Marshall:1963:LE

[MOP^{+}63] Thomas H. P. Marshall, James L. Olsen, Jr., Robert L. Patrick, T. R. Savage, and J. A. Cordero. Letters to the Editor. American Documentation, 14(3):271–272, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Morris:1954:DCS

[Mor54] Jack C. Morris. The duality concept in subject analysis. American Documentation, 5(3):117–146, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Morsch:1955:TSL

[Mor55] Lucile M. Morsch. Technical services in libraries. American Documentation, 6(2):103–104, April 1955. CODEN AMDOA7. ISSN 0096-946X.

Moran:1960:IMT

[Mor60] Marguerite K. Moran. Indexing of Metal and Thermit Corporation research records on keysort cards. American Documentation, 11(3): 222–228, August 1960. CODEN AMDOA7. ISSN 0096-946X.

Morse:1962:CIR

[Mor62] R. D. Morse. A.I.Ch.E. information-retrieval activities. American Documentation, 13(1):69–70, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Mostecky:1956:SSR

[Mos56] Vaclav Mostecky. Study of the see-also reference structure in relation to the subject of international law. American Documentation, 7(4):294–314, October 1956. CODEN AMDOA7. ISSN 0096-946X.

Moss:1964:CRO

[Mos64] R. Moss. Categories and relations: Origins of two classification theories. American Documentation, 15(4):296–301, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Mount:1965:BRB

[Mou65] Ellis Mount. Book review: Citations of physics literature. (Experimental issue) 1964. Ben-Ami Lipetz. American Institute of Physics, New York. AIP/DRP-CI 1964. American Documentation, 16(3): 242–243, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Mount:1967:BRB

[Mou67] Ellis Mount. Book review: 2/67-3R Information Management in Engineering Education. (Proceedings and recommendations of the Conference on Information Sources, Systems and Media in Engineering Education, held at Lehigh University on May 19–20, 1966). 1966. Robert S. Taylor, Editor. Lehigh University, Bethlehem, Pa.

76 pp. American Documentation, 18(2):116–117, April 1967. CO-DEN AMDOA7. ISSN 0096-946X.

Montgomery:1962:MIP

[MS62] Christine Montgomery and Don R. Swanson. Machinelike indexing by people. American Documentation, 13(4):359–366, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Marron:1966:CDA

[MS66] Harvey Marron and Martin Snyderman, Jr. Cost distribution and analysis in computer storage and retrieval. American Documentation, 17(2):89–95, April 1966. CODEN AMDOA7. ISSN 0096-946X.

Marron:1967:CDA

[MS67] H. Marron and M. Snyderman. Cost distribution and analysis in computer storage and retrieval. II. American Documentation, 18 (3):162–164, July 1967. CODEN AMDOA7. ISSN 0096-946X.

MacMillan:1961:SIP

[MW61] Judith T. MacMillan and Isaac D. Welt. A study of indexing procedures in a limited area of the medical sciences. American Documentation, 12(1):27–31, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Neeland:1962:TRS

[Nee62] Frances Neeland. Treasurer’s report: Statement of cash receipts and disbursements. January 1, 1961 thru October 31, 1961. American Documentation, 13(1):130–131, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Nemec:1957:DCL

[Nem57] Jaroslav Nemec. Documentation of Czechoslovak law. American Documentation, 8(3):181–201, July 1957. CODEN AMDOA7. ISSN 0096-946X.

Newman:1961:CPS

[New61] Simon M. Newman. Classified patent search files, a proposed base for technical information centers. American Documentation, 12(3): 206–212, August 1961. CODEN AMDOA7. ISSN 0096-946X.

Newman:1963:BRB

[New63a] Simon M. Newman. Book review: Basic technical writing. By Herman M. Weisman. Charles E. Merrill Books, Inc., Columbus, Ohio.

1962. 512 pages. $10.60. American Documentation, 14(1):82, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Newman:1963:RBI

[New63b] Simon M. Newman. Reviews: 10/63-2 Proceedings of the 1962 Institute in Technical and Industrial Communications. 1962. Herman W. Weisman, ed. Institute in Technical and Industrial Communications, Colorado State University, Fort Collins, Colo. 163 pp. $5.00. American Documentation, 14(4):312–313, October 1963. CODEN AMDOA7. ISSN 0096-946X.

Newman:1969:BRB

[New69a] Simon M. Newman. Book review: 3/69-4R Thesaurus of information science terminology. 1968. Claire K. Schultz. Rev. Ed. Communication Service Corporation, Washington, D.C. 246 pp. $4.00. American Documentation, 20(3):285, July 1969. CODEN AM-DOA7. ISSN 0096-946X.

Newman:1969:LEB

[New69b] Simon M. Newman. Letters to the Editors: Bibliographies. American Documentation, 20(2):172, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Nitecki:1967:BRB

[Nit67] Andre Nitecki. Book review: 4/67-4R Anglo-American Cataloging Rules. 1967. Prepared by the American Library Association, The Library of Congress, The Library Association and the Canadian Library Association. North American text. American Library Association, Chicago. American Documentation, 18(4):255–257, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Nolan:1959:ISR

[Nol59] J. J. Nolan. Information storage and retrieval using a large scale random access memory. American Documentation, 10(1):27–35, January 1959. CODEN AMDOA7. ISSN 0096-946X.

Nicholson:1958:SJM

[NT58] Natalie N. Nicholson and William Thurston. Serials and journals in the M.I,T. library. American Documentation, 9(4):304–307, October 1958. CODEN AMDOA7. ISSN 0096-946X.

Oates:1968:CTL

[Oat68] James F. Oates, Jr. Convention theme: Levels of interaction between man and information. American Documentation, 19(3):290– 294, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Opler:1959:DCS

[OB59] Ascher Opler and Norma Baird. Display of chemical structural formulas as digital computer output. American Documentation, 10 (1):59–63, January 1959. CODEN AMDOA7. ISSN 0096-946X.

OConnor:1961:SSM

[O’C61] John O’Connor. Some suggested mechanized indexing investigations which require no machines. American Documentation, 12(3): 198–203, August 1961. CODEN AMDOA7. ISSN 0096-946X.

OConnor:1962:SCI

[O’C62a] John O’Connor. The scan-column index. American Documentation, 13(2):204–209, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Orr:1962:SPC

[OC62b] Richard H. Orr and Eleanor M. Crouse. Secondary publication in cardiovascular, endocrine and psychopharmacologic research. American Documentation, 13(2):197–203, April 1962. CODEN AM-DOA7. ISSN 0096-946X.

OConnor:1963:RBL

[O’C63] John O’Connor. Reviews: 4/63-1 Searching legal literature electronically: Results of a test program. 1960. By Jessica S. Melton and Robert C. Bensing. 45 Minnesota Law Review, 229–48. American Documentation, 14(2):167–168, April 1963. CODEN AMDOA7. ISSN 0096-946X.

OConnor:1964:BCT

[O’C64a] John O’Connor. Brief communications: Thesaurus. American Documentation, 15(3):226–227, July 1964. CODEN AMDOA7. ISSN 0096-946X.

OConnor:1964:CIH

[O’C64b] John O’Connor. Correlation of indexing headings and title words in three medical indexing systems. American Documentation, 15(2): 96–104, April 1964. CODEN AMDOA7. ISSN 0096-946X.

OConnor:1967:RDU

[O’C67] John O’Connor. Relevance disagreements and unclear request forms. American Documentation, 18(3):165–177, July 1967. CO-DEN AMDOA7. ISSN 0096-946X.

OConnor:1968:OPS

[O’C68a] John O’Connor. Opinion papers: Some Questions concerning “information need”. American Documentation, 19(2):200–203, April 1968. CODEN AMDOA7. ISSN 0096-946X.

OConnor:1968:RAP

[O’C68b] John O’Connor. Retrieval of answer-providing documents. American Documentation, 19(4):381–386, October 1968. CODEN AM-DOA7. ISSN 0096-946X.

OConnor:1969:SIA

[O’C69] John O’Connor. Some independent agreements and resolved disagreements about answer-providing documents. American Documentation, 20(4):311–319, October 1969. CODEN AMDOA7. ISSN 0096-946X.

OConnor:1968:LE

[ODVB68] John O’Connor, P. S. Davison, James O. Vann, and Charles L. Berner. Letters to the Editors. American Documentation, 19(1): 104–106, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Oatfield:1958:SAS

[OE58] Harold Oatfield and Betty Reynolds Emilio. Some aspects of searching in the pharmaceutical literature reference fringe benefits. American Documentation, 9(4):238–272, October 1958. CODEN AM-DOA7. ISSN 0096-946X.

Oettinger:1968:LI

[Oet68] A. G. Oettinger. Language and information. American Documentation, 19(3):295–298, July 1968. CODEN AMDOA7. ISSN 0096-946X.

OConnor:1966:LE

[OHBS66] John O’Connor, Karl F. Heumann, Harry Baum, and Claire K. Schultz. Letters to the Editor. American Documentation, 17(4): 216–217, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Ohlman:1957:LCP

[Ohl57] Herbert Ohlman. Low-cost production of marginal punched cards on accounting machines. American Documentation, 8(2):123–126, April 1957. CODEN AMDOA7. ISSN 0096-946X.

Opler:1959:AMR

[Opl59] Ascher Opler. On the automatic manipulation of representations of chemical structures. American Documentation, 10(2):130–134, April 1959. CODEN AMDOA7. ISSN 0096-946X.

Orne:1952:MRC

[Orn52] Jerrold Orne. Maintenance and reduction of a collection. American Documentation, 3(2):101–105, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Ornsen:1954:PCV

[Orn54] Kate C. Ornsen. The petroleum chemist’s vernacular. American Documentation, 5(4):218–222, October 1954. CODEN AMDOA7. ISSN 0096-946X.

Orne:1963:RFS

[Orn63] Jerrold Orne. Resources of foreign scientific literature: Acquisition on a national scale. American Documentation, 14(3):229–233, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Orr:1959:IAD

[Orr59] Richard H. Orr. An integrated approach to documentation. American Documentation, 10(3):214–220, July 1959. CODEN AMDOA7. ISSN 0096-946X.

Orr:1961:MNS

[Orr61] Richard H. Orr. The metabolism of new scientific information: a preliminary report. American Documentation, 12(1):15–19, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Orton:1961:USI

[Ort61] Floyd E. Orton. A unified subject index to science literature. American Documentation, 12(1):32–34, January 1961. CODEN AM-DOA7. ISSN 0096-946X.

Oster:1961:P

[Ost61] Gerald Oster. Photoreproduction. American Documentation, 12 (4):285–291, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Overmyer:1962:TPE

[Ove62] Lavahn Overmyer. Test program for evaluating procedures for the exploitation of literature of interest to metallurgists. American Documentation, 13(2):210–222, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Overmyer:1963:AOC

[Ove63] La Vahn Overmyer. An analysis of output costs and procedures for an operational searching service. American Documentation, 14(2): 123–142, April 1963. CODEN AMDOA7. ISSN 0096-946X.

Parthasarathy:1962:IIN

[Par62] S. Parthasarathy. INSDOC (Indian National Scientific Documentation Centre) and its regional activities. American Documentation, 13(3):334–337, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Parker:1966:UPI

[Par66] Edwin B. Parker. The user’s place in an information system. American Documentation, 17(1):26–27, January 1966. CODEN AM-DOA7. ISSN 0096-946X.

Perry:1954:MLSb

[PBK54] J. W. Perry, M. M. Berry, and Allen Kent. Machine literature searching VI. Class definition and code construction. American Documentation, 5(4):238–244, October 1954. CODEN AMDOA7. ISSN 0096-946X.

Pelta:1969:BRB

[Pel69] Harold Nils Pelta. Book review: 2/69-3 Computers in information science. 1968. Susan Artandi. Scarecrow Press, Metuchen, N.J., 145 pp. $4.00. American Documentation, 20(2):174–175, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Perry:1950:IAM

[Per50] James W. Perry. Information analysis for machine searching. American Documentation, 1(3):133–139, August 1950. CODEN AM-DOA7. ISSN 0096-946X.

Perry:1951:SPN

[Per51] James W. Perry. Superimposed punching of numerical codes on hand-sorted punch cards. American Documentation, 2(4):205–212, October 1951. CODEN AMDOA7. ISSN 0096-946X.

Perry:1953:SMF

[Per53] J. W. Perry. Specialized machines of the future for handling broad ranges of subject matter. American Documentation, 4(1):22–26, January 1953. CODEN AMDOA7. ISSN 0096-946X.

Perreault:1965:BCN

[Per65] Jean M. Perreault. Brief communications: A new device for achieving hospitality in array. American Documentation, 16(3):245–246, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Perreault:1966:BRB

[Per66a] Jean M. Perreault. Book review: 10/66-1R Faceted Classification Schemes. 1966. Brian C. Vickery. Rutgers Series on Systems for the Intellectual Organization of Information, vol. 5. Graduate School of Library Service, Rutgers University, New Brunswick, N. J. 108 pp. American Documentation, 17(4):218–219, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Perreault:1966:DRS

[Per66b] Jean M. Perreault. Documentary relevance and structural hierarchy. American Documentation, 17(3):136–139, July 1966. CODEN AMDOA7. ISSN 0096-946X.

Perry:1968:CGC

[Per68] Peter Perry. Combined grouping for coordinate indexes. American Documentation, 19(2):142–145, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Petty:1959:ACA

[Pet59] Gerald M. Petty. Abstracts in Chemical Abstracts of articles and “letters to the Editor” from Physical Review. American Documentation, 10(2):144–150, April 1959. CODEN AMDOA7. ISSN 0096-946X.

Phelps:1958:ESL

[Phe58] Ralph H. Phelps. The engineering societies library. American Documentation, 9(3):165–167, August 1958. CODEN AMDOA7. ISSN 0096-946X.

Piez:1964:BCL

[Pie64] Gladys T. Piez. Brief communications: Library technology project. American Documentation, 15(4):304, October 1964. CODEN AM-DOA7. ISSN 0096-946X.

Pings:1957:OTB

[Pin57] Vern M. Pings. Observations on theories of bibliographic organization. American Documentation, 8(4):292–295, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Pings:1967:SUM

[Pin67] Vern Pings. A study of the use of materials circulated from an engineering library, March to May 1956. American Documentation, 18(3):178–184, July 1967. CODEN AMDOA7. ISSN 0096-946X.

Perry:1956:RMI

[PK56] James W. Perry and Allen Kent. The Russians have a machine for it, a review. American Documentation, 7(3):233–234, July 1956. CODEN AMDOA7. ISSN 0096-946X.

Perry:1954:MLSa

[PKB54] J. W. Perry, Allen Kent, and M. M. Berry. Machine literature searching I. A general approach. American Documentation, 5(1): 18–22, January 1954. CODEN AMDOA7. ISSN 0096-946X.

Perry:1955:MLSa

[PKB55a] J. W. Perry, Allen Kent, and Madeline M. Berry. Machine literature searching IX. Operational functions of automatic equipment. American Documentation, 6(3):166–178, July 1955. CODEN AM-DOA7. ISSN 0096-946X.

Perry:1955:MLSb

[PKB55b] James W. Perry, Allen Kent, and Madeline M. Berry. Machine literature searching X. Machine language; factors underlying its design and development. American Documentation, 6(4):242–254, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Pfeffer:1959:FAP

[PKM59] Harold Pfeffer, Herbert R. Koller, and Ethel C. Marden. A first approach to patent searching procedures on Standard’s Electronic Automatic Computer (SEAC). American Documentation, 10(1): 20–26, January 1959. CODEN AMDOA7. ISSN 0096-946X.

Plankeel:1960:ADM

[Pla60a] F. H. Plankeel. Automation in documentation A mechanised coordinate index system. American Documentation, 11(2):128–134, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Platau:1960:TPA

[Pla60b] Gerard O. Platau. Training of patent abstracters for Chemical Abstracts. American Documentation, 11(1):40–43, January 1960. CO-DEN AMDOA7. ISSN 0096-946X.

Perry:1959:PMS

[PM59] James W. Perry and John L. Melton. Principles of machine and system design with special reference to the indexing and analysis of historical literature. American Documentation, 10(4):278–285, October 1959. CODEN AMDOA7. ISSN 0096-946X.

Pohl:1965:ISF

[Poh65] Frederik Pohl. Information science-fiction or fact? American Documentation, 16(2):101–104, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Polak:1950:ATD

[Pol50] H. Polak. The activities of the Technical Documentation and Information Center for the Netherlands. American Documentation, 1 (3):152–153, August 1950. CODEN AMDOA7. ISSN 0096-946X.

Pollock:1968:MCI

[Pol68] Stephen M. Pollock. Measures for the comparison of information retrieval systems. American Documentation, 19(4):387–397, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Potocko:1968:IST

[Pot68] Richard J. Potocko. An index system of terms with related and associated words. American Documentation, 19(2):146–150, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Poulin:1956:PCS

[Pou56] Roger J. Poulin. A punched card shelf list for discarding books. American Documentation, 7(3):229–230, July 1956. CODEN AM-DOA7. ISSN 0096-946X.

Power:1950:UHR

[Pow50] Eugene B. Power. The use of high reduction microfilm in libraries. American Documentation, 1(3):139–143, August 1950. CODEN AMDOA7. ISSN 0096-946X.

Power:1951:MSB

[Pow51a] Eugene B. Power. Microfilm as a substitute for binding. American Documentation, 2(1):33–39, January 1951. CODEN AMDOA7. ISSN 0096-946X.

Power:1951:MEP

[Pow51b] Eugene B. Power. Microfilm as an edition process of documentary reproduction. American Documentation, 2(3):157–159, August 1951. CODEN AMDOA7. ISSN 0096-946X.

Power:1958:PBL

[Pow58] Eugene Power. O–P books; a library breakthrough. American Documentation, 9(4):273–276, October 1958. CODEN AMDOA7. ISSN 0096-946X.

Pratt:1963:RBA

[Pra63] Lorraine Pratt. Reviews: 7/63-1 Practical administration in public libraries. 1962. Joseph L. Wheeler and Herbert Goldhor. Harper, New York. 571 p. $7.50. American Documentation, 14(3):244, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Prentice:1969:BRH

[Pre69] Ann E. Prentice. Book review: 3/69-3R. H. Scher and W. L. Saunders. Education and training for scientific and technological library and information work. 1968. Her Majesty’s Stationery Office, London. 79 pp. American Documentation, 20(3):284–285, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Price:1969:RCC

[Pri69] C. E. Price. Representing characters to computers. American Documentation, 20(1):50–60, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Perez-Vitoria:1963:DUN

[PV63] A. Perez-Vitoria. Document UNESCO/NS/177 — a code of good practice for scientific publications. American Documentation, 14 (3):241–243, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Prabhavi-Vadhana:1967:BCD

[PV67] Chun Prabhavi-Vadhana. Brief communications: Documentation in Thailand. American Documentation, 18(4):249–250, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Patterson:1958:PCC

[PW58] Helen A. Patterson and Isaac D. Welt. A punch card code for cardiovascular pathology. American Documentation, 9(2):77–83, April 1958. CODEN AMDOA7. ISSN 0096-946X.

Quenzel:1963:SPH

[Que63] Carrol H. Quenzel. Some proposals for handling the information problem — a brief bibliographical essay. American Documentation, 14(2):145–148, April 1963. CODEN AMDOA7. ISSN 0096-946X.

Radziejowski:1969:BRB

[Rad69] Jerzy Radziejowski. Book review: 4/69-2R Referativnyi Zhurnal, Nauchnaya I Tekhnicheskaya Informatsiya, Otdyelnyi Vypusk. 1963. Vsesoyuznyi Institut Nauchnoi i Tekhnicheskoi Informatsii, Moskva. American Documentation, 20(4):388–389, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Randall:1967:BRB

[Ran67] G. E. Randall. Book review: 4/67-3R A Checklist for the Organization, Operation and Evaluation of a Company Library. 2d rev. ed. 1966. Eva Lou Fisher. Special Libraries Association, New York. 61 pp. American Documentation, 18(4):255, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Rauseo:1967:BCT

[Rau67] Michael J. Rauseo. Brief communications: The training implications of automated personnel systems. American Documentation, 18(4):248–249, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Rosenberg:1968:CRK

[RB68] Kenyon C. Rosenberg and Charles L. M. Blocher. A comparison of the relevance of key-word-in-context versus descriptor indexing terms. American Documentation, 19(1):27–29, January 1968. CO-DEN AMDOA7. ISSN 0096-946X.

Rees:1962:RPI

[Ree62] Alan M. Rees. Relevancy and pertinency in indexing. American Documentation, 13(1):93–94, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Rees:1965:ACT

[Ree65] Alan M. Rees. The Aslib–Cranfield test of the Western Reserve University indexing system for metallurgical literature: a review of the final report. American Documentation, 16(2):73–76, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Reid:1954:CS

[Rei54] J. B. Reid. Chronological sigils. American Documentation, 5(1): 26–28, January 1954. CODEN AMDOA7. ISSN 0096-946X.

Resnick:1961:PIR

[Res61] A. Resnick. Part II. The reliability of people in selecting sentences. American Documentation, 12(2):141–143, April 1961. CODEN AM-DOA7. ISSN 0096-946X.

Reynolds:1961:SCG

[Rey61] Walter L. Reynolds. The Senate Committee on Government Operations and Documentation. American Documentation, 12(2):93–97, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Resnick:1963:UDI

[RH63] A. Resnick and C. B. Hensley. The use of diary and interview techniques in evaluating a system for disseminating technical information. American Documentation, 14(2):109–116, April 1963. CODEN AMDOA7. ISSN 0096-946X.

Richmond:1954:SMP

[Ric54] Phyllis Allen Richmond. Some multi-plane classification schemes. American Documentation, 5(2):61–71, April 1954. CODEN AM-DOA7. ISSN 0096-946X.

Richmond:1956:DCT

[Ric56] Phyllis Allen Richmond. A divided catalog — Then what? American Documentation, 7(4):315–319, October 1956. CODEN AM-DOA7. ISSN 0096-946X.

Richmond:1960:HD

[Ric60] Phyllis A. Richmond. Hierarchical definition. American Documentation, 11(2):91–96, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Richmond:1961:RFS

[Ric61] Phyllis A. Richmond. Recommendations for further study in classification. American Documentation, 12(1):35–37, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Richmond:1963:PDP

[Ric63a] Phyllis A. Richmond. A proposal for dual publication of scientific journals. American Documentation, 14(1):54–55, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Richmond:1963:RCP

[Ric63b] Phyllis A. Richmond. Review of the Cranfield Project. American Documentation, 14(4):307–311, October 1963. CODEN AMDOA7. ISSN 0096-946X.

Richmond:1967:BCS

[Ric67] Phyllis A. Richmond. Brief communications: Suggestions on how to read experimental material in information science. American Documentation, 18(4):247–248, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Richmond:1968:BRB

[Ric68] Phyllis A. Richmond. Book review: 3/68–2R PANDEX. 1967. Pandex, Inc., New York. Microfiche, $460 (educational rate, $390); magnetic tape $6,500. American Documentation, 19(3):357–358, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Rider:1950:AMM

[Rid50] Fremont Rider. Archival materials on microcards. American Documentation, 1(1):42–45, January 1950. CODEN AMDOA7. ISSN 0096-946X.

Rider:1951:MVC

[Rid51a] Fremont Rider. Microcards vs. the cost of book storage. American Documentation, 2(1):39–44, January 1951. CODEN AMDOA7. ISSN 0096-946X.

Rider:1951:PCA

[Rid51b] Fremont Rider. The possible correlation of all forms of microtext. American Documentation, 2(3):152–157, August 1951. CODEN AMDOA7. ISSN 0096-946X.

Rees:1958:MSE

[RK58] Janet Rees and Allen Kent. Mechanized searching experiments using the WRU searching selector. American Documentation, 9(4): 277–303, October 1958. CODEN AMDOA7. ISSN 0096-946X.

Rogers:1967:BRB

[Rog67] Frank B. Rogers. Book review: 4/67-2R Symbolic Shorthand System. 1966. (Rutgers Series on Systems for the Intellectual Orgnization of Information, Volume VI.) Hans Selye. New Brunswick. 89 pp. American Documentation, 18(4):254–255, October 1967. CO-DEN AMDOA7. ISSN 0096-946X.

Roper:1968:CBS

[Rop68] Fred W. Roper. A computer-based serials control system for a large biomedical library. American Documentation, 19(2):151–157, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Rosenblum:1965:BCF

[Ros65a] Marcus Rosenblum. Brief communications: Factors of selection in defining science information for the Federal Government. American Documentation, 16(1):33–34, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Ross:1965:BCN

[Ros65b] Laurence W. Ross. Brief communications: A note on value functions. American Documentation, 16(2):122, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Roth:1956:HRC

[Rot56] Robert J. Roth. How readable is chemical literature? American Documentation, 7(3):215–221, July 1956. CODEN AMDOA7. ISSN 0096-946X.

Roy:1957:NPS

[Roy57] K. K. Guha Roy. A new problem in scientific documentation. American Documentation, 8(1):53–55, January 1957. CODEN AMDOA7. ISSN 0096-946X.

Rath:1961:CFT

[RRS61a] G. J. Rath, A. Resnick, and T. R. Savage. Comparisons of four types of lexical indicators of content. American Documentation, 12 (2):126–130, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Rath:1961:FAS

[RRS61b] G. J. Rath, A. Resnick, and T. R. Savage. The formation of abstracts by the selection of sentences. Part I. Sentence selection by men and machines. American Documentation, 12(2):139–141, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Resnick:1964:CHJ

[RS64] A. Resnick and T. R. Savage. The consistency of human judgments of relevance. American Documentation, 15(2):93–95, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Reichardt:1967:CTM

[RSF^{+}67] R. W. Reichardt, S. P. Sher, R. Ford, R. B. Anderson, and E. E. Vogin. Coding and tabulating machine processing of physical signs in toxicity tests. American Documentation, 18(2):81–86, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Rubin:1966:BRB

[Rub66] Audrey Rubin. Book review: 4/66-1R Textbook on Mechanized Information Retrieval. 1962. Allen Kent. Interscience, New York. 268 pp. American Documentation, 17(2):108–109, April 1966. CODEN AMDOA7. ISSN 0096-946X.

Ruhl:1964:CDT

[Ruh64] Mary Jane Ruhl. Chemical documents and their titles: Human concept indexing vs. KWIC-machine indexing. American Documentation, 15(2):136–141, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Rupp:1964:LE

[RVS64] Walter H. Rupp, Charles A. Vertanes, and Daniel Shively. Letters to the Editors. American Documentation, 15(1):76–77, January 1964. CODEN AMDOA7. ISSN 0096-946X.

Sage:1965:AID

[SAF65] C. R. Sage, R. R. Anderson, and D. R. Fitzwater. Adaptive information dissemination. American Documentation, 16(3):185–200, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Sage:1966:CDC

[Sag66] C. R. Sage. Comprehensive dissemination of current literature. American Documentation, 17(4):155–177, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Salton:1963:SHM

[Sal63] Gerard Salton. Some hierarchical models for automatic document retrieval. American Documentation, 14(3):213–222, July 1963. CO-DEN AMDOA7. ISSN 0096-946X.

Salton:1965:LE

[Sal65a] Gerald Salton. Letters to the Editors. American Documentation, 16 (4):340–341, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Salton:1965:EAR

[Sal65b] Gerard Salton. The evaluation of automatic retrieval procedures — selected test results using the SMART system. American Documentation, 16(3):209–222, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Salton:1969:CBM

[Sal69] Gerard Salton. A comparison between manual and automatic indexing methods. American Documentation, 20(1):61–71, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Santiago:1968:SKE

[San68] Antony Santiago. Systematics of knowledge with enneametry and data processing machines. American Documentation, 19(2):158– 162, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Saracevic:1968:LRT

[Sar68a] Tefko Saracevic. Linking research and teaching. American Documentation, 19(4):398–403, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Sarett:1968:SSD

[Sar68b] Lewis H. Sarett. The scientist and scientific data. American Documentation, 19(3):299–304, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Savage:1963:NEM

[Sav63] T. R. Savage. A note on the evaluation of methods for systematically abbreviating English words. American Documentation, 14(1):26–27, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Savage:1965:BCB

[Sav65] Terry R. Savage. Brief communications: A Boyle’s Law for index consistency. American Documentation, 16(1):33, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Savage:1966:BCU

[Sav66] T. R. Savage. Brief communications: Users versus documents. American Documentation, 17(3):141, July 1966. CODEN AM-DOA7. ISSN 0096-946X.

Savage:1967:OPI

[Sav67] T. R. Savage. Opinion paper: The interpretation of SDI data. American Documentation, 18(4):242–246, October 1967. CODEN AMDOA7. ISSN 0096-946X.

Skallerup:1959:BCG

[SB59] Harry R. Skallerup and Angela Battaglia. Bibliographic control in general science publications as illustrated by the transactions of the Illinois State Academy of Science. American Documentation, 10(3): 232–233, July 1959. CODEN AMDOA7. ISSN 0096-946X.

Shank:1966:LE

[SB66] Russell Shank and Robert L. Birch. Letters to the Editors. American Documentation, 17(1):45–46, January 1966. CODEN AM-DOA7. ISSN 0096-946X.

Sargent:1965:LE

[SBF65] Charles W. Sargent, Mrs. Lea M. Bohnert, and S. W. Frank. Letters to the Editors. American Documentation, 16(1):36–37, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Schoendorff:1954:CBD

[Sch54] Ellen Schoendorff. Classification for building documentation. American Documentation, 5(2):79–92, April 1954. CODEN AMDOA7. ISSN 0096-946X.

Schwegmann:1957:RPT

[Sch57] George A. Schwegmann, Jr. The rationale, planning, and technique of the published national union catalog. American Documentation, 8(4):296–299, October 1957. CODEN AMDOA7. ISSN 0096-946X.

Schultz:1962:IAI

[Sch62a] Claire K. Schultz. Inaugural address of the incoming President. American Documentation, 13(1):129–130, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Schultz:1962:IST

[Sch62b] Louise Schultz. The information system: Too big and growing. American Documentation, 13(3):288–294, July 1962. CODEN AM-DOA7. ISSN 0096-946X.

Schultz:1963:GCM

[Sch63] Claire K. Schultz. A generalized computer method for information retrieval. American Documentation, 14(1):39–48, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Schultz:1964:LE

[Sch64a] Claire K. Schultz. Letter to the Editor. American Documentation, 15(4):306, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Schweizer:1964:DDA

[Sch64b] Robert F. E. Schweizer. Direct documentation, analogue documentation, and information-retrieval. American Documentation, 15(1): 23–25, January 1964. CODEN AMDOA7. ISSN 0096-946X.

Schlueter:1969:ISS

[Sch69a] R. A. Schlueter. Information science: Some questions and answers. American Documentation, 20(4):366–368, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Schultz:1969:GEB

[Sch69b] Claire K. Schultz. Guest editorial: Biographical sketches of information scientists. American Documentation, 20(2):109–110, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Scott:1951:RDR

[Sco51] Eugene W. Scott. The role of documentation in research administration. American Documentation, 2(1):44–47, January 1951. CODEN AMDOA7. ISSN 0096-946X.

Scott:1953:NPS

[Sco53] Eugene W. Scott. New patterns in scientific research and publication. American Documentation, 4(3):90–95, August 1953. CODEN AMDOA7. ISSN 0096-946X.

Shera:1955:PPC

[SD55] Jesse H. Shera and Barbara Denison. Preliminary planning conference on information processing and correlation. March 19–20, 1955 summary statement. American Documentation, 6(3):162–166, July 1955. CODEN AMDOA7. ISSN 0096-946X.

Shera:1950:DUS

[SE50a] Jesse H. Shera and Margaret E. Egan. Documentation in the United States. American Documentation, 1(1):8–12, January 1950. CO-DEN AMDOA7. ISSN 0096-946X.

Shera:1950:USR

[SE50b] Jessie H. Shera and Margaret E. Egan. The United States report on national and international bibliographic problems. American Documentation, 1(3):146–151, August 1950. CODEN AMDOA7. ISSN 0096-946X.

Sebring:1953:MPC

[Seb53] M. W. Sebring. A marginal punched card system for a specialized information collection. American Documentation, 4(1):18–22, January 1953. CODEN AMDOA7. ISSN 0096-946X.

Seidell:1956:DMC

[Sei56] Atherton Seidell. Documentation and microfilm copying. American Documentation, 7(3):231–232, July 1956. CODEN AMDOA7. ISSN 0096-946X.

Sewell:1962:ASA

[Sew62] Winifred Sewell. Atherton Seidell: an appreciation (1878–1961). American Documentation, 13(1):123–124, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Seybold:1966:BRB

[Sey66] John W. Seybold. Book review: 10/66-2R Computer Typesetting: Experiments and Prospects. 1965. Michael T. Barnett. MIT Press, Cambridge, Mass. 245 pp. American Documentation, 17(4):219– 220, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Simms:1967:LE

[SFMO67] Daniel M. Simms, D. J. Foskett, Barbara A. Montague, and John O’Connor. Letters to the Editors. American Documentation, 18(3): 185–186, July 1967. CODEN AMDOA7. ISSN 0096-946X.

Schultz:1969:HAD

[SG69] Claire K. Schultz and Paul L. Garwig. History of the American Documentation Institute — a sketch. American Documentation, 20 (2):152–160, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Shaw:1950:HSI

[Sha50] Ralph R. Shaw. High speed intermittent camera. American Documentation, 1(4):194–196, October 1950. CODEN AMDOA7. ISSN 0096-946X.

Shaw:1953:SBH

[Sha53] Ralph R. Shaw. A simple book holder. American Documentation, 4(2):65–66, April 1953. CODEN AMDOA7. ISSN 0096-946X.

Shaw:1954:FMD

[Sha54] Ralph R. Shaw. FID manual on documentation reproduction and selection — A review. American Documentation, 5(3):174–175, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Sharp:1956:EAD

[Sha56] Elmer M. Sharp. Experiences with automatic digital data processing. American Documentation, 7(1):22–28, January 1956. CODEN AMDOA7. ISSN 0096-946X.

Shank:1959:LSR

[Sha59] Russell Shank. Library service to research laboratories of a large university. American Documentation, 10(3):221–223, July 1959. CODEN AMDOA7. ISSN 0096-946X.

Shaw:1961:LE

[Sha61] Ralph R. Shaw. Letters to the Editor. American Documentation, 12(4):305, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Shaw:1962:PMH

[Sha62] Ralph R. Shaw. Parameters for machine handling of alphabetic information. American Documentation, 13(3):267–269, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Shaw:1965:FCR

[Sha65] Ralph R. Shaw. F^{1} a^{1} c^{1} e^{2} i^{2} t^{3} s^{1} n^{3} o^{1} r^{2} m^{1}. American Documentation, 16(2):77–80, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Sharp:1966:BCS

[Sha66] John R. Sharp. Brief communications: The SLIC index. American Documentation, 17(1):41–44, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Shera:1950:UCI

[She50] Jesse H. Shera. The UNESCO Conference on the Improvement of Bibliographic Services. A preliminary report. American Documentation, 1(3):144–146, August 1950. CODEN AMDOA7. ISSN 0096-946X.

Shera:1951:BM

[She51] Jesse H. Shera. Bibliographic management. American Documentation, 2(1):47–54, January 1951. CODEN AMDOA7. ISSN 0096-946X.

Shera:1952:EMM

[She52] J. H. Shera. Effect of machine methods on the organization of knowledge. American Documentation, 3(1):15–20, ???? 1952. CO-DEN AMDOA7. ISSN 0096-946X.

Shera:1953:ENI

[She53] Jesse H. Shera. Emergence of a new institutional structure for the dissemination of specialized information. American Documentation, 4(4):163–173, October 1953. CODEN AMDOA7. ISSN 0096-946X.

Shera:1955:TR

[She55] Jesse H. Shera. The technical report. American Documentation, 6 (2):104–105, April 1955. CODEN AMDOA7. ISSN 0096-946X.

Shera:1956:UL

[She56] Jesse H. Shera. The university library. American Documentation, 7(2):137, April 1956. CODEN AMDOA7. ISSN 0096-946X.

Shera:1959:NTE

[She59] Jesse H. Shera. New tools for easing the burden of historical research. American Documentation, 10(4):274–277, October 1959. CODEN AMDOA7. ISSN 0096-946X.

Shera:1967:BRB

[She67] J. H. Shera. Book review: 2/67-2R Looking Forward in Documentation, Papers and Discussion, Aslib 38th Annual Conference, University of Exeter. 1964. 1966. Aslib, London. 109 pp. American Documentation, 18(2):115–116, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Shilling:1963:RSM

[Shi63] Charles W. Shilling. Requirements for a scientific mission-oriented information center. American Documentation, 14(1):49–53, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Shoffner:1962:TOL

[Sho62] Ralph M. Shoffner. A technique for the organization of large files. American Documentation, 13(1):95–103, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Silver:1950:FC

[Sil50a] Henry Silver. Film composition. American Documentation, 1(4): 177–180, October 1950. CODEN AMDOA7. ISSN 0096-946X.

Silver:1950:POR

[Sil50b] Henry M. Silver. The publication of original research materials. American Documentation, 1(1):13–23, January 1950. CODEN AM-DOA7. ISSN 0096-946X.

Silver:1951:NMP

[Sil51] Henry M. Silver. New methods of printing and reproducing scholarly materials. American Documentation, 2(1):54–58, January 1951. CODEN AMDOA7. ISSN 0096-946X.

Simpson:1962:SIC

[Sim62] G. S. Simpson, Jr. Scientific information centers in the United States. American Documentation, 13(1):43–57, January 1962. CO-DEN AMDOA7. ISSN 0096-946X.

Sprenkle:1963:QSC

[SK63] Peter M. Sprenkle and Frederick G. Kilgour. A quantitative study of characters on biomedical catalogue cards — a preliminary investigation. American Documentation, 14(3):202–206, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Simmons:1964:IDL

[SKM64] Robert F. Simmons, Sheldon Klein, and Keren McConlogue. Indexing and dependency logic for answering English questions. American Documentation, 15(3):196–204, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Smith:1968:POM

[SL68] Michael J. Smith and Richard P. Levy. A physician oriented method for clinical information retrieval. American Documentation, 19(1): 90–94, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Slamecka:1963:CAA

[Sla63] Vladimir Slamecka. Classificatory, alphabetical, and associative schedules as aids in coordinate indexing. American Documentation, 14(3):223–228, July 1963. CODEN AMDOA7. ISSN 0096-946X.

Slamecka:1964:MCE

[Sla64] Vladimir Slamecka. Machine compilation and editing of printed alphabetical subject indexes. American Documentation, 15(2):132– 135, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Slamecka:1969:PIP

[Sla69] Vladimir Slamecka. Part II papers generated by the conference on the methodology of curricular studies. American Documentation, 20(4):355–357, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Slote:1968:AWC

[Slo68] Stanley J. Slote. An approach to weeding criteria for newspaper libraries. American Documentation, 19(2):168–172, April 1968. CO-DEN AMDOA7. ISSN 0096-946X.

Simmons:1963:MDI

[SM63] Robert F. Simmons and Keren L. McConlogue. Maximum-depth indexing for computer retrieval of English language data. American Documentation, 14(1):68–73, January 1963. CODEN AMDOA7. ISSN 0096-946X.

Simpson:1967:BCS

[SM67] Gustavus S. Simpson, Jr. and John W. Murdock. Brief communications: $’s and Secrets. American Documentation, 18(2):110, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Smith:1955:LBT

[Smi55] Julian F. Smith. Literature’s buried treasures. American Documentation, 6(2):86–88, April 1955. CODEN AMDOA7. ISSN 0096-946X.

Snyder:1966:BRI

[Sny66] Samuel S. Snyder. Book review: 7/66-IR INTREX: Report of a Planning Conference on Information Transfer Experiments. September 3, 1965. Carl F. J. Overhage and R. Joyce Harman, ed. MIT Press, Cambridge, Mass. 276 pp. American Documentation, 17(3):149–150, July 1966. CODEN AMDOA7. ISSN 0096-946X.

Sophar:1962:PM

[Sop62] Gerald J. Sophar. A presidential message. American Documentation, 13(1):126–128, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Sophar:1968:DLF

[Sop68] Gerald J. Sophar. The determination of legal facts and economic guideposts with respect to the dissemination of scientific and educational information as it is affected by copyright — A status report. American Documentation, 19(3):317–321, July 1968. CODEN AM-DOA7. ISSN 0096-946X.

Speedie:1960:SW

[Spe60] Gordon Alexander Speedie. Speed words. American Documentation, 11(1):23–31, January 1960. CODEN AMDOA7. ISSN 0096-946X.

Spencer:1967:SSS

[Spe67] Carol C. Spencer. Subject searching with Science Citation Index: Preparation of a drug bibliography using Chemical Abstracts, Index Medicus, and Science Citation Index 1961 and 1964. American Documentation, 18(2):87–96, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Schless:1968:RGI

[SPS68] Arthur P. Schless, Marie E. Purdy, and Robert E. Stevenson. Reading guide and index to the cancer-virology literature. American Documentation, 19(2):163–167, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Stone:1968:BCS

[SR68] Don C. Stone and Morris Rubinoff. Brief communications: Statistical generation of a technical vocabulary. American Documentation, 19(4):411–412, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Szebenyi-Sigmond:1959:LIS

[SS59] Judith Szebenyi-Sigmond. Libraries and information services behind the iron curtain. American Documentation, 10(2):108–115, April 1959. CODEN AMDOA7. ISSN 0096-946X.

Schultz:1961:CAM

[SS61] Claire K. Schultz and Clayton A. Shepherd. A computer analysis of the Merck Sharp & Dohme Research Laboratories indexing system. American Documentation, 12(2):83–92, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Schultz:1962:GCM

[SS62] Claire K. Schultz and Phyllis A. Schwartz. A generalized computer method for index production. American Documentation, 13(4):420– 432, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Schultz:1965:CIT

[SSO65] Claire K. Schultz, Wallace L. Schultz, and Richard H. Orr. Comparative indexing: Terms supplied by biomedical authors and by document titles. American Documentation, 16(4):299–312, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Schultz:1961:CDU

[SSS61] Claire K. Schultz, Phyllis D. Schwartz, and Leon Steinberg. A comparison of dictionary use within two information retrieval systems. American Documentation, 12(4):247–253, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Stanto:1961:BRB

[Sta61] Herbert A. Stanto. Book review: Introducere în documentarea ştiinţifică (Introduction to scientific documentation). by Aurel Avramescu and Virgil Candea. Academy of the Rumanian People’s Republic, Bucharest, 1960. 519 pp. Illus. American Documentation, 12(2):144, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Stevens:1955:RFD

[Ste55] David H. Stevens. The role of the foundation in documentation. American Documentation, 6(2):57–62, April 1955. CODEN AM-DOA7. ISSN 0096-946X.

Stevens:1961:CST

[Ste61] Norman D. Stevens. A comparative study of three systems of information retrieval: a summary. American Documentation, 12(4): 243–246, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Stein:1969:BRB

[Ste69] Rose Stein. Book review: 4/69-3R Klassifikationssysteme und Thesauri; eine Anleitung zur Herstellung von Klassifikationssystemen und Thesauri im Bereich der Dokumentation. (Classifica (Classification systems and thesauri; a guide for the building of classification systems and thesauri in the field of documentation.)) Soergel, Dagobert. 1969. Hrsg. v d. Deutschen Gesellschaft fur Dokumentation, e.V., Frankfurt a/M. 224 p. illus., flow diagrs., tables. Bibliography, sources, glossary. American Documentation, 20(4):389, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Stiles:1958:ICV

[Sti58] H. Edmund Stiles. Identification of the conditions for valid application of machines to bibliographic control. American Documentation, 9(1):42–49, January 1958. CODEN AMDOA7. ISSN 0096-946X.

Stiassny:1960:MAV

[Sti60] Simon Stiassny. Mathematical analysis of various superimposed coding methods. American Documentation, 11(2):155–169, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Stone:1969:ERC

[Sto69] C. Walter Stone. Educational resources and curriculum development. American Documentation, 20(4):332–334, October 1969. CO-DEN AMDOA7. ISSN 0096-946X.

Strada:1956:DCG

[Str56] Francis Strada. The decimal classification in Germany. American Documentation, 7(1):1–8, January 1956. CODEN AMDOA7. ISSN 0096-946X.

Strieby:1957:PLC

[Str57] Irene M. Strieby. The problem of literature chemists in industry. American Documentation, 8(2):89–97, April 1957. CODEN AM-DOA7. ISSN 0096-946X.

Sturdivant:1969:LES

[Stu69] Clarence A. Sturdivant. Letters to the Editor: Search laundry. American Documentation, 20(3):281, July 1969. CODEN AM-DOA7. ISSN 0096-946X.

Sable:1967:LE

[SVS^{+}67] Martin H. Sable, Charles A. Vertanes, Herbeet A. Simon, W. T. Brandhorst, and Howard P. Iker. Letters to the Editors. American Documentation, 18(2):112–114, April 1967. CODEN AMDOA7. ISSN 0096-946X.

Swanson:1965:BCR

[Swa65] Rowena Swanson. Brief communications: Report on a conference on the education of science information personnel held July 27–28, 1964, in Cleveland, Ohio. American Documentation, 16(1):34–35, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Swanson:1968:BCI

[Swa68] Rowena W. Swanson. Brief communications: Information services for small-scale industry. American Documentation, 19(4):412–413, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Swanson:1969:BCU

[Swa69] Rowena W. Swanson. Brief communications: User-oriented information systems. American Documentation, 20(3):279–280, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Starke:1968:GDS

[SWCT68] A. C. Starke, F. R. Whaley, E. C. Carson, and W. B. Thompson. GAF document storage and retrieval system. American Documentation, 19(2):173–180, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Swets:1969:EIR

[Swe69] John A. Swets. Effectiveness of information retrieval methods. American Documentation, 20(1):72–89, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Tague:1965:MQA

[Tag65] Jean Tague. Matching of question and answer terminology in an education research file. American Documentation, 16(1):26–32, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Tareev:1962:MDS

[Tar62] Boris M. Tareev. Methods of disseminating scientific information, and science information activities in the USSR. American Documentation, 13(3):338–343, July 1962. CODEN AMDOA7. ISSN 0096-946X.

Tate:1950:AM

[Tat50a] Vernon D. Tate. An appraisal of microfilm. American Documentation, 1(2):91–99, April 1950. CODEN AMDOA7. ISSN 0096-946X.

Tate:1950:IBD

[Tat50b] Vernon D. Tate. Introducing American Documentation. A quarterly review of ideas, techniques, problems and achievements in documentation. American Documentation, 1(1):3–7, January 1950. CODEN AMDOA7. ISSN 0096-946X.

Tate:1950:SNF

[Tat50c] Vernon D. Tate. Special notice to former subscribers of the Journal of Documentary Reproduction. American Documentation, 1(1):1–2, January 1950. CODEN AMDOA7. ISSN 0096-946X.

Tate:1951:F

[Tat51] Vernon D. Tate. Foreword. American Documentation, 2(1):iii–iv, January 1951. CODEN AMDOA7. ISSN 0096-946X.

Tate:1954:AMN

[Tat54] Vernon D. Tate. Annual meeting, National Microfilm Association. American Documentation, 5(3):173–174, August 1954. CODEN AMDOA7. ISSN 0096-946X.

Taube:1952:PCW

[Tau52a] Mortimer Taube. Possibilities for cooperative work in subject controls. American Documentation, 3(1):21–28, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Taube:1952:SLD

[Tau52b] Mortimer Taube. Special librarianship and documentation. American Documentation, 3(3):166–167, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Taube:1953:IPO

[Tau53] Mortimer Taube. Implications for professional organization and training. American Documentation, 4(3):122–125, August 1953. CODEN AMDOA7. ISSN 0096-946X.

Taube:1961:BRB

[Tau61a] Mortimer Taube. Book review: Automatic language translation. By Anthony G. Oettinger. Cambridge, Massachusetts: Harvard University Press, 1960. Pp. xix + 380. $10.00. American Documentation, 12(1):62–63, January 1961. CODEN AMDOA7. ISSN 0096-946X.

Taube:1961:NUR

[Tau61b] Mortimer Taube. Notes on the use of roles and links in coordinate indexing. American Documentation, 12(2):98–100, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Taube:1962:NEW

[Tau62] Mortimer Taube. A note on the evaluation of the WRU semantic code as an example of generic coding. American Documentation, 13(2):185–186, April 1962. CODEN AMDOA7. ISSN 0096-946X.

Taube:1965:BRB

[Tau65a] Mortimer Taube. Book review: 1/65-1R Natural Language and the Computer. 1963. Paul L. Garvin, ed. McGraw-Hill, New York. 398 pp. $8.95. American Documentation, 16(1):38–39, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Taube:1965:NPM

[Tau65b] Mortimer Taube. A note on the pseudo-mathematics of relevance. American Documentation, 16(2):69–72, April 1965. CODEN AM-DOA7. ISSN 0096-946X.

Tauber:1968:BCI

[Tau68] Stephen J. Tauber. Brief communications: Imprecision — problems for information processing. American Documentation, 19(4):413– 414, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Taylor:1962:PAQ

[Tay62] Robert S. Taylor. The process of asking questions. American Documentation, 13(4):391–396, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Thompson:1968:PSD

[TBW68] David A. Thompson, Lawrence A. Bennigson, and David Whitman. A proposed structure for displayed information to minimize search time through a data base. American Documentation, 19(1):80–84, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Tinker:1964:BCC

[TEC64] William Tinker, Benjamin Evers, and Paul N. Chance. Brief communications: A case history in selective elimination as a solution to the information crises. American Documentation, 15(4):304–305, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Taube:1952:UTC

[TGW52] Mortimer Taube, C. D. Gull, and Irma S. Wachtel. Unit terms in coordinate indexing. American Documentation, 3(4):213–218, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Tucci:1969:DFT

[TH69] Valerie A. Tucci and Norman Horrocks. Dictionary of facsimile transmission terminology. American Documentation, 20(3):280, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Thomas:1962:RTC

[Tho62] William Thomas. Report of the Teller’s Committee. American Documentation, 13(1):132, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Thompson:1968:OI

[Tho68] Frederick B. Thompson. The organization is the information. American Documentation, 19(3):305–308, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Tinker:1966:IMM

[Tin66] John F. Tinker. Imprecision in meaning measured by inconsistency of indexing. American Documentation, 17(2):96–102, April 1966. CODEN AMDOA7. ISSN 0096-946X.

Tinker:1968:IIP

[Tin68] John F. Tinker. Imprecision in indexing. Part II. American Documentation, 19(3):322–330, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Taylor:1969:BCC

[TJ69] Leonard B. Taylor and Gerald G. Johnson, Jr. Brief communications: Computerized numeric proofreading for a Linofilm phototypesetter. American Documentation, 20(1):94–97, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Tauber:1962:PMI

[TM62] Alfred S. Tauber and Wilbur C. Myers. Photochromic microimages, a key to practical microdocument storage and dissemination. American Documentation, 13(4):403–409, October 1962. CO-DEN AMDOA7. ISSN 0096-946X.

Tyler:1955:AKM

[TMK55] A. W. Tyler, W. L. Myers, and J. W. Kuipers. The application of the Kodak minicard system to problems of documentation. American Documentation, 6(1):18–30, January 1955. CODEN AMDOA7. ISSN 0096-946X.

Tancredi:1968:APT

[TN68] Samuel A. Tancredi and Owen D. Nichols. Air pollution technical information processing — the microthesaurus approach. American Documentation, 19(1):66–70, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Tober:1959:DOB

[Tob59a] Edward P. Tober. Determining optimal back number inventories. American Documentation, 10(3):224–227, July 1959. CODEN AM-DOA7. ISSN 0096-946X.

Tober:1959:POL

[Tob59b] Edward P. Tober. Print ordering for learned journals. American Documentation, 10(4):316–318, October 1959. CODEN AMDOA7. ISSN 0096-946X.

Taube:1955:SNL

[TPBK55] Mortimer Taube, J. W. Perry, Madeline M. Berry, and Allen Kent. Some notes on logical products and associations. American Documentation, 6(4):239–241, October 1955. CODEN AMDOA7. ISSN 0096-946X.

Trejo:1966:BRB

[Tre66] Arnulfo D. Trejo. Book review: 1/66-4R Technical Dictionary of Librarianship, English–Spanish. 1964. Beatriz Massa de Gil, Ray Trautman. and Peter Goy. Editorial F. Trillas, S. A., Mexico. 387 pp. American Documentation, 17(1):49, January 1966. CODEN AMDOA7. ISSN 0096-946X.

Treu:1968:BRG

[Tre68] Siegried Treu. The browser’s retrieval game. American Documentation, 19(4):404–410, October 1968. CODEN AMDOA7. ISSN 0096-946X.

Tritschler:1964:EIS

[Tri64] R. J. Tritschler. Effective information-searching strategies without “perfect” indexing. American Documentation, 15(3):179–184, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Trueswell:1965:QMU

[Tru65] Richard W. Trueswell. A quantitative measure of user circulation requirements and its possible effect on stack thinning and multiple copy determination. American Documentation, 16(1):20–25, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Truelson:1967:BRB

[Tru67] Stanley D. Truelson, Jr. Book review: 3/67-2R National Library of Medicine Current Catalog. Volume 1. Jan. 1–14, 1966. U.S. Public Health Service, Washington, D.C. Biweekly, cumulated quarterly from the first of the current year; annual cumulation case-bound. Supersedes National Library of Medicine Catalog. Sold by the Superintendent of Documents: 1966 price, $15.00 a year ($20.00 foreign) with annual cumulation also sold separately for $4.50; price increase to be announced. American Documentation, 18(3):189–190, July 1967. CODEN AMDOA7. ISSN 0096-946X.

Tryon:1969:BRB

[Try69] Jonathan S. Tryon. Book review: 3/69-1R Annual review of information science and technology. Volume 3. 1968. Carlos A. Cuadra, Editor. American Society for Information Science. Published by Encyclopaedia Britannica, Inc. 457 pp. $15.00. American Documentation, 20(3):283, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Tumbleson:1953:RDI

[Tum53] Robert Tumbleson. Restricted dissemination of information and its social implications. American Documentation, 4(3):115–122, August 1953. CODEN AMDOA7. ISSN 0096-946X.

Taube:1953:LSC

[TW53] Mortimer Taube and Irma S. Wachtel. The logical structure of coordinate indexing. American Documentation, 4(2):67–68, April 1953. CODEN AMDOA7. ISSN 0096-946X.

Tyaganatarajan:1961:SDC

[Tya61] T. Tyaganatarajan. A study in the developments of colon classification. American Documentation, 12(4):270–278, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Uhlmann:1964:ARS

[Uhl64] W. Uhlmann. The application of random superimposed coding and chain spelling to peek-a-boo cards. American Documentation, 15 (2):89–92, April 1964. CODEN AMDOA7. ISSN 0096-946X.

Uhlmann:1968:DSS

[Uhl68] Wolfram Uhlmann. Document specification and search strategy using basic intersections and the probability measure of sets. American Documentation, 19(3):240–246, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Uhlmann:1969:GMM

[Uhl69] Wolfram Uhlmann. A general method for matching arbitrary logical statements in mechanized retrieval systems. American Documentation, 20(3):253–258, July 1969. CODEN AMDOA7. ISSN 0096-946X.

Ullman:1953:CTP

[Ull53] Morris B. Ullman. Contemporary trends in the production and use of social data. American Documentation, 4(4):137–146, October 1953. CODEN AMDOA7. ISSN 0096-946X.

Urquhart:1951:AI

[Urq51] D. J. Urquhart. American impressions. American Documentation, 2(2):100–102, April 1951. CODEN AMDOA7. ISSN 0096-946X.

Urquhart:1956:DPU

[Urq56] D. J. Urquhart. The documentation problem in the United Kingdom. American Documentation, 7(4):276–282, October 1956. CO-DEN AMDOA7. ISSN 0096-946X.

Vaillancourt:1967:BRBa

[Vai67a] Pauline M. Vaillancourt. Book review: 1/67-2R L’Automatisation des Récherches Documentaires: Un Modèle Général — le Syntol. 1964. R. C. Cros, J. C. Gardin, and F. Lévy. Gauthier-Villars, Paris. 260 pp. American Documentation, 18(1):50, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Vaillancourt:1967:BRBb

[Vai67b] Pauline M. Vaillancourt. Book review: 1/67-3R L’Organization de la Documentation Scientifique. 1964. J. C. Gardin, E. de Grolier and F. Levéry. Gauthier-Villars, Paris. 269 pp. American Documentation, 18(1):51, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Vanby:1963:MDD

[Van63] L. Vanby. A minor Devil’s documentation dictionary. American Documentation, 14(2):143–144, April 1963. CODEN AMDOA7. ISSN 0096-946X.

VanLuik:1965:BCC

[Van65] James Van Luik. Brief communications: The Chemical Abstracts KWIC index. American Documentation, 16(2):122–123, April 1965. CODEN AMDOA7. ISSN 0096-946X.

Vassilion:1969:BRB

[Vas69] Epsevia Vassilion. Book review: 4/69-1R Computer-based retrieval systems. 1969. Bernard Houghton (Ed.) Archon Books and Clive Bingley, Hamden, Connecticut. 136 pp. No. index. $5.00. American Documentation, 20(4):388, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Venezky:1968:SRE

[Ven68] Richard L. Venezky. Storage, retrieval, and editing of information for a dictionary. American Documentation, 19(1):71–79, January 1968. CODEN AMDOA7. ISSN 0096-946X.

Vickery:1951:SCM

[Vic51] Brian C. Vickery. Some comments on mechanical selection. American Documentation, 2(2):102–107, April 1951. CODEN AMDOA7. ISSN 0096-946X.

Vickery:1959:ISR

[Vic59a] B. C. Vickery. Information storage and retrieval, a review. American Documentation, 10(4):319–320, October 1959. CODEN AMDOA7. ISSN 0096-946X.

Vickery:1959:SSC

[Vic59b] B. C. Vickery. The structure of “semantic coding,” a review. American Documentation, 10(3):234–241, July 1959. CODEN AMDOA7. ISSN 0096-946X.

Vickery:1960:MMA

[Vic60] B. C. Vickery. The Marlis a multi-aspect relevance linkage information system. — Present position and future needs. American Documentation, 11(2):97–101, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Virgo:1969:LE

[VMB69] Julie A. Virgo, Russell R. Miller, and W. T. Brandhorst. Letters to the Editors. American Documentation, 20(1):98–99, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Voress:1965:IPT

[Vor65] Hugh E. Voress. Improvements in a permuted title index. American Documentation, 16(2):97–100, April 1965. CODEN AMDOA7. ISSN 0096-946X.

VanCott:1968:EBS

[VZ68] Harold P. Van Cott and Albert Zavala. Extracting the basic structure of scientific literature. American Documentation, 19(3):247– 262, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Weisman:1962:WA

[WA62] Herman M. Weisman and Maurice L. Albertson. Water Abstracts. American Documentation, 13(1):71–82, January 1962. CODEN AMDOA7. ISSN 0096-946X.

Wachtel:1952:PCI

[Wac52] Irma Wachtel. A punched card index for nuclear data. American Documentation, 3(1):56–57, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Wadington:1958:UCC

[Wad58] John P. Wadington. Unit concept coordinate indexing. American Documentation, 9(2):107–113, April 1958. CODEN AMDOA7. ISSN 0096-946X.

Waddell:1968:BRB

[Wad68] Samuel J. Waddell. Book review: 4/68–4R Union catalog of medical periodicals. Listing medical and paramedical periodicals in existence in 1950 and new titles published since, with Holdings of 68 Libraries in the New York Metropolitan Area, as of December 31, 1966. 1967. medical Library Center of New York. 589 pp. $10.00 to members of the Center or to those outside the metropolitan New York area. American Documentation, 19(2):209–210, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Wheeler:1958:SCO

[WAF^{+}58] Keith W. Wheeler, E. R. Andrews, F. Fallon, G. L. Krueger, F. P. Palopoli, and E. L. Schumann. A structure code for organic compounds. American Documentation, 9(3):198–207, August 1958. CO-DEN AMDOA7. ISSN 0096-946X.

Wagner:1960:DDT

[Wag60] Frank S. Wagner, Jr. A dictionary of documentation terms. American Documentation, 11(2):102–119, April 1960. CODEN AMDOA7. ISSN 0096-946X.

Waahlin:1966:CST

[Wåh66] Ejnar Wåhlin. Classification Systems and Their Subjects. A general analysis of different kinds of classification systems characterized by different types of subject. American Documentation, 17(4):199–215, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Waite:1969:LEN

[Wai69] David P. Waite. Letters to the Editors: Nasa thesaurus. American Documentation, 20(2):172, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Wall:1967:RAI

[Wal67] Eugene Wall. A rationale for attacking information problems. American Documentation, 18(2):97–103, April 1967. CODEN AM-DOA7. ISSN 0096-946X.

Wall:1968:PAI

[Wal68] Eugene Wall. Possibilities of articulation of information systems into a network. American Documentation, 19(2):181–187, April 1968. CODEN AMDOA7. ISSN 0096-946X.

Wall:1969:VBC

[Wal69] Eugene Wall. Vocabulary building and control techniques. American Documentation, 20(2):161–164, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Warheit:1952:BIO

[War52] I. A. Warheit. Bibliographic identification and organization. American Documentation, 3(2):105–110, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Warheit:1956:ELT

[War56] I. A. Warheit. Evaluation of library techniques for the control of research materials. American Documentation, 7(4):267–275, October 1956. CODEN AMDOA7. ISSN 0096-946X.

Ward:1966:BRB

[War66] Edith Ward. Book review: 4/66-3R Indexing Theory, Indexing Methods and Search Devices. 1964. Frederick Jonker. Scarecrow, New York. 124 pp. American Documentation, 17(2):109–111, April 1966. CODEN AMDOA7. ISSN 0096-946X.

Watt:1968:H

[Wat68] W. C. Watt. Habitability. American Documentation, 19(3):338– 351, July 1968. CODEN AMDOA7. ISSN 0096-946X.

Way:1968:SHA

[Way68] William Way. Subject heading authority list, computer prepared. American Documentation, 19(2):188–199, April 1968. CODEN AM-DOA7. ISSN 0096-946X.

Weil:1961:CDR

[WBO^{+}61] B. H. Weil, E. H. Brenner, Kate Ornsen, Rita Paddock, Jurgen Raymond, and Sidney Winn. Coordination of documentation research and information services by the American Petroleum Institute. American Documentation, 12(1):56–60, January 1961. CO-DEN AMDOA7. ISSN 0096-946X.

Weinstein:1959:ACI

[WD59] Shirley Jane Weinstein and Raymond J. Drozda. Adaptation of coordinate indexing system to a general literature and patent file: Machine posting. American Documentation, 10(2):122–129, April 1959. CODEN AMDOA7. ISSN 0096-946X.

Weaver:1955:POP

[Wea55] Warren Weaver. The patent office problem. American Documentation, 6(3):129–133, July 1955. CODEN AMDOA7. ISSN 0096-946X.

Weber:1958:TRP

[Web58] John Weber. The technical reports program of the technology division of the office of technical services. American Documentation, 9(4):308–311, October 1958. CODEN AMDOA7. ISSN 0096-946X.

Weber:1965:BRB

[Web65] David C. Weber. Book review: 7/65-3R Catalog Card Reproduction; Report on a Study. 1965. George Fry and Associates. American Library Association (Library Technology Project Publication #9), Chicago. 81 pp. American Documentation, 16(3):243–244, July 1965. CODEN AMDOA7. ISSN 0096-946X.

Weinstein:1966:BDP

[Wei66] S. Jane Weinstein. Biological dictionary preparation, control, and maintenance. American Documentation, 17(4):190–198, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Welt:1961:BRB

[Wel61] Isaac D. Welt. Book review: The National Library of Medicine index mechanization project. 1961. National Library of Medicine U.S. Department of Health, Education, and Welfare Washington, D.C. American Documentation, 12(2):144, April 1961. CODEN AMDOA7. ISSN 0096-946X.

Welt:1964:E

[Wel64] Isaac D. Welt. Editorial. American Documentation, 15(4):249, October 1964. CODEN AMDOA7. ISSN 0096-946X.

Welt:1965:E

[Wel65] Isaac D. Welt. Editorial. American Documentation, 16(4):282, October 1965. CODEN AMDOA7. ISSN 0096-946X.

Welt:1966:E

[Wel66] Isaac D. Welt. Editorial. American Documentation, 17(4):154, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Wepsiec:1961:DP

[Wep61] Jan Wepsiec. Documentation in Poland. American Documentation, 12(4):279–284, October 1961. CODEN AMDOA7. ISSN 0096-946X.

Warheit:1963:LE

[WFY^{+}63] I. A. Warheit, R. A. Fairthorne, S. Yagil, Rinehart S. Potts, Mortimer Taube, and Kingsbury T. Jackson. Letter to Editor. American Documentation, 14(2):163–166, April 1963. CODEN AMDOA7. ISSN 0096-946X.

Waldo:1958:RRW

[WGP58] W. H. Waldo, R. S. Gordon, and J. D. Porter. Routine report writing by computer. American Documentation, 9(1):28–31, January 1958. CODEN AMDOA7. ISSN 0096-946X.

Whaley:1961:MNI

[Wha61] Fred R. Whaley. The manipulation of nonconventional indexing systems. American Documentation, 12(2):101–107, April 1961. CO-DEN AMDOA7. ISSN 0096-946X.

White:1966:BRB

[Whi66] Herbert S. White. Book review: 7/66–3R Principles of Automated Information Retrieval. 1965. William F. Williams. The Business Press. Elmhurst, Ill. 439 pp. American Documentation, 17(3):150– 151, July 1966. CODEN AMDOA7. ISSN 0096-946X.

Wilcox:1951:PVL

[Wil51a] Jerome K. Wilcox. The point of view of the librarian. American Documentation, 2(3):162–166, August 1951. CODEN AMDOA7. ISSN 0096-946X.

Wilson:1951:PLU

[Wil51b] Baxter D. Wilson. The polygraph in the library of the university of Virginia. American Documentation, 2(2):97–100, April 1951. CO-DEN AMDOA7. ISSN 0096-946X.

Wilcox:1952:RCT

[Wil52] Harold E. Wilcox. Reaching the consumer through the field representative and the research information center. American Documentation, 3(3):155–160, ???? 1952. CODEN AMDOA7. ISSN 0096-946X.

Wilde:1969:ISD

[Wil69] Daniel U. Wilde. Iterative strategy design. American Documentation, 20(1):90–91, January 1969. CODEN AMDOA7. ISSN 0096-946X.

Wise:1952:MWC

[Wis52] Carl S. Wise. Multiple word coding vs. random coding for the rapid selector. A reply to Calvin N. Mooers. American Documentation, 3 (4):223–225, ???? 1952. CODEN AMDOA7. ISSN 0096-946X. See [Moo51].

Wolfe:1959:SCI

[Wol59] William Wolfe. Some comments on the IRIA retrieval system. American Documentation, 10(1):74–77, January 1959. CODEN AMDOA7. ISSN 0096-946X.

Wood:1957:DBC

[Woo57] G. Congdon Wood. A detailed biology code for storing, retrieving, and correlating chemical-biological data. American Documentation, 8(3):168–180, July 1957. CODEN AMDOA7. ISSN 0096-946X.

Wooster:1964:BCO

[Woo64] Harold Wooster. Brief communications: Ode to illiteracy. American Documentation, 15(3):226, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Worley:1969:BCU

[Wor69] William E. Worley. Brief communications: Use of an on-line, videodisplay computer system in medical research. American Documentation, 20(2):165–166, April 1969. CODEN AMDOA7. ISSN 0096-946X.

Wise:1950:MCR

[WP50] Carl S. Wise and James W. Perry. Multiple coding and the rapid selector. American Documentation, 1(2):76–83, April 1950. CODEN AMDOA7. ISSN 0096-946X.

Westbrook:1962:ASM

[WP62] J. H. Westbrook and A. J. Peat. An analysis of Soviet metallurgical literature. American Documentation, 13(4):367–376, October 1962. CODEN AMDOA7. ISSN 0096-946X.

Weinstein:1964:BSC

[WS64] Edward A. Weinstein and Joan Spry. Boeing SLIP: Computer produced and maintained printed book catalogs. American Documentation, 15(3):185–190, July 1964. CODEN AMDOA7. ISSN 0096-946X.

Wildhack:1954:DI

[WSS54] William A. Wildhack, Joshua Stern, and Julian Smith. Documentation in instrumentation. American Documentation, 5(4):223–237, October 1954. CODEN AMDOA7. ISSN 0096-946X.

Wilkinson:1967:BCD

[WW67] W. A. Wilkinson and W. H. Waldo. Brief communications: A decentralized national chemical information system. American Documentation, 18(1):47–48, January 1967. CODEN AMDOA7. ISSN 0096-946X.

Yagi:1965:ATM

[Yag65] Eri Yagi. An application of a type of matrix to analyze citations of scientific papers. American Documentation, 16(1):10–19, January 1965. CODEN AMDOA7. ISSN 0096-946X.

Yovits:1969:IST

[Yov69] M. C. Yovits. Information science: Toward the development of a true scientific discipline. American Documentation, 20(4):369–376, October 1969. CODEN AMDOA7. ISSN 0096-946X.

Zunde:1969:ICQ

[ZD69] Pranas Zunde and Margaret E. Dexter. Indexing consistency and quality. American Documentation, 20(3):259–267, July 1969. CO-DEN AMDOA7. ISSN 0096-946X.

Zunde:1967:DIT

[ZS67] Pranas Zunde and Vladimir Slamecka. Distribution of indexing terms for maximum efficiency of information transmission. American Documentation, 18(2):104–108, April 1967. CODEN AMDOA7. ISSN 0096-946X.